### v1

# Analysis of Semi-Supervised Learning on Hypergraphs

Abstract
Hypergraphs provide a natural framework for modeling higher-order interactions, yet their theoretical underpinnings in semi-supervised learning remain limited. We provide an asymptotic consistency analysis of variational learning on random geometric hypergraphs, precisely characterizing the conditions ensuring the well-posedness of hypergraph learning as well as showing convergence to a weighted $p$-Laplacian equation. Motivated by this, we propose Higher-Order Hypergraph Learning (HOHL), which regularizes via powers of Laplacians from skeleton graphs for multiscale smoothness. HOHL converges to a higher-order Sobolev seminorm. Empirically, it performs strongly on standard baselines.
Keywords and phrases. hypergraphs, non-parametric regression, semi-supervised learning, asymptotic consistency, multiscale problems
Mathematics Subject Classification. 49J55, 49J45, 62G20, 65N12
1 Introduction
Given a set of $n$ feature vectors $\Omega_{n}=\{x_{i}\}_{i=1}^{n}\subset\Omega\subset\mathbb{R}^{d}$, where we assume $x_{i}\overset{\text{iid}}{\sim}\mu\in\mathcal{P}(\Omega)$, and a set of binary labels $\{y_{i}\}_{i=1}^{N}\subset\{0,1\}$, the goal of semi-supervised learning is to infer the missing labels for $\{x_{i}\}_{i=N+1}^{n}$ by leveraging both labeled and unlabeled data. Graph-based learning (see, e.g., [61, 16] for non-deep learning methods and [91, 53] for deep learning approaches) has become a widely used framework for this task, owing to the ability of graphs to effectively encode the underlying geometry of the data. Typically, one constructs a graph $G_{n}=(\Omega_{n},W_{n})$, where the nodes correspond to the feature vectors $\Omega_{n}$ and $W_{n}=(w_{ij})_{i,j=1}^{n}$ is an edge-weight matrix reflecting pairwise similarities. A variational problem is then posed on this graph to recover a labeling function on the vertices [99, 80, 90, 8]. When $W_{n}$ is not intrinsic to the the data, the main subtlety lies in the edge-construction. For most applications, the assumption of graph homophily is made, i.e. connected points should share similar labels and therefore, one uses a local approach for the design of $W_{n}$. Most prominently, $k$-nearest neighbour graphs [85, 17] or random geometric graphs [68, 80, 25, 90] are employed.
We now detail the latter model. Given a parameter $\varepsilon$ and a non-increasing function $\eta:[0,\infty)\mapsto[0,\infty)$, weights $w_{\varepsilon,ij}$ between vertices $x_{i}$ and $x_{j}$ are defined as
$w_{\varepsilon,ij}=\eta\left(\frac{|x_{i}-x_{j}|}{\varepsilon}\right).$
The canonical choice of $\eta$ is $\mathds{1}_{[0,1]}$ which ensures that only points within an $\varepsilon$-distance are connected. In order to capture the geometry in a more delicate manner, one can go beyond pairwise interactions and decide to more strongly connect all points within the same $\varepsilon$-ball: e.g. if $x_{0},x_{1},x_{2}$ are all connected pairwise, then we could add an additional connection for the tuple $(x_{0},x_{1},x_{2})$. This observation leads to the considerations of hypergraphs (see Figure 1). Another alternative is, for example, to consider multiscale Laplace learning as in [61] and we refer to Section 3.3 for a thorough discussion about the link between multiscale Laplace learning and hypergraph learning.
Figure 1: From graphs to hypergraphs. Left: In the graph, the vertices $v_{1}$, $v_{2}$, and $v_{3}$ are all connected pairwise. Right: A single hyperedge is added connecting all three vertices, transitioning from a graph to a hypergraph representation.
A hypergraph $G$ is defined as $G=(V,E)$ where $V$ is a set of objects and $E$ a family of subsets ee—called hyperedges—of $V$. Hypergraphs extend graphs by allowing hyperedges to connect arbitrary subsets of nodes, and hypergraph-based methods have found success in various areas of science as in [32, 92, 72, 50, 74, 69, 21, 31, 94, 65, 96, 58, 28]. We denote the weight of hyperedge $e$ by $w_{0}(e)$ and its degree/size by $|e|$. In our setting, we will be considering $V=\Omega_{n}$. Learning on hypergraphs is, for example, considered in [96] where the following optimization problem (we keep the scaling with the edge degree but remove the vertex degree below) is proposed as a relaxation of the hypergraph cut problem (similar to what is done in [85] for the graph setting): feature vectors are clustered based on the value of the function $u$ which solves
(1) $\operatorname*{\mathrm{argmin}}_{v:\Omega_{n}\to\mathbb{R}}\sum_{e\in E}\sum_{\{x_{i},x_{j}\}\subseteq e}\frac{w_{0}(e)}{|e|}(v(x_{i})-v(x_{j}))^{2}\quad\text{such that $\sum_{x_{r}\in\Omega_{n}}v(x_{r})^{2}=1$ and $\sum_{x_{r}\in\Omega_{n}}v(x_{r})=0$.}$
Continuing the graph analogy, it is also shown that $u$ is an eigenvector of the hypergraph Laplacian matrix. Similarly to what is considered in [99] (Laplace learning) or [80] ($p$-Laplacian learning), it is natural to extend (1) to the semi-supervised learning regime, namely to consider the solution to
(2) $\operatorname*{\mathrm{argmin}}_{v:\Omega_{n}\to\mathbb{R}}\sum_{e\in E}\sum_{\{x_{i},x_{j}\}\subseteq e}\frac{w_{0}(e,x_{i},x_{j})}{|e|}(v(x_{i})-v(x_{j}))^{2}\quad\text{such that $v(x_{i})=y_{i}$ for $i\leq N$}$
where the weight function $w_{0}$ now can also depend on the vertices $x_{i}$ and $x_{j}$.
Several points are apparent from the formulation (2), which naturally decomposes into two components: the mechanism of label interaction, encoded via the finite difference term, and the support of these interactions, governed by the hypergraph’s hyperedge weights and connectivity pattern.
  1. P.1
    Since the dependence on $v$ in the objective function arises solely from pairwise interactions between vertices, the resulting regularization is inherently first-order. Consequently, the hypergraph itself contributes only to defining the support of interactions—that is, which and how strongly vertex pairs are connected—not to specifying the order or nature of the regularization being applied. We show in Theorems 3.3 and 3.4 that the solution of (2), in an appropriate limit, behaves as the sum of first order energies.
  2. P.2
    Given the previous point, the hyperedge construction is essential to the success of semi-supervised learning as in (2). We refer to [32] for a review of commonly used weight models in hypergraph learning.
  3. P.3
    In order to go beyond simple pairwise interaction, in Section 3.3, we propose another hypergraph learning model that penalizes higher-order derivatives similar to what can be found in [25, 90] for the graph setting. By doing so, we also establish a link between hypergraph learning and multiscale Laplace learning [61].
Figure 2: Classification of several algorithms based on their continuum limit. Edges indicate convergence to a Sobolev-type seminorm in the continuum limit, with each algorithm linked to its associated $\mathrm{W}^{(k,p)}$ space. Methods from left-to-right: [18], [99], [96], [26], [74], [25], [61], this work.
In order to make these insights rigorous, we provide a new consistency analysis of hypergraph learning (2) in the large-data regime. Asymptotic consistency analysis is a widely used framework for studying graph-based learning algorithms, see [7, 23, 45, 78, 82, 14]. It serves two main purposes: (1) identifying the continuum limit as $n\to\infty$ enables a principled classification of algorithmic behavior—see Figure 2 for a visual summary of several algorithms and their associated continuum limits; (2) analyzing the limiting continuum problem provides insights into fundamental properties of the discrete model, such as well-posedness and ill-posedness, i.e. convergence to nontrivial labelling functions or constants, on large datasets (see for example [80, 25, 90]). In contrast, consistency analysis for hypergraph-based regularization methods remains largely unexplored, with the exception of [74], and this work addresses that gap.
Consistency analysis between discrete energies $\mathcal{E}_{n}$ defined for discrete functions $v_{n}:\Omega_{n}\mapsto\mathbb{R}$ and an appropriate continuum energy $\mathcal{E}_{\infty}$ defined for continuum functions $v:\Omega\mapsto\mathbb{R}$, can be performed in various ways:
  • in spectral convergence [7, 33, 17, 86, 67, 79, 87], one analyses the convergence of the eigenpairs of the discrete operator appearing in $\mathcal{E}_{n}$ to the eigenpairs of the corresponding continuum operator in $\mathcal{E}_{\infty}$;
  • in pointwise convergence [7, 23, 45, 49, 47, 78, 82], one studies whether $\mathcal{E}_{n}(v\lfloor_{\Omega_{n}})$ converge to $\mathcal{E}_{\infty}(v)$ as $n\to\infty$ and $\varepsilon\to 0$, for a sufficiently smooth function $v:\Omega\to\mathbb{R}$—an alternative approach is to analyze the corresponding operators in the Euler–Lagrange equations and study their convergence instead [89];
  • in variational convergence [14, 24, 25, 37, 40, 39, 38, 80, 81, 83], one is interested in the convergence of minimizers of $\mathcal{E}_{n}$ to the minimizers of $\mathcal{E}_{\infty}$.
In this paper, we study the consistency of (2) through the lens of the latter two complementary notions of convergence. Pointwise convergence allows us to show that the operator appearing in the discrete Euler–Lagrange equation of (2) converges to a non-trivial continuum operator, which behaves qualitatively like a $p$-Laplace operator. Variational convergence, on the other hand, enables a precise characterization of the well- and ill-posedness of hypergraph learning in the semi-supervised setting, as a function of the hypergraph construction parameters. The appropriate analytical framework for convergence of minimizers is $\Gamma$-convergence [11], which requires a common metric space in which both discrete and continuum functions can be embedded. To this end, we rely on the $\mathrm{TL}^{p}$ space introduced in [37], which is based on tools from optimal transport [84, 73]. We review these methodologies in Section 2.
A notable observation is that the non-trivial continuum operator obtained via pointwise convergence differs from the operator arising in the Euler–Lagrange equation of the limiting energy obtained through variational convergence. To the best of our knowledge, this phenomenon has not been previously documented in the context of graph-based algorithms. This discrepancy highlights a subtle but important distinction between the two modes of convergence and underscores the need to consider both when analyzing the continuum limits of graph- or hypergraph-based learning models.
Another conclusion of our analysis is that hypergraph learning, in its standard variational formulation, is inherently a first-order method. This confirms our intuition discussed in Point P.1. Indeed, both the non-trivial operator obtained via pointwise convergence and the limiting energy identified through variational convergence are associated with first-order regularization: the former behaves like a $p$-Laplace operator, while the latter corresponds to a weighted Sobolev $\mathrm{W}^{1,p}$-seminorm. To move beyond first-order regularization and capture richer geometric structures encoded by the hypergraph, we therefore propose a new approach called Higher-Order Hypergraph Learning (HOHL) that penalizes higher-order derivatives of the labeling function. This is achieved by introducing powers of graph Laplacians computed at multiple scales, each induced by the underlying hypergraph structure.
We also study the variational convergence of the proposed HOHL model and provide a rigorous characterization of its well- and ill-posedness. This analysis also shows that HOHL converges to a genuinely higher-order Sobolev-type energy in the continuum limit, thereby distinguishing it fundamentally from classical hypergraph learning frameworks (see Figure 2) and addressing Point P.3. Furthermore, we establish a formal connection between HOHL and multiscale Laplacian learning, showing that our method can be interpreted as a principled extension of existing multiscale approaches. We support this interpretation with empirical results that demonstrate the effectiveness of HOHL in leveraging higher-order regularization for semi-supervised learning.
1.1 Contributions
Our main contributions (summarized in Table 1) are as follows:
  1. 1.
    Pointwise Convergence to Continuum Operator: We establish the pointwise convergence of the discrete Euler–Lagrange operators arising in hypergraph learning to a non-trivial continuum operator that behaves like a $p$-Laplacian.
  2. 2.
    Well-/Ill-Posedness via Variational Convergence: We characterize the well- and ill-posedness of classical hypergraph learning via variational convergence, explicitly as a function of the hypergraph construction parameters.
  3. 3.
    Mismatch Between Pointwise and Variational Limits: We show that the limiting operator obtained through pointwise convergence differs from the variational limit, highlighting a previously undocumented distinction between these two modes of convergence in graph-based algorithms.
  4. 4.
    Classical Hypergraph Learning is First-Order: We demonstrate that both the limiting operator and the limiting energy in classical hypergraph learning correspond to first-order regularization, thereby confirming that this framework is inherently first-order and effectively reduces to reweighted graph-based learning.
  5. 5.
    HOHL and Multiscale Higher-Order Regularization: We introduce the HOHL framework, which explicitly penalizes higher-order derivatives of the labeling function. Furthermore, we establish a formal link between HOHL and multiscale Laplacian learning and characterize the well-/ill-posedness of this model via variational convergence.
  6. 6.
    Empirical Validation of HOHL: We empirically validate HOHL on standard semi-supervised learning benchmarks, demonstrating its ability to exploit hypergraph geometric information.
  7. 7.
    Continuum-Based Framework for Algorithm Classification: We propose a principled framework for comparing and classifying graph- and hypergraph-based learning algorithms via their continuum limits. This asymptotic perspective reveals structural similarities between methods that may appear distinct in the discrete setting, as illustrated in Figure 2.
Classical Hypergraph learning (2) HOHL (5)
Pointwise convergence Theorem 3.3
Well/Ill-posedness characterization Theorem 3.4 Theorem 3.5
First-order method
Higher-order method
Table 1: Summary of contributions.
1.2 Related works
Laplace learning [99] is a widely used algorithm defined on undirected graphs and has served as a foundational tool for label propagation and semi-supervised learning. Its success has led to numerous extensions, including to directed graphs [95, 48] and hypergraphs [96]. However, it has been observed that Laplace learning performs poorly in low-label regimes, especially when the number of unlabeled data points is large [64, 26]. The asymptotic consistency analysis in [80] provides a theoretical explanation for this phenomenon and, by precisely characterizing the well-posedness of the algorithm, allows practitioners to circumvent limitations of Laplace learning (or of its variants – in the case of [80], the authors consider the more general $p$-Laplacian method). The work presented in this paper is partly concerned with the analogous analysis for the hypergraph learning algorithm of [96].
Our first contribution is to establish a continuum limit and characterize the well- and ill-posedness of hypergraph learning in the semi-supervised setting. This can be viewed as a continuation of recent efforts in discrete-to-continuum analysis of graph-based methods. For instance, the authors in [37] deal with the asymptotic consistency of graph total variation for clustering problems. Other examples include graph cut problems and their continuum counterparts in [40, 39, 35] and [66], the Mumford-Shah functional in [20] and an application in empirical risk minimization [34]. In the semi-supervised setting, the most relevant papers are [80, 90] dealing with $p$-Laplacian learning and fractional Laplacian regularization respectively. The latter two methods were developed as an alternative to Laplace learning [99], in order to loosen the strict well-posedness requirements of Laplace learning (other variants for this purpose are Lipschitz learning [55, 15, 71, 13], game theoretic $p$-Laplacian regularization [14], Poisson learning [16, 12], re-weighting of the graph Laplacian in [75, 76, 77, 18] or truncated energies in [5, 6]). Extending these graph-based techniques to the hypergraph setting remains an important avenue for future research.
Our second contribution is an asymptotic characterization of the role of hypergraph structure in semi-supervised learning. Previous works demonstrated, that hypergraphs can offer meaningful advantages over graphs — see, for example, [93, 65], where the richer combinatorial structure of hypergraphs is argued to better capture complex geometric relationships. This has led to the development of a variety of hypergraph-based learning methodologies [96, 28, 58, 32, 50, 72], often accompanied by discrete probabilistic or combinatorial comparisons with graph-based analogues (e.g., [3, 51, 22, 52, 63]). In contrast, our work focuses on the asymptotic regime, where $n\to\infty$, and studies the continuum limits of hypergraph learning objectives. To the best of our knowledge, such asymptotic analysis has only been carried out in [74], and only for their specific hypergraph regularization model. Extending this line of analysis to other hypergraph learning methodologies mentioned above remains an important and largely open direction. We believe that such asymptotic characterizations are highly valuable: as illustrated in Figure 2, identifying continuum limits provides a principled framework for comparing and classifying a broad range of algorithms that may appear quite different at the discrete level. Notably, our analysis reveals that certain hypergraph models, while distinct at the discrete level, asymptotically behave like graph-based models, revealing underlying structural equivalences.
We note that asymptotic analysis of hypergraphs has also been studied in the context of stochastic block models, where the hypergraph is itself a random object sampled from a generative process conditioned on the node labels [43, 44]. This setting is fundamentally different from ours which focuses on the regularization of label functions on a fixed hypergraph.
Our proposed HOHL model builds on prior work on nonlocal Laplacian-based regularization [25, 90, 98, 61], and introduces higher-order penalties by applying powers of the graph Laplacian at multiple scales induced by hypergraph structure. This unifies and generalizes ideas that have emerged in the graph neural network literature [42, 2, 62]. Although (hyper)graph neural networks (see [9] for a recent review) achieve strong empirical performance, they typically require extensive supervision, hyperparameter tuning, and often lack theoretical guarantees. In contrast, HOHL provides a theoretically grounded, architecture-free alternative that is well-suited for low-label regimes.
On the technical side, our proofs rely on nonlocal approximation results for $\mathrm{W}^{1,p}$ initially formulated in [10] and then considered under the $\Gamma$-convergence lens in [70]. The latter were extended to the discrete-to-continuum setting in [37, 80]. For other discrete-to-continuum results approximating $\mathrm{W}^{s,2}$ we refer to [25, 90, 36]. Finally, we note that our results are purely asymptotic. Discrete-to-continuum rates have for example been obtained on similar graph learning problems in [19, 89, 27].
The remainder of the paper is structured as follows: in Section 2, we review the theoretical foundations of our approach; in Section 3, we state our main results; in Section 4, we provide detailed proofs; in Section 5, we present numerical experiments; in Section 6, we conclude and propose directions for future research.
2 Background
In this section, we introduce the $\mathrm{TL}^{p}$ topology and provide a brief overview of $\Gamma$-convergence. These two notions are essential for studying variational convergence, which concerns the convergence of minimizers of our machine learning objective. This perspective is particularly relevant in the semi-supervised learning setting, where the minimizers are often the primary objects of interest. In practice, label predictions are frequently obtained by thresholding the minimizer: for instance, if the discrete minimizer $u_{n}$ takes real values and the observed labels satisfy $\{y_{i}\}_{i=1}^{N}\subset\{0,1\}$, then for $i>N$, predicted labels are set via thresholding as $y_{i}=\mathds{1}_{[0,0.5]}(u_{n}(x_{i}))$. For more than two classes the methodology is extended using 1-hot encoding.
2.1 The $\mathrm{TL}^{p}$ Space
Let $\mathcal{P}(\Omega)$ be the set of probability measures on $\Omega$ and $\mathcal{P}_{p}(\Omega)$ be the set of probability measures on $\Omega$ with finite $p$th-moment. We denote by $\mathrm{L}^{p}(\mu)$ the set of functions $u$ that are measurable with respect to $\mu$ and such that $\int_{\Omega}|u(x)|^{p}\,\mathrm{d}x<+\infty$ . The pushforward of a measure $\mu\in\mathcal{P}(\Omega)$ by a map $T:\Omega\to\mathcal{Z}$ is the measure $\nu\in\mathcal{P}(\mathcal{Z})$ defined by
$\nu(A)=T_{\#}\mu(A):=\mu(T^{-1}(A))=\mu\left(\left\{x\,|\,T(x)\in A\right\}\right)\qquad\text{for all measurable sets }A.$
For $\mu,\nu\in\mathcal{P}_{p}(\Omega)$ we denote by $\Pi(\mu,\nu)$ the set of all probability measures on $\Omega\times\Omega$ such that the first marginal is $\mu$ and the second marginal is $\nu$, i.e. $(P_{X})_{\#}\pi=\mu$ and $(P_{Y})_{\#}\pi=\nu$ where $P_{X}:\Omega\times\Omega\ni(x,y)\mapsto x\in\Omega$ and $P_{Y}:\Omega\times\Omega\ni(x,y)\mapsto y\in\Omega$. The following definition of the $\mathrm{TL}^{p}$ space and metric can be found in [37].
Definition 2.1.
For an underlying domain $\Omega$, define the set
$\mathrm{TL}^{p}=\left\{(\mu,u)\,|\,\mu\in\mathcal{P}_{p}(\Omega),u\in\mathrm{L}^{p}(\mu)\right\}.$
For $(\mu,u),(\nu,v)\in\mathrm{TL}^{p}$, we define the $\mathrm{TL}^{p}$ distance $d_{\mathrm{TL}^{p}}$ as follows:
$d_{\mathrm{TL}^{p}}((\mu,u),(\nu,v))=\inf_{\pi\in\Pi(\mu,\nu)}\left(\int_{\Omega\times\Omega}|x-y|^{p}+|u(x)-v(y)|^{p}\,\mathrm{d}\pi(x,y)\right)^{\frac{1}{p}}.$
The $\mathrm{TL}^{p}$ distance is related to the $p$-Wasserstein [73, 84] distance between the measures $\mu$ and $\nu$ and we refer to [37] for more details. In particular, from the latter property, we can characterize convergence in the $\mathrm{TL}^{p}$ space as follows.
Proposition 2.2.
[37, Proposition 3.12] Let $(\mu,u)\in\mathrm{TL}^{p}$ where $\mu$ is absolutely continuous with respect to Lebesgue measure and let $\{(\mu_{n},u_{n})\}_{n=1}^{\infty}$ be a sequence in $\mathrm{TL}^{p}$. The following are equivalent:
  1. 1.
    $(\mu_{n},u_{n})$ converges to $(\mu,u)$ in $\mathrm{TL}^{p}$;
  2. 2.
    $\mu_{n}$ converges weakly to $\mu$ and there exists a sequence of transport maps $\{T_{n}\}_{n=1}^{\infty}$ with $(T_{n})_{\#}\mu=\mu_{n}$ and $\int_{\Omega}|x-T_{n}(x)|\,\mathrm{d}x\to 0$ such that
    $\int_{\Omega}|u(x)-u(T_{n}(x))|^{p}\,\mathrm{d}\mu(x)\to 0;$
In order to compare discrete functions to continuum ones, we will let $\mu_{n}$ be the empirical measures of our samples $\{x_{i}\}_{i=1}^{n}$ and $\mu$ will be the measure from which the points are sampled. Furthermore, $\{u_{n}\}_{n=1}^{\infty}$ and $u$ will respectively be the minimizers of our discrete and continuum objectives. In order to use the above result, we need certain transport maps $T_{n}$ whose existence is guaranteed by the next theorem [41, Theorem 1.1].
Theorem 2.3 (Existence of transport maps).
Let $\Omega\subset\mathbb{R}^{d}$ be open, connected and bounded with Lipschitz boundary. Let $\mu$ be a probability measure on $\Omega$ with a density that is bounded above and below by positive constants. Let $x_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mu\in\mathcal{P}(\Omega)$ and we denote the empirical measure of $\{x_{i}\}_{i=1}^{n}$ by $\mu_{n}$. Then, there exists a constant $C>0$ such that $\mathbb{P}$-a.s., there exists a sequence of transport maps $\{T_{n}:\Omega\mapsto\Omega_{n}\}_{n=1}^{\infty}$ from $\mu$ to $\mu_{n}$ such that:
$\begin{cases}\limsup_{n\to\infty}\frac{n^{1/2}\|\mathrm{Id}-T_{n}\|_{\mathrm{L}^{\infty}}}{\log(\log(n))}\leq C&\text{if }d=1;\\ \limsup_{n\to\infty}\frac{n^{1/2}\|\mathrm{Id}-T_{n}\|_{\mathrm{L}^{\infty}}}{\log(n)^{3/4}}\leq C&\text{if }d=2;\\ \limsup_{n\to\infty}\frac{n^{1/d}\|\mathrm{Id}-T_{n}\|_{\mathrm{L}^{\infty}}}{\log(n)^{1/d}}\leq C&\text{if }d\geq 3.\end{cases}$
The probability measure $\mathbb{P}$ is defined in Section 3. In terms of the assumptions we introduce later, the conditions in the above theorem are given by S.1, M.1, M.2 and D.1.
2.2 $\Gamma$-Convergence
The appropriate framework to describe the convergence of variational problems is $\Gamma$-convergence from the calculus of variations. We only recall the key properties used in this paper similarly to what be found in [37, 80, 90] and refer to [11] for more details.
Definition 2.4.
Let $(Z,d_{Z})$ be a metric space and $F_{n}:Z\to\mathbb{R}$ a sequence of functionals. We say that $F_{n}$ $\Gamma$-converges to $F$ with respect to $d_{Z}$ if:
  1. 1.
    For every $z\in Z$ and every sequence $\{z_{n}\}$ with $d_{Z}(z_{n},z)\to 0$:
    $\liminf_{n\to\infty}F_{n}(z_{n})\geq F(z);$
  2. 2.
    For every $z\in Z$, there exists a sequence $\{z_{n}\}$ with $d_{Z}(z_{n},z)\to 0$ and
    $\limsup_{n\to\infty}F_{n}(z_{n})\leq F(z).$
The notion of $\Gamma$-convergence allows one to derive the convergence of minimizers from compactness.
Definition 2.5.
We say that a sequence of functionals $F_{n}:Z\to\mathbb{R}$ has the compactness property if the following holds: if $\{n_{k}\}_{k\in\mathbb{N}}$ is an increasing sequence of integers and $\{z_{k}\}_{k\in\mathbb{N}}$ is a bounded sequence in $Z$ for which $\sup_{k\in\mathbb{N}}F_{n_{k}}(z_{k})<\infty$, then the closure of $\{z_{k}\}$ has a convergent subsequence.
Proposition 2.6.
Convergence of minimizers. Let $F_{n}:Z\mapsto[0,\infty]$ be a sequence of functionals which are not identically equal to $\infty$. Suppose that the functionals satisfy the compactness property and that they $\Gamma$-converge to $F:Z\mapsto[0,\infty]$. Then
$\lim_{n\to\infty}\inf_{z\in Z}F_{n}(z)=\min_{z\in Z}F(z).$
Furthermore, the closure of every bounded sequence $\{z_{n}\}$ for which
(3) $\lim_{n\to\infty}\left(F_{n}(z_{n})-\inf_{z\in Z}F_{n}(z)\right)=0$
has a convergent subsequence and each of its cluster points is a minimizer of $F$. In particular, if $F$ has a unique minimizer, then any sequence satisfying (3) converges to the unique minimizer of $F$.
In this paper, we show that our discrete objectives $\Gamma$-converge (with respect to the $\mathrm{TL}^{p}$-topology) to the appropriate continuum objectives. Then, we will show that the sequence of discrete minimizers are precompact in $\mathrm{TL}^{p}$ and, using Proposition 2.6, deduce that the latter converge to the continuum minimizers.
3 Main results
In this section, we present our main results as well the relevant notation and assumptions used for our proofs.
3.1 General notation
For $z\in\mathbb{R}^{d}$ and $A\in\mathbb{R}^{d\times d}$, we denote by $(z)_{i}$ the $i$-th coordinate of $z$ and by $(A)_{ij}$ the $ij$-th element of $A$. We denote the Sobolev space of functions in $\mathrm{L}^{p}$ with $k$-th order derivatives as $\mathrm{W}^{k,p}$ [57].
3.2 Hypergraph setting
We will use the same probabilistic setting as in [90]. In particular, the idea is to consider a probability space with measure $\mathbb{P}$ in which elements are sequences $\{x_{i}\}_{i=1}^{\infty}$. Our results will be formulated in terms of $\mathbb{P}$, showing that certain properties holds for a set $\Psi$ of sequences $\{x_{i}\}_{i=1}^{\infty}$ with $\mathbb{P}(\Psi)=1$.
Given a set of $n$ feature vectors $\Omega_{n}=\{x_{i}\}_{i=1}^{n}\subset\Omega\subset\mathbb{R}^{d}$ where we assume that $x_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mu\in\mathcal{P}(\Omega)$, a length-scale $\varepsilon>0$ and a function $\eta:[0,\infty)\mapsto[0,\infty)$, we can define weights $w_{\varepsilon,ij}$ between vertices $x_{i}$ and $x_{j}$ as follows:
$w_{\varepsilon,ij}=\eta\left(\frac{|x_{i}-x_{j}|}{\varepsilon}\right).$
The graph $(\Omega_{n},W_{n,\varepsilon})$ where $W_{\varepsilon,n}=\{w_{\varepsilon,ij}\}_{i,j=1}^{n}$ is called a random geometric graph [68].
We now define essential matrices related to such graphs. Let $D_{n,\varepsilon}$ be the diagonal matrix with entries $d_{n,\varepsilon,ii}=\sum_{j=1}^{n}w_{\varepsilon,ij}$ and define $\sigma_{\eta}=\frac{1}{d}\int_{\mathbb{R}^{d}}\eta(|h|)|h|^{2}\,\mathrm{d}h<\infty$. The graph Laplacian is defined as
$\Delta_{n,\varepsilon}:=\frac{2}{\sigma_{\eta}n\varepsilon^{d+2}}(D_{n,\varepsilon}-W_{n,\varepsilon}).$
The latter can be interpreted as a matrix $\Delta_{n,\varepsilon}\in\mathbb{R}^{n\times n}$ or as an operator $\Delta_{n,\varepsilon}:\mathrm{L}^{2}(\mu_{n})\to\mathrm{L}^{2}(\mu_{n})$ where $\mu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$ is the empirical measure.
Given functions $u_{n},v_{n}:\Omega_{n}\to\mathbb{R}$, we also define the $\mathrm{L}^{2}(\mu_{n})$ inner product:
$\langle u_{n},v_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}=\frac{1}{n}\sum_{i=1}^{n}u_{n}(x_{i})v_{n}(x_{i}).$
Such functions can be considered vectors in $\mathbb{R}^{n}$ and we will understand $u_{n}$ as both a function $u_{n}:\Omega_{n}\to\mathbb{R}$ and a vector $\mathbb{R}^{n}$.
The eigenpairs of $\Delta_{n,\varepsilon}$ are denoted by $\{(\lambda_{n,\varepsilon,k},\psi_{n,\varepsilon,k})\}_{k=1}^{n}$ where $\lambda_{n,\varepsilon,k}$ are in increasing order, $0=\lambda_{n,\varepsilon,1}<\lambda_{n,\varepsilon,2}\leq\lambda_{n,\varepsilon,3}\leq\ldots\leq\lambda_{n,\varepsilon,n}$, (where strict inequality between $\lambda_{n,\varepsilon,1}$ and $\lambda_{n,\varepsilon,2}$ follows when the graph $G_{n,\varepsilon}$ is connected) and we note that $\{\psi_{n,\varepsilon,k}\}_{k=1}^{n}$ form for a basis of $\mathrm{L}^{2}(\mu_{n})$.
We can generalize the random geometric graph weight model to create random geometric hypergraphs. In particular, we define the weight of a hyperedge of size $k+1$ by aggregating pairwise interactions as
(4) $w_{\varepsilon,i_{0}\cdots i_{k}}=\prod_{j=1}^{k}\prod_{r=0}^{j-1}w_{\varepsilon,i_{j}i_{r}}.$
This construction biases the model toward hyperedges whose constituent nodes lie within a shared neighborhood, effectively encoding a finer notion of locality. For instance, choosing $\eta=\mathbf{1}_{[0,1]}$ yields $w_{\varepsilon,i_{0}\cdots i_{k}}>0$ if and only if the entire tuple $(x_{i_{0}},\dots,x_{i_{k}})$ lies within a common ball of radius $\varepsilon$. In this sense, $\varepsilon$ should be thought of as the length-scale of interaction between vertices. We denote by $t(k)$ the number of terms in the product $\prod_{j=1}^{k}\prod_{r=0}^{j-1}w_{\varepsilon,i_{j}i_{r}}$.
The weight construction introduced in (4) serves as the foundation for all subsequent theoretical analysis. In particular, we reformulate the hypergraph learning energy (2) using this weight model in (7). With $\eta_{\mathrm{p}}(x_{i_{0}},\dots,x_{i_{k}})=\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{i_{j}}-x_{i_{r}}|}{\varepsilon}\right)$, we define the discrete $(k,p)$-Laplacian operators which we relate to the hypergraph learning energy (7) in Proposition 3.2 as
$\Delta_{n,\varepsilon}^{(k,p)}(u)(x_{i_{0}})$ $=\frac{1}{n^{k}\varepsilon^{p+kd}}\sum_{i_{1},\dots,i_{k}=1}^{n}\biggl[\eta_{\mathrm{p}}(x_{i_{0}},\dots,x_{i_{k}})|u(x_{i_{1}})-u(x_{i_{0}})|^{p-2}(u(x_{i_{1}})-u(x_{i_{0}}))\biggr].$
We note that the $(1,2)$-Laplacian is just $\Delta_{n,\varepsilon}$.
In order to introduce the continuum counterpart of $\Delta^{(k,p)}_{n,\varepsilon}$, we first define
$\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})=\left[\prod_{s=1}^{k}\eta(|z_{s}|)\right]\left[\prod_{j=2}^{k}\prod_{r=1}^{j-1}\eta(|z_{j}-z_{r}|)\right],$
and the constants
$\sigma_{\eta}^{(k,p)}$ $=\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})|(\tilde{z}_{1})_{d}|^{p}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1},$
$\sigma_{\eta}^{(k,p,1)}$ $=\int_{(\mathbb{R}^{d})^{k}}\hskip-8.53581pt\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})|(\tilde{z}_{1})_{d}|^{p-2}(\tilde{z}_{1})_{1}^{2}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1},$
$\sigma_{\eta}^{(k,p,2)}$ $=\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})|(\tilde{z}_{1})_{d}|^{p-2}(\tilde{z}_{1})_{d}(\tilde{z}_{2})_{d}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}$
In Theorem 3.3, we will establish the precise asymptotic relationship between $\Delta^{(k,p)}_{n,\varepsilon}$ and the following operator:
$\Delta_{\infty}^{(k,p)}(u)(x)=\biggl(\|\nabla u(x)\|_{2}^{p-2}\rho(x)^{k}\nabla\rho(x)\cdot\nabla u(x)\times\frac{2(\sigma_{\eta}^{(k,p)}+(k-1)\sigma_{\eta}^{(k,p,2)})}{(p-1)\sigma_{\eta}^{(k,p,1)}}$
$+\rho(x)^{k+1}\|\nabla u(x)\|_{2}^{p-2}\biggl[\Delta u(x)+\left(\frac{\sigma_{\eta}^{(k,p)}}{\sigma_{\eta}^{(k,p,1)}}-1\right)\frac{\nabla u(x)^{\top}\nabla^{2}u(x)\nabla u(x)}{\|\nabla u(x)\|_{2}^{2}}\biggr]\biggr)\frac{\sigma_{\eta}^{(k,p,1)}(p-1)}{2\rho(x)}$
where $\Delta$ denotes the regular continuum Laplacian operator.
Remark 3.1 (Relationship between $\Delta_{\infty}^{(k,p)}$ and the $p$-Laplacian operator).
For $k=1$, we have
$\Delta_{\infty}^{(1,p)}(u)(x)=\frac{\sigma_{\eta}^{(1,p)}}{2\rho(x)}\mathrm{div}(\|\nabla u(x))\|_{2}^{p-2}\nabla u(x)\rho(x)^{2})$
so that $\Delta_{\infty}^{(1,p)}$ is a weighted $p$-Laplacian operator. The key identity used in the previous computation is $\sigma_{\eta}^{(1,p,1)}(p-1)=\sigma_{\eta}^{(1,p)}$. For $k>1$, Corollary 4.2 yields an equivalence bound of the form $\frac{\sigma_{\eta}^{(1,p,1)}(p-1)}{\sigma_{\eta}^{(1,p)}}\in[c,C]$. This implies that qualitatively, up to constants, $\Delta_{\infty}^{(k,p)}$ acts like a $p$-Laplacian operator.
3.3 Higher-order hypergraph learning
To penalize higher-order derivatives in a discrete setting, one must go beyond pairwise interactions in the energy functional (as in (2)). In this section, we introduce the Higher-Order Hypergraph Learning (HOHL) model, which is specifically designed to achieve this. In particular, we recall that $v^{\top}\Delta_{n,\varepsilon}^{s}v$, with $s\in\mathbb{R}$, corresponds to a discrete Sobolev $\mathrm{W}^{s,2}$ semi-norm applied to $v$ [25, 90].
Let $(V,E)$ be a hypergraph (independently of its weight model), and define $q=\max_{e\in E}|e|-1$ as the maximum hyperedge size minus one. For each $k\in\{1,\dots,q\}$, we define a corresponding skeleton graph $G^{(k)}=(V,E^{(k)})$ by
$E^{(k)}=\left\{\{v_{i},v_{j}\}\,\middle|\,\exists e\in E\text{ with }|e|=k+1\text{ and }\{v_{i},v_{j}\}\subset e\right\},$
that is, $G^{(k)}$ includes all pairwise edges induced by hyperedges of size $k+1$. Let $L^{(k)}$ denote the graph Laplacian associated with $G^{(k)}$. In hypergraph models where larger hyperedges connect increasingly closer points—a locality principle satisfied by our random model defined via (4)—the skeleton graphs $G^{(k)}$ become more selective as $k$ increases, capturing finer local interactions.
We define the HOHL energy as
(5) $v^{\top}\left[\sum_{k=1}^{q}\lambda_{k}(L^{(k)})^{p_{k}}\right]v$
for $v\in\mathbb{R}^{n}$, where $0<p_{1}<\ldots<p_{q}$ are powers and $\lambda_{1},\ldots,\lambda_{q}>0$ are tuning parameters. In practice, we often set $p_{k}=k$ for simplicity, although the same reasoning applies to any positive and increasing sequence $\{p_{k}\}_{k=1}^{q}$. For hypergraph models based on the locality principle detailed above, this energy imposes a hierarchical, scale-aware regularization: small $k$ enforces global smoothness, while large $k$ imposes fine-grained regularity on hyperedges of large sizes. Figure 3 illustrates this mechanism. In this paper, we restrict our attention to the HOHL energy (5) applied to the random hypergraph model where the vertex set is $\Omega_{n}\subset\mathbb{R}^{d}$ and hyperedges are constructed via (4). For a generalization of (5) to non-geometric datasets and arbitrary hypergraphs, as well as an analysis of the computational properties of HOHL, we refer the reader to [88].
Figure 3: Illustration of the HOHL energy with $p_{k}=k$. Left: For $q=2$, the energy imposes hierarchical regularization by penalizing $v^{\top}L^{(1)}v$ on skeleton edges $E^{(1)}$ and $v^{\top}(L^{(2)})^{2}v$ on $E^{(2)}$. Right: With the random hypergraph model of (4), in high-density regions, hyperedges of large size capture finer structural details, and HOHL imposes stronger smoothness to exploit this local structure.
The idea behind HOHL—imposing higher-order regularity based on sample closeness—naturally reduces, in point cloud settings where locality is geometrically defined, to a multiscale approach. In particular, our framework provides a theoretical justification for the multiscale Laplace model proposed in [61]. In the latter, the authors consider the energy
(6) $v^{\top}\left[\sum_{k=1}^{q}\lambda_{k}\Delta_{n,\varepsilon^{(k)}}^{p_{k}}\right]v,$
where $\varepsilon^{(1)}>\cdots>\varepsilon^{(q)}$, and $p_{k}>0$ controls the regularity imposed at each scale. When using the random hypergraph model, our HOHL formulation recovers this multiscale behavior via hypergraph structure and offers a principled explanation for choosing $p_{k}=k$ (or any increasing sequence): since $v^{\top}\Delta^{s}_{n,\varepsilon}v$ converges (under suitable conditions) to a $\mathrm{W}^{s,2}$ seminorm in the continuum, increasing $p_{k}$ enforces higher regularity in high-density regions where large hyperedges—and hence skeleton graphs with large kk—are more likely to form. The main distinction between (5) and (6) lies in how the Laplacians are obtained: in (6), they are constructed explicitly from multiple scaled kernels, while in (5), they are derived from the hypergraph structure.
While the hypergraph construction in (4) is theoretically appealing, it becomes computationally expensive due to the combinatorial growth in hyperedge enumeration. To address this, we propose using the multiscale Laplace model (6), which preserves the hierarchical regularization structure of HOHL while remaining computationally efficient and analytically tractable. Although (6) does not exactly approximate (5) with weights from (4)—since the corresponding limiting Laplacians may differ in density scaling—it provides a practical surrogate for point clouds embedded in a metric space, where hypergraphs are constructed based on proximity and we do not have access to the (skeleton) Laplacians. In contrast, for general hypergraphs where the Laplacians $L^{(k)}$ are known, we apply (5) directly.
3.4 Variational problems for hypergraph learning
For some $p>1$ and $k\geq 1$, the classical hypergraph energy can be written as
(7) $\mathcal{E}_{n,\varepsilon}^{(k,p)}(v)=\frac{1}{n^{k+1}\varepsilon^{p+kd}}\sum_{i_{0},\cdots,i_{k}=1}^{n}\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{i_{j}}-x_{i_{r}}|}{\varepsilon}\right)\right]\left|v(x_{i_{1}})-v(x_{i_{0}})\right|^{p}$
for $v:\Omega\to\mathbb{R}$ while the associated discrete semi-supervised learning objective is
$\mathcal{F}_{n,\varepsilon}^{(k,p)}((\nu,v))=\begin{cases}\mathcal{E}_{n,\varepsilon}^{(k,p)}(v)&\text{if }\nu=\mu_{n}\text{ and for }i\leq N,v(x_{i})=y_{i}\\ +\infty&\text{else}\end{cases}$
for $(\nu,v)\in\mathrm{TL}^{p}(\Omega)$ and where $\{y_{i}\}_{i=1}^{N}\subset\{0,1\}$ are binary labels.
For $\rho$ the density of $\mu$ with respect to Lebesgue measure, in the continuum, we define
$\mathcal{E}_{\infty}^{(k,p)}(v)$ $=\int_{\Omega}\int_{(\mathbb{R}^{d})^{k}}\left[\prod_{s=1}^{k}\eta(|z_{s}|)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(|z_{j}-z_{r}|\right)\right]\left|\nabla v(x_{0})\cdot z_{1}\right|^{p}\rho(x_{0})^{k+1}\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\mathrm{d}x_{0}$
(8) $=\int_{(\mathbb{R}^{d})^{k}}\left[\prod_{s=1}^{k}\eta(|z_{s}|)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(|z_{j}-z_{r}|\right)\right]\left|e\cdot z_{1}\right|^{p}\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\int_{\Omega}\|\nabla v(x_{0})\|_{2}^{p}\,\rho(x_{0})^{k+1}\,\mathrm{d}x_{0}$
$=:\sigma_{\eta}^{(k)}\int_{\Omega}\|\nabla v(x_{0})\|_{2}^{p}\,\rho(x_{0})^{k+1}\,\mathrm{d}x_{0}$
where $e\in\mathbb{R}^{d}$ is any vector with $\|e\|_{2}=1$ and (8) follows by isotropy of the kernels. The corresponding semi-supervised learning objectives are:
$\mathcal{F}_{\infty}^{(k,p)}((\nu,v))=\begin{cases}\mathcal{E}_{\infty}^{(k,p)}(v)&\text{if }\nu=\mu,\,v\in\mathrm{W}^{1,p}(\Omega)\text{ and for }i\leq N,v(x_{i})=y_{i}\\ +\infty&\text{else,}\end{cases}$
$\mathcal{G}_{\infty}^{(k,p)}((\nu,v))=\begin{cases}\mathcal{E}_{\infty}^{(k,p)}(v)&\text{if }\nu=\mu\text{ and }v\in\mathrm{W}^{1,p}(\Omega)\\ +\infty&\text{else.}\end{cases}$
Our final objective, for $q\geq 1$ and a positive sequence $\{\lambda_{k}\}_{k=1}^{q}\subseteq\mathbb{R}$, is to consider the sums
$(\mathcal{S}\mathcal{F})_{n,\varepsilon}^{(q,p)}((\nu,v))=\sum_{k=1}^{q}\lambda_{k}\mathcal{F}_{n,\varepsilon}^{(k,p)}((\nu,v)),$
$(\mathcal{S}\mathcal{F})_{\infty}^{(q,p)}((\nu,v))=\sum_{k=1}^{q}\lambda_{k}\mathcal{F}_{\infty}^{(k,p)}((\nu,v))$
and
$(\mathcal{S}\mathcal{G})_{\infty}^{(q,p)}((\nu,v))=\sum_{k=1}^{q}\lambda_{k}\mathcal{G}_{\infty}^{(k,p)}((\nu,v)).$
3.5 Variational problems for higher-order hypergraph learning
For HOHL and $p>0$, we define the discrete energies
$\mathcal{I}_{n,\Delta_{n,\varepsilon}}^{(p)}(v)=\langle v,\Delta_{n,\varepsilon}^{p}v\rangle_{\mathrm{L}^{2}(\mu_{n})}$
for $v:\Omega\to\mathbb{R}$ and their associated semi-supervised learning objectives
$\mathcal{J}_{n,\Delta_{n,\varepsilon}}^{(p)}((\nu,v))=\begin{cases}\mathcal{I}_{n,\Delta_{n,\varepsilon}}^{(p)}(v)&\text{if }\nu=\mu_{n}\text{ and for }i\leq N,v(x_{i})=y_{i}\\ +\infty&\text{else}\end{cases}$
for $(\nu,v)\in\mathrm{TL}^{p}(\Omega)$.
The latter have continuum analogues. Indeed, let $\Delta_{\rho}$ be the continuum weighted Laplacian operator defined by
$\Delta_{\rho}u(x)=-\frac{1}{\rho(x)}\mathrm{div}(\rho^{2}\nabla u)(x),\,x\in\Omega\quad\quad\quad\frac{\partial u}{\partial n}=0,\,x\in\partial\Omega$
and let $\{(\beta_{i},\psi_{i})\}_{i=1}^{\infty}$ be its associated eigenpairs where $\beta_{1}=0<\beta_{2}\leq\beta_{3}\leq\ldots$. We note that $\{\psi_{i}\}_{i=1}^{\infty}$ form a basis of $\mathrm{L}^{2}(\mu)$. The continuum energy is then defined as
$\mathcal{I}_{\infty}^{(p)}(v)=\langle v,\Delta^{p}_{\rho}v\rangle_{\mathrm{L}^{2}(\mu)}$
for $v:\Omega\to\mathbb{R}$ and we consider the following well-posed and ill-posed semi-supervised learning objectives:
$\mathcal{J}_{\infty}^{(p)}((\nu,v))=\begin{cases}\mathcal{I}_{\infty}^{(p)}(v)&\text{if }\nu=\mu,\,v\in\mathcal{H}^{p}(\Omega)\text{ and for }i\leq N,v(x_{i})=y_{i}\\ +\infty&\text{else,}\end{cases}$
$\mathcal{K}_{\infty}^{(p)}((\nu,v))=\begin{cases}\mathcal{I}_{\infty}^{(p)}(v)&\text{if }\nu=\mu\text{ and }v\in\mathcal{H}^{p}(\Omega)\\ +\infty&\text{else}\end{cases}$
for $(\nu,v)\in\mathrm{TL}^{p}(\Omega)$ and where
$\mathcal{H}^{p}(\Omega)=\{h\in\mathrm{L}^{2}(\mu)\,|\,\mathcal{I}_{\infty}^{(p)}(h)<+\infty\}.$
The latter set can be shown to be very closely related to the Sobolev space $\mathrm{W}^{p,2}(\Omega)$ [25, Lemma 17]. Finally, for $q\geq 1$ and positive sequences $\{\lambda_{k}\}_{k=1}^{q}\subseteq\mathbb{R}$, $P:=\{p_{k}\}_{k=1}^{q}\subseteq\mathbb{R}$ and $E:=\{\varepsilon^{(k)}\}_{k=1}^{q}$ with $\varepsilon^{(1)}>\cdots>\varepsilon^{(q)}$, we consider the sums
$(\mathcal{S}\mathcal{J})_{n,E}^{(q,P)}((\nu,v))=\sum_{k=1}^{q}\lambda_{k}\mathcal{J}_{n,\Delta_{n,\varepsilon^{(k)}}}^{(p_{k})}((\nu,v)),$
$(\mathcal{S}\mathcal{J})_{\infty}^{(q,P)}((\nu,v))=\sum_{k=1}^{q}\lambda_{k}\mathcal{J}_{\infty}^{(p_{k})}((\nu,v))$
and
$(\mathcal{S}\mathcal{K})_{\infty}^{(q,P)}((\nu,v))=\sum_{k=1}^{q}\lambda_{k}\mathcal{K}_{\infty}^{(p_{k})}((\nu,v)).$
We will also index our length-scales by the number of vertices, i.e. $\varepsilon^{(k)}=\varepsilon_{n}^{(k)}$, and in this case, we write $E_{n}:=\{\varepsilon_{n}^{(k)}\}_{k=1}^{q}$. The above sums correspond to the multiscale model for HOHL as detailed in Section 3.3.
3.6 Assumptions
In this section, we list the assumptions used throughout the paper.
Assumptions 1.
Assumption on the space. We assume either S.1 or S.2.
  1. S.1
    The feature vector space $\Omega$ is an open, connected and bounded subset of $\mathbb{R}^{d}$ with Lipschitz boundary.
  2. S.2
    The feature vector space $\Omega$ is the unit torus $\nicefrac{{\mathbb{R}^{d}}}{{\mathbb{Z}^{d}}}$.
Assumptions 2.
Assumptions on the measure. In most cases we need both M.1 and M.2.
  1. M.1
    The measure $\mu$ is a probability measure on $\Omega$.
  2. M.2
    There is a continuous Lebesgue density $\rho$ of $\mu$ which is bounded from above and below by strictly positive constants, i.e. $0<\min_{x\in\Omega}\rho(x)\leq\max_{x\in\Omega}\rho(x)<+\infty.$
The data consists of feature vectors $\{x_{i}\}_{i=1}^{n}$ and labels $\{y_{i}\}_{i=1}^{N}$ and we make the following assumptions.
Assumptions 3.
Assumptions on the data. Assumption D.1 is needed for consistency results and D.2 is needed in the semi-supervised setting.
  1. D.1
    Feature vectors $\Omega_{n}=\{x_{i}\}_{i=1}^{n}$ are iid samples from a measure $\mu$ satisfying M.1.
  2. D.2
    There are NN labels $\{y_{i}\}_{i=1}^{N}\subset\mathbb{R}$ corresponding to the first NN feature vectors $\{x_{i}\}_{i=1}^{N}$.
The weight function $\eta$ is assumed to satisfy the following assumptions.
Assumptions 4.
Assumptions on the weight function or kernel.
  1. W.1
    The function $\eta:[0,\infty)\to[0,\infty)$ is non-increasing, has compact support, is continuous and positive at $x=0$.
The compactness of the support of $\eta$ corresponds to the setting in most applications where, for computational purposes, one wants to restrict the range of interactions between vertices in our hypergraph. Theoretically however, the compact support assumption is not strictly necessary and we can extend our results to the non-compactly supported case as in done in [37, 80].
Finally, we make the following assumption on the length scale $\varepsilon=\varepsilon_{n}$ which we scale with $n$.
Assumptions 5.
Assumptions on the length-scale. For our consistency results we will need one of L.1, L.2 or L.3.
  1. L.1
    The length scale $\varepsilon=\varepsilon_{n}$ is positive, converges to 0, i.e. $0<\varepsilon_{n}\to 0$.
  2. L.2
    The length scale $\varepsilon=\varepsilon_{n}$ is positive, converges to 0, i.e. $0<\varepsilon_{n}\to 0$, and satisfies the following lower bound:
    $\begin{split}\lim_{n\to\infty}\frac{\log(n)}{n\varepsilon_{n}^{d}}&=0\qquad\text{if }d\geq 3;\\ \lim_{n\to\infty}\frac{(\log(n))^{3/2}}{n\varepsilon_{n}^{2}}&=0\qquad\text{if }d=2;\\ \lim_{n\to\infty}\frac{(\log(\log(n))}{n\varepsilon_{n}^{2}}&=0\qquad\text{if }d=1.\end{split}$
  3. L.3
    The length scale $\varepsilon=\varepsilon_{n}$ is positive, converges to 0, i.e. $0<\varepsilon_{n}\to 0$ and satisfies the following lower bound:
    $\lim_{n\to\infty}\frac{\log(n)}{n\varepsilon_{n}^{d+4}}=0.$
Assumption L.2 guarantees that (with probability one – measured with $\mathbb{P}$) that there exists $N_{1}$ such that for all $n\geq N_{1}$ the graph $G_{n,\varepsilon_{n}}=(\Omega_{n},W_{n,\varepsilon_{n}})$ is connected (see [46] or [68]). We also note that the condition in the $d=2$ case in Assumption L.2 can be tightened by removing the $\log$-term (using the techniques from [20, 17]), i.e. $\lim_{n\to\infty}\frac{\log(n)}{n\varepsilon_{n}^{d}}=0$ for $d\geq 2$, so that $\varepsilon_{n}$ can be chosen to be any sequence asymptotically greater than the connectivity radius for all $d\geq 2$.
3.7 Main results
We give our results for (classical) hypergraph learning (2) in Subsection 3.7.1, and our results for higher order hypergraph learning (6) in Subsection 3.7.2.
3.7.1 Hypergraph learning
We start by determining the corresponding Euler–Lagrange equations of (7). The result implies that the gradient of the energy decomposes into a sum of discrete operators, each tied to a hyperedges of a given size. The proof is in Section 4.2.1.
Proposition 3.2 (Discrete Euler-Lagrange equations of hypergraph learning).
The energy
$v\mapsto\sum_{k=1}^{q}\lambda_{k}\mathcal{E}_{n,\varepsilon}^{(k,p)}(v)$
is minimized by $u$ if and only if $u$ satisfies $\sum_{k=1}^{q}\lambda_{k}\Delta_{n,\varepsilon_{n}}^{(k,p)}(u)=0.$
We now study the asymptotic behavior as $n\to\infty$. The next result shows pointwise convergence of the discrete hypergraph operator to its continuum analogue.
Theorem 3.3 (Pointwise consistency).
Assume that Assumptions S.1, M.1, M.2, D.1 and L.1 hold. Furthermore, assume that $\rho\in\mathrm{C}^{2}(\Omega)$. Let $\Omega^{\prime}$ be compactly contained in $\Omega$, $q\geq 1$, $\{\lambda_{k}\}_{k=1}^{q}\subset(0,\infty)$, $p\in\{2\}\cup[3,\infty)$, $\varepsilon_{n}\leq\delta$ and $u\in\mathrm{C}^{3}$. Then, for $n$ large enough, we have that
$\biggl|\left(\sum_{k=1}^{q}\lambda_{k}\Delta_{n,\varepsilon_{n}}^{(k,p)}\right)(u)(x_{i_{0}})-\left(\sum_{k=1}^{q}\lambda_{k}\Delta_{\infty}^{(k,p)}\right)(u)(x_{i_{0}})\biggr|=\mathcal{O}\left(\delta\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
for $x_{i_{0}}\in\Omega_{n}\cap\Omega^{\prime}$, with probability $1-Cn\exp\left(-Cn\varepsilon_{n}^{2(1+qd)}\delta^{2}\right)$ where $C>0$ is a constant independent of $n$ and $\delta$.
The proof uses a Taylor expansion of $|t|^{p-2}t$ for $p\geq 3$, relying on bounded second derivatives. For $p=2$, the expansion is exact. The details are in Section 4.2.2.
For the case $k=1$, the limiting operator $\Delta_{\infty}^{(1,p)}$ is a weighted $p$-Laplacian operator. For $k>1$, we refer to Remark 3.1 where we argue that $\Delta_{\infty}^{(k,p)}$ continues to act qualitatively like a weighted $p$-Laplacian. The main conclusion of Theorem 3.3 therefore is that hypergraph learning asymptotically behaves like a weighted variant of $p$-Laplace learning in the continuum limit. Indeed, the weighted $p$-Laplacian naturally appears as the Euler–Lagrange operator of the variational problem
(9) $\min_{v:\Omega\mapsto\mathbb{R}}\int\|\nabla v(x_{0})\|_{2}^{p}\rho(x_{0})^{2}\,\mathrm{d}x_{0},$
which defines a weighted $\mathrm{W}^{1,p}$ seminorm and serves as the continuum limit of $p$-Laplace learning on graphs, see [80].
Next, we precisely characterize the asymptotic consistency of hypergraph learning as a function of the length-scale $\varepsilon_{n}$. We refer to Figure 4 for a visual summary of the result.
Figure 4: Well- and Ill-posedness characterization of hypergraph learning as a function of the length-scale $\varepsilon_{n}$.
Theorem 3.4 (Asymptotic consistency analysis of hypergraph learning).
Assume that S.1, M.1, M.2, D.1, D.2, W.1, and L.2 hold. Let $(\mu_{n},u_{n})$ be minimizers of $(\mathcal{S}\mathcal{F})_{n,\varepsilon_{n}}^{(q,p)}$.
  1. 1.
    (Well-posed case) Assume that $n\varepsilon_{n}^{p}\to 0$. Then, $\mathbb{P}$-a.s., there exists a continuous function $u$ such that $(\mu_{n},u_{n})\to(\mu,u)$ in $\mathrm{TL}^{p}(\Omega)$ and for any $\Omega^{\prime}\subset\subset\Omega$, $\max_{\{r\leq n\,|\,x_{r}\in\Omega^{\prime}\}}|u(x_{r})-u_{n}(x_{r})|\to 0$. In particular, $(\mu,u)$ is a minimizer of $(\mathcal{S}\mathcal{F})_{\infty}^{(q,p)}$.
  2. 2.
    (Ill-posed case) Assume that $n\varepsilon_{n}^{p}\to\infty$. Then, $\mathbb{P}$-a.s., there exists $u\in\mathrm{W}^{1,p}(\Omega)$ and a subsequence $\{n_{r}\}_{r=1}^{\infty}$ such that $(\mu_{n_{r}},u_{n_{r}})\to(\mu,u)$ in $\mathrm{TL}^{p}(\Omega)$ and $(\mu,u)$ is a minimizer of $(\mathcal{S}\mathcal{G})_{\infty}^{(q,p)}$.
The limiting energy identified in Theorem 3.4 is
(10) $\sum_{k=1}^{q}\lambda_{k}\sigma_{\eta}^{(k)}\int_{\Omega}\|\nabla v(x_{0})\|_{2}^{p}\,\rho(x_{0})^{k+1}\,\mathrm{d}x_{0}=\int_{\Omega}\|\nabla v(x_{0})\|_{2}^{p}\,\left(\sum_{k=1}^{q}\lambda_{k}\sigma_{\eta}^{(k)}\rho(x_{0})^{k+1}\right)\,\mathrm{d}x_{0}.$
In particular, it only differs from the limiting energy of $p$-Laplacian learning (9) by a weighting of the density. This also means that, asymptotically and similarly to our conclusion from Theorem 3.3, these hypergraph models are equivalent to graph models. We thus obtain an identical characterization of well-/ill-posedness in terms of $\varepsilon_{n}$ as in [80, Theorem 2.1], now in the hypergraph context. In particular, the well-posedness, in which minimizers smoothly interpolate the known labels, is ensured if and only if $\varepsilon_{n}$ satisfies the lower bound L.2 and the upper bound $n\varepsilon_{n}^{p}\to 0$.
Similarly to [90, Remark 3.1], the above-mentioned bounds also imply that $p>d$ and we recover an intuition stemming from Sobolev spaces. Indeed, in the continuum, our functions in $\mathrm{W}^{1,p}(\Omega)$ must be at least continuous in order to satisfy pointwise constraints, i.e. be in the well-posed case: by Sobolev inequalities, this can only the case whenever $p>d$. Our results show that this condition is necessary but not sufficient as $\varepsilon_{n}$ also has to satisfy an upper bound. We also note that in practice, the condition $p>d$ often leads to $p\geq 3$, which satisfies the requirements for pointwise convergence in Theorem 3.3.
For the ill-posed case, we note that minimizers of $(\mathcal{S}\mathcal{G})_{\infty}^{(q,p)}$ are constants and therefore, for large $n$, we expect our discrete minimizers to be almost constant with spikes at the known labels (this is observed for $q=1$ in [64, 26]). The labelling problem relying on the thresholding of our minimizers is therefore rendered nonsensical, hence our denomination of ill-posed. The case $p\leq d$ is also covered by our characterization of our ill-posed case (see [90, Remark 3.3]), linking our results back to the Sobolev embedding intuition.
We now compute the Euler–Lagrange equations associated with the limiting energy functional (10), yielding the PDE:
$\sum_{k=1}^{q}\lambda_{k}\sigma_{\eta}^{(k)}\,\frac{1}{\rho(x)}\mathrm{div}\left(\rho(x)^{k+1}\left\|\nabla v(x)\right\|^{p-2}\nabla v(x)\right)=0.$
[…]
Figure 5: Pointwise and variational convergence yield distinct limiting operators, a phenomenon not previously reported for graph-based algorithms.
3.7.2 Higher Order Hypergraph Learning
We observe from the definition of $\Delta_{\infty}^{(k,p)}$ and the limiting energy functional (10) that the notion of locality encoded by hyperedges of size $k+1$ appears through higher powers of the sampling density $\rho$. This has the beneficial effect of amplifying the contribution of high-density regions during learning. As mentioned in the introduction, this illustrates that hypergraph learning—at least in the classical formulation—modifies where pairwise interactions occur, but not how they occur. In contrast, the HOHL framework fundamentally alters the nature of interactions by introducing higher-order terms. For instance, the term $v^{\top}(L^{(2)})^{2}v$ incorporates nested finite differences of $L^{(2)}(x_{i})$ (see [85]), which already represent aggregated information from multiple neighbors. Such terms approximate second-order derivatives (see Section 3.3) and cannot be decomposed into purely pairwise interactions. In this way, HOHL better leverages the hypergraph structure by simultaneously modifying the support of interactions (via multiscale decompositions) and the mechanism of interaction (through higher-order regularization).
We obtain the following asymptotic consistency result for HOHL/ the multiscale Laplacian learning energy (6). Figure 6 provides a visual summary of the result.
Figure 6: Well- and Ill-posedness characterization of HOHL as a function of the length-scale $\varepsilon_{n}^{(q)}$. The striped regions are conjectured results.
Theorem 3.5 (Asymptotic consistency analysis of higher-order hypergraph learning).
Assume that S.2, M.1, M.2, D.1, D.2 and W.1 hold. Let $(\mu_{n},u_{n})$ be minimizers of $(\mathcal{S}\mathcal{J})_{n,\varepsilon_{n}}^{(q,p)}$. Let $q\geq 1$, $P=\{p_{k}\}_{k=1}^{q}\subseteq\mathbb{R}$ with $p_{1}\leq\cdots\leq p_{q}$ and $E_{n}=\{\varepsilon_{n}^{(k)}\}_{k=1}^{q}$ with $\varepsilon_{n}^{(1)}>\cdots>\varepsilon_{n}^{(q)}$. Assume that $\rho\in\mathrm{C}^{\infty}$.
  1. 1.
    (Well-posed case) Assume that $\varepsilon_{n}^{(q)}$ satisfies L.3, that $n\cdot(\varepsilon_{n}^{(q)})^{p_{q}/2-1/2}$ is bounded and that $p_{q}>\frac{5}{2}d+4$. Then, $\mathbb{P}$-a.s., there exists a continuous function $u$ such that $(\mu_{n},u_{n})\to(\mu,u)$ in $\mathrm{TL}^{2}(\Omega)$ and $\max_{\{r\leq n\}}|u(x_{r})-u_{n}(x_{r})|\to 0$. In particular, $(\mu,u)$ is a minimizer of $(\mathcal{S}\mathcal{J})_{\infty}^{(q,p)}$.
  2. 2.
    (Ill-posed case) Assume that $\varepsilon_{n}^{(q)}$ satisfies L.2 as well as $n(\varepsilon_{n}^{(q)})^{2p_{q}}\to\infty$. Furthermore, assume that $\sup_{n\geq 1}\|u_{n}\|_{\mathrm{L}^{2}(\mu_{n})}$ is bounded. Then, $\mathbb{P}$-a.s., there exists $u$ and a subsequence $\{n_{r}\}_{r=1}^{\infty}$ such that $(\mu_{n_{r}},u_{n_{r}})\to(\mu,u)$ in $\mathrm{TL}^{2}(\Omega)$ and $(\mu,u)$ is a minimizer of $(\mathcal{S}\mathcal{K})_{\infty}^{(q,p)}$.
Similarly to Theorem 3.4, this result shows how the choice of scale and regularity governs the transition between expressive interpolation (i.e. minimizers are smooth-functions satisfying the pointwise constraints in the well-posed case) and trivial smoothing (i.e. minimizers are constants in the ill-posed case). Theorem 3.5 identifies the precise parameter regime needed for well-posedness in semi-supervised learning with HOHL. Notably, we remark that the characterization mostly depends on the parameters of the finest scale, i.e. $\varepsilon_{n}^{(q)}$ and $p_{q}$.
In contrast to standard hypergraph learning, which converges to a $\mathrm{W}^{1,p}$ seminorm, as explained in Section 3.3, the continuum limiting enery identified through Theorem 3.5 indicates that HOHL converges to a $\mathrm{W}^{p_{q},2}$ seminorm, as illustrated in Figure 2. This underscores the distinct regularity structure induced by our higher-order formulation.
Furthermore, the same Sobolev intuition developed for hypergraph learning prevails for HOHL. In particular, our result implies that $p_{q}>d/2$—or equivalently that $\mathrm{W}^{p_{q},2}$ is embedded in $\mathrm{C}^{0}$—is necessary for well-posedness. Similarly $p_{q}\leq d/2$ also partly characterizes the ill-posed case.
4 Proofs
4.1 Limiting operator $\Delta_{\infty}^{(k,p)}$
In this section, we present the proofs leading to Remark 3.1. We denote the $d$-dimensional unit sphere by $S^{d-1}$ and parametrize it by the parameters $(r,\phi_{1},\dots,\phi_{d-1})$ in spherical coordinates. We write $\mathrm{d}\Phi$ for its surface element and let $A=[0,\pi]^{d-2}\times[0,2\pi]\subset\mathbb{R}^{d-1}$ be the range of the parameters $(\phi_{1},\dots,\phi_{d-1})$. We also write $\mathrm{Area}(S^{d-1})$ for the surface area of $S^{d-1}$. Finally, in this section, $\Gamma(x)$ denotes the classical gamma function.
Lemma 4.1 (Gamma identity).
Assume that Assumption W.1 holds and let $k>1$. Then,
$C_{\mathrm{lb}}(p_{1},\dots,p_{d})\times\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma(p_{1})\cdots\Gamma(p_{d})}{\Gamma(p_{1}+\cdots+p_{d})}$
$\leq\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})|(z_{1})_{1}|^{2p_{1}-1}\cdots|(z_{1})_{d}|^{2p_{d}-1}\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}$
$\leq C_{\mathrm{ub}}(p_{1},\dots,p_{d})\times\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma(p_{1})\cdots\Gamma(p_{d})}{\Gamma(p_{1}+\cdots+p_{d})}$
where
$C_{\mathrm{ub}}(p_{1},\dots,p_{d})=\int_{[0,\infty]^{k}}\left[\prod_{s=1}^{k}\eta(r_{s})\right]\left[\prod_{j=1}^{k}\prod_{\ell=1}^{j-1}\eta(|r_{j}-r_{\ell}|)\right]r_{1}^{2(p_{1}+\cdots+p_{d}-1/2)}r_{2}^{d-1}\cdots r_{k}^{d-1}\,\mathrm{d}r_{k}\cdots\mathrm{d}r_{1}$
and
$C_{\mathrm{lb}}(p_{1},\dots,p_{d})=\int_{[0,\infty]^{k}}\left[\prod_{s=1}^{k}\eta(r_{s})\right]\left[\prod_{j=1}^{k}\prod_{\ell=1}^{j-1}\eta(|r_{j}+r_{\ell}|)\right]r_{1}^{2(p_{1}+\cdots+p_{d}-1/2)}r_{2}^{d-1}\cdots r_{k}^{d-1}\,\mathrm{d}r_{k}\cdots\mathrm{d}r_{1}.$
For $k=1$, $C_{\mathrm{ub}}(p_{1},\dots,p_{d})=C_{\mathrm{lb}}(p_{1},\dots,p_{d})$ and therefore
$\frac{2\Gamma(p_{1})\cdots\Gamma(p_{d})}{\Gamma(p_{1}+\cdots+p_{d})}\int_{0}^{\infty}\eta(r)r^{2(p_{1}+\cdots+p_{d}-1/2)}\,\mathrm{d}r=\int_{\mathbb{R}^{d}}\eta(|z|)|(z_{1})_{1}|^{2p_{1}-1}\cdots|(z_{1})_{d}|^{2p_{d}-1}\,\mathrm{d}z_{1}.$
Proof.
Let us define
$f(p_{1},\dots,p_{d};\phi_{1},\dots,\phi_{d-1})=|\cos(\phi_{1})|^{2p_{1}-1}\prod_{s=2}^{d-1}\left|\left(\prod_{\ell=1}^{s-1}\sin(\phi_{\ell})\right)\cos(\phi_{s})\right|^{2p_{s}-1}\cdot\left|\prod_{\ell=1}^{d-1}\sin(\phi_{\ell})\right|^{2p_{d}-1}.$
We recall from [14] that
$\int_{\mathbb{R}^{d}}\eta(|x|^{2})|(x)_{1}|^{2p_{1}-1}\cdots|(x)_{d}|^{2p_{d}-1}\,\mathrm{d}x=\frac{\Gamma(p_{1})\cdots\Gamma(p_{d})}{\Gamma(p_{1}+\cdots+p_{d})}\int_{0}^{\infty}\eta(t)t^{p_{1}+\cdots+p_{d}-1}\,\mathrm{d}t.$
Using spherical coordinates, with $A=[0,\pi]^{d-2}\times[0,2\pi]$, this implies that
$\int_{\mathbb{R}^{d}}\eta(|x|^{2})|(x)_{1}|^{2p_{1}-1}\cdots|(x)_{d}|^{2p_{d}-1}\,\mathrm{d}x$
$=\int_{0}^{\infty}\eta(r^{2})r^{2(p_{1}+\cdots+p_{d}-1/2)}\,\mathrm{d}r\int_{A}f(p_{1},\dots,p_{d};\phi_{1},\dots,\phi_{d-1})\,\mathrm{d}\Phi$
$=\int_{0}^{\infty}\eta(t)t^{p_{1}+\cdots+p_{d}-1}\,\mathrm{d}t\cdot\frac{1}{2}\int_{A}f(p_{1},\dots,p_{d};\phi_{1},\dots,\phi_{d-1})\,\mathrm{d}\Phi$
from which we conclude:
(11) $\frac{\Gamma(p_{1})\cdots\Gamma(p_{d})}{\Gamma(p_{1}+\cdots+p_{d})}=\frac{1}{2}\int_{A}f(p_{1},\dots,p_{d};\phi_{1},\dots,\phi_{d-1})\,\mathrm{d}\Phi.$
For each $1\leq j\leq k$, we denote by $(r_{j},\phi_{1}^{j},\dots,\phi_{d-1}^{j})$ the spherical coordinates and by $\mathrm{d}\Phi_{j}$ the corresponding surface element. We now estimate as follows:
$\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})|(z_{1})_{1}|^{2p_{1}-1}\cdots|(z_{1})_{d}|^{2p_{d}-1}\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}$
$\leq\int_{([0,\infty]\times A)^{k}}\tilde{\eta}_{\mathrm{p}}(r_{1},\dots,r_{k})r_{1}^{2(p_{1}+\cdots+p_{d}-1/2)}$
(12) $\times r_{2}^{d-1}\cdots r_{k}^{d-1}f(p_{1},\dots,p_{d};\phi_{1}^{1},\dots,\phi_{d-1}^{1})\,\mathrm{d}r_{k}\cdots\mathrm{d}r_{1}\mathrm{d}\Phi_{1}\cdots\mathrm{d}\Phi_{k}$
(13) $=C_{\mathrm{ub}}(p_{1},\dots,p_{d})\times\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma(p_{1})\cdots\Gamma(p_{d})}{\Gamma(p_{1}+\cdots+p_{d})}$
where we used the spherical change of variables $z_{j}\mapsto(r_{j},\phi_{1}^{j},\dots,\phi_{d-1}^{j})$ for $1\leq j\leq k$, the fact that $|r_{j}-r_{\ell}|\leq|z_{j}-z_{\ell}|$ and Assumption W.1 for (12) and (11) for (13).
Using the identity, $|r_{j}+r_{\ell}|\geq|z_{j}-z_{\ell}|$, we can similarly obtain
$C_{\mathrm{lb}}(p_{1},\dots,p_{d})\times\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma(p_{1})\cdots\Gamma(p_{d})}{\Gamma(p_{1}+\cdots+p_{d})}$
$\leq\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})|(z_{1})_{1}|^{2p_{1}-1}\cdots|(z_{1})_{d}|^{2p_{d}-1}\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}.\qed$
Corollary 4.2.
Assume that Assumption W.1 holds and let $k>1$. Then,
$(p-1)\frac{\mathrm{LB}(\sigma_{\eta}^{(k,p,1)})}{\mathrm{UB}(\sigma_{\eta}^{(k,p,1)})}\leq\frac{\sigma_{\eta}^{(k,p)}}{\sigma_{\eta}^{(k,p,1)}}\leq(p-1)\frac{\mathrm{LB}(\sigma_{\eta}^{(k,p)})}{\mathrm{UB}(\sigma_{\eta}^{(k,p)})}.$
where
$\mathrm{LB}(\sigma_{\eta}^{(k,p)})=C_{\mathrm{lb}}\left(\frac{p+1}{2},\frac{1}{2},\dots,\frac{1}{2}\right)\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma\left(\frac{p+1}{2}\right)\Gamma\left(\frac{1}{2}\right)^{d-1}}{\Gamma\left(\frac{p+d}{2}\right)},$
$\mathrm{UB}(\sigma_{\eta}^{(k,p)})=C_{\mathrm{ub}}\left(\frac{p+1}{2},\frac{1}{2},\dots,\frac{1}{2}\right)\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma\left(\frac{p+1}{2}\right)\Gamma\left(\frac{1}{2}\right)^{d-1}}{\Gamma\left(\frac{p+d}{2}\right)},$
$\mathrm{LB}(\sigma_{\eta}^{(k,p,1)})=C_{\mathrm{lb}}\left(\frac{p-1}{2},\frac{1}{2},\dots,\frac{1}{2},\frac{3}{2}\right)\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma\left(\frac{p-1}{2}\right)\Gamma\left(\frac{1}{2}\right)^{d-2}\Gamma\left(\frac{3}{2}\right)}{\Gamma\left(\frac{p+d}{2}\right)}$
and
$\mathrm{UB}(\sigma_{\eta}^{(k,p,1)})=C_{\mathrm{ub}}\left(\frac{p-1}{2},\frac{1}{2},\dots,\frac{1}{2},\frac{3}{2}\right)\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma\left(\frac{p-1}{2}\right)\Gamma\left(\frac{1}{2}\right)^{d-2}\Gamma\left(\frac{3}{2}\right)}{\Gamma\left(\frac{p+d}{2}\right)}.$
For $k=1$, we have that $\mathrm{UB}(\sigma_{\eta}^{(k,p,1)})=\mathrm{LB}(\sigma_{\eta}^{(k,p,1)})$ and $\mathrm{UB}(\sigma_{\eta}^{(k,p)})=\mathrm{LB}(\sigma_{\eta}^{(k,p)})$ and therefore
$\frac{\sigma_{\eta}^{(1,p)}}{\sigma_{\eta}^{(1,p,1)}}=p-1.$
Proof.
This is just an application of Lemma 4.1. In particular,
$\mathrm{LB}(\sigma_{\eta}^{(k,p)})$ $=C_{\mathrm{lb}}\left(\frac{p+1}{2},\frac{1}{2},\dots,\frac{1}{2}\right)\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma\left(\frac{p+1}{2}\right)\Gamma\left(\frac{1}{2}\right)^{d-1}}{\Gamma\left(\frac{p+d}{2}\right)}$
$\leq\sigma_{\eta}^{(k,p)}$
$\leq C_{\mathrm{ub}}\left(\frac{p+1}{2},\frac{1}{2},\dots,\frac{1}{2}\right)\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma\left(\frac{p+1}{2}\right)\Gamma\left(\frac{1}{2}\right)^{d-1}}{\Gamma\left(\frac{p+d}{2}\right)}=\mathrm{UB}(\sigma_{\eta}^{(k,p)})$
and
$\mathrm{LB}(\sigma_{\eta}^{(k,p,1)})$ $=C_{\mathrm{lb}}\left(\frac{p-1}{2},\frac{1}{2},\dots,\frac{1}{2},\frac{3}{2}\right)\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma\left(\frac{p-1}{2}\right)\Gamma\left(\frac{1}{2}\right)^{d-2}\Gamma\left(\frac{3}{2}\right)}{\Gamma\left(\frac{p+d}{2}\right)}$
$\leq\sigma_{\eta}^{(k,p,1)}$
$\leq C_{\mathrm{ub}}\left(\frac{p-1}{2},\frac{1}{2},\dots,\frac{1}{2},\frac{3}{2}\right)\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma\left(\frac{p-1}{2}\right)\Gamma\left(\frac{1}{2}\right)^{d-2}\Gamma\left(\frac{3}{2}\right)}{\Gamma\left(\frac{p+d}{2}\right)}=\mathrm{UB}(\sigma_{\eta}^{(k,p,1)}).$
We note that $\Gamma(t+1)=t\Gamma(t)$ as well as
$C_{\mathrm{lb}}\left(\frac{p+1}{2},\frac{1}{2},\dots,\frac{1}{2}\right)=C_{\mathrm{lb}}\left(\frac{p-1}{2},\frac{1}{2},\dots,\frac{1}{2},\frac{3}{2}\right)$
and
$C_{\mathrm{ub}}\left(\frac{p+1}{2},\frac{1}{2},\dots,\frac{1}{2}\right)=C_{\mathrm{ub}}\left(\frac{p-1}{2},\frac{1}{2},\dots,\frac{1}{2},\frac{3}{2}\right).$
This allows us to conclude that $\mathrm{LB}(\sigma_{\eta}^{(k,p)})=(p-1)\mathrm{LB}(\sigma_{\eta}^{(k,p,1)})$ and $\mathrm{UB}(\sigma_{\eta}^{(k,p)})=(p-1)\mathrm{UB}(\sigma_{\eta}^{(k,p)})$. This leads to
$(p-1)\frac{\mathrm{LB}(\sigma_{\eta}^{(k,p,1)})}{\mathrm{UB}(\sigma_{\eta}^{(k,p,1)})}\leq\frac{\sigma_{\eta}^{(k,p)}}{\sigma_{\eta}^{(k,p,1)}}\leq(p-1)\frac{\mathrm{LB}(\sigma_{\eta}^{(k,p)})}{\mathrm{UB}(\sigma_{\eta}^{(k,p)})}.\qed$
4.2 Hypergraph learning
4.2.1 Euler-Lagrange equations
In this section, we present the proof for the derivation of the Euler-Lagrange equations of hypergraph learning.
Proof of Proposition 3.2.
We want to obtain the Euler-Lagrange equation of $\sum_{k=1}^{q}\lambda_{k}\mathcal{E}_{n,\varepsilon}^{(k,p)}(\cdot)$ and to this purpose, let us evaluate
$\left(\frac{d}{dt}\sum_{k=1}^{q}\lambda_{k}\mathcal{E}_{n,\varepsilon}^{(k,p)}(u+tv)\right)\lfloor_{t=0}=\sum_{k=1}^{q}\lambda_{k}\frac{d}{dt}\mathcal{E}_{n,\varepsilon}^{(k,p)}(u+tv)\lfloor_{t=0}.$
In particular, we proceed as follows:
$T$ $:=\frac{d}{dt}\mathcal{E}_{n,\varepsilon}^{(k,p)}(u+tv)\lfloor_{t=0}$
$=\frac{p}{n^{k+1}\varepsilon^{p+kd}}\sum_{i_{0},\cdots,i_{k}=1}^{n}\eta_{\mathrm{p}}(x_{i_{0}},\dots,x_{i_{k}})\left|u(x_{i_{1}})-u(x_{i_{0}})\right|^{p-2}(u(x_{i_{1}})-u(x_{i_{0}}))(v(x_{i_{1}})-v(x_{i_{0}}))$
$=:\sum_{i_{0},i_{1}=1}^{n}g(x_{i_{0}},x_{i_{1}})(v(x_{i_{1}})-v(x_{i_{0}}))$
$=\sum_{i_{0},i_{1}=1}^{n}g(x_{i_{0}},x_{i_{1}})v(x_{i_{1}})-\sum_{i_{0},i_{1}=1}^{n}g(x_{i_{0}},x_{i_{1}})v(x_{i_{0}})$
$=\sum_{i_{0},i_{1}=1}^{n}g(x_{i_{1}},x_{i_{0}})v(x_{i_{0}})-\sum_{i_{0},i_{1}=1}^{n}g(x_{i_{0}},x_{i_{1}})v(x_{i_{0}})$
(14) $=-2\sum_{i_{0},i_{1}=1}^{n}g(x_{i_{0}},x_{i_{1}})v(x_{i_{0}})$
$=\langle-2p\Delta_{n,\varepsilon}^{(k,p)}(u),v\rangle_{\mathrm{L}^{2}(\mu_{n})}$
where we used the fact that the function $f(x,y)=\eta_{\mathrm{p}}(x,y,x_{i_{2}},\dots,x_{i_{k}})$ satisfies $f(x,y)=f(y,x)$ for all fixed $x_{i_{2}},\dots,x_{i_{k}}$ implying that g(x,y)=g(y,x)g(x,y)=-g(y,x) for (14). From this, we deduce that $u$ minimizing $\sum_{k=1}^{q}\lambda_{k}\mathcal{E}_{n,\varepsilon}^{(k,p)}$ must satisfy
$\sum_{k=1}^{q}\lambda_{k}\Delta_{n,\varepsilon_{n}}^{(k,p)}(u)=0.$
Conversely, by convexity any $u$ satisfying $\sum_{k=1}^{q}\lambda_{k}\Delta_{n,\varepsilon_{n}}^{(k,p)}(u)=0$ must be a minimizer. ∎
4.2.2 Pointwise convergence
In this section, we present the proofs related to Theorem 3.3. We start with several auxiliary results. First, the following inequality [60] will be useful in the proof of Theorem 3.3.
Theorem 4.3 (McDiarmid/Azuma Inequality).
Let $X_{1},\dots,X_{n}$ be iid random variables satisfying $|X_{i}|\leq M$ almost surely. Let $Y_{n}=f(X_{1},\dots,X_{n})$ for some function $f$. If there exists $b>0$ such that $f$ satisfies
$|f(x_{1},\dots,x_{i},\dots,x_{n})-f(x_{1},\dots,\tilde{x_{i}},\dots,x_{n})|\leq b$
for all $x_{i}$ and $\tilde{x}_{i}$, $1\leq i\leq n$, then for all $t>0$,
${\mathbb{P}}(|Y_{n}-{\mathbb{E}}(Y_{n})|\geq t)\leq 2\exp\left(-\frac{t^{2}}{2nb^{2}}\right).$
Lemma 4.4.
Let $S^{(n,k)}(i)=\#\{(\alpha_{1},\dots,\alpha_{k})\in\{1,\dots,n\}^{k}\,|\,\exists 1\leq\ell\leq k\text{ such that }\alpha_{\ell}=i\}$. Then, for $1\leq i\leq n$, $S^{(n,k)}(i)=n^{k-1}+(n-1)S^{(n,k-1)}(i)$ and $S^{(n,k)}(i)=\leq kn^{k-1}$.
Proof.
Let $(\alpha_{1},\dots,\alpha_{k})\in\{1,\dots,n\}^{k}$. If we fix $\alpha_{1}=i$, then there exists $n^{k-1}$ tuples of the form $(i,\dots,\alpha_{k})$. Now, if $\alpha_{1}=j\neq i$, there exist $S^{(n,k-1)}(i)$ tuples of the form $(j,\alpha_{2},\dots,\alpha_{k})$ that contain at least one $i$. Since, $j$ can take n1n-1 values, we conclude that $S^{(n,k)}(i)=n^{k-1}+(n-1)S^{(n,k-1)}(i)$. The second claim can be proven simply by induction. ∎
The following lemma can be proven by induction and Taylor’s expansion.
Lemma 4.5 (Product identity).
Let $\rho\in\mathrm{C}^{2}(\mathbb{R}^{d})$. Then, for $k\geq 1$, we have
$\prod_{\ell=1}^{k}\rho(x_{i_{0}}+\varepsilon_{n}z_{\ell})=\rho(x_{i_{0}})^{k}+\varepsilon_{n}\rho(x_{i_{0}})^{k-1}\nabla\rho(x_{i_{0}})(z_{1}+\cdots+z_{k})+\mathcal{O}(\varepsilon_{n}^{2})$
for $x_{i_{0}},z_{1},\dots,z_{k}\in\mathbb{R}^{d}$ and $\varepsilon_{n}\in\mathbb{R}$.
We recall the following lemma from [89].
Lemma 4.6 (Asymptotics of domain of integration).
Assume that $\Omega\subset\mathbb{R}^{d}$ is a bounded open domain. Let $\varepsilon_{n}>0$ be a sequence that tends to 0, $\Omega^{\prime}$ be compactly contained in $\Omega$ and $C\subset\mathbb{R}^{d}$ be a compact subset. Then, for $n$ large enough, for all $x_{i_{0}}\in\Omega^{\prime}$, the set $S_{\varepsilon_{n}}(x_{i_{0}})=\{z\in\mathbb{R}^{d}\,|\,x_{i_{0}}+\varepsilon_{n}z\in\Omega\}\cap C$ is equal to $C$.
We recall that $f:(\mathbb{R}^{d})^{k}\mapsto\mathbb{R}$ is odd symmetric if $f(-x_{1},\dots,-x_{k})=-f(x_{1},\dots,x_{k})$ and that for such a function, $\int_{A}f(x_{1},\dots,x_{k})\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{1}=0$ if $A$ is symmetric.
Proof of Theorem 3.3.
In the proof $C>0$ will denote a constant that can be arbitrarily large, independent of $n$ and that may change from line to line. We will roughly follow the strategy in [14].
Let us start by assuming that $p\geq 3$. By Taylor’s theorem, if $\psi(t)=|t|^{p-2}t$, then, we have that $\psi(t)=\psi(a)+\psi^{\prime}(a)(t-a)+\mathcal{O}(C_{b}^{p-3}|t-a|^{2})$ for $a,t\in[-C_{b},C_{b}]$. For $u\in\mathrm{C}^{3}(\mathbb{R}^{d})$, let t=u(x+z)u(x)t=u(x+z)-u(x) and $a=\nabla u(x)\cdot z$. Then, using the previous identity, we obtain that
$\psi(u(x+z)-u(x))$
$=|\nabla u(x)\cdot z|^{p-2}\nabla u(x)\cdot z+(p-1)|\nabla u(x)\cdot z|^{p-2}(u(x+z)-u(x)-\nabla u(x)\cdot z)$
$+\mathcal{O}(C_{b}^{p-3}|u(x+z)-u(x)-\nabla u(x)\cdot z|^{2}).$
Noting that $u(x+z)-u(x)-\nabla u(x)\cdot z=z^{\top}\nabla^{2}u(x)z/2+\mathcal{O}(\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}|z|^{3})$ and also $u(x+z)-u(x)-\nabla u(x)\cdot z=\mathcal{O}(\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}|z|^{2})$, we continue the above computation:
$\psi(u(x+z)-u(x))$ $=|\nabla u(x)\cdot z|^{p-2}\nabla u(x)\cdot z+\frac{(p-1)}{2}|\nabla u(x)\cdot z|^{p-2}z^{\top}\nabla^{2}u(x)z$
$+\mathcal{O}(\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}|z|^{p+1})+\mathcal{O}(C_{b}^{p-3}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{2}|z|^{4}).$
Finally, we note that $\max\{|u(x+z)-u(x)|,|\nabla u(x)\cdot z|\}\leq C_{b}$ means that we can pick $C_{b}=\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}|z|$ which allows us to conclude that
$\psi(u(x+z)-u(x))$ $=|\nabla u(x)\cdot z|^{p-2}\nabla u(x)\cdot z+\frac{(p-1)}{2}|\nabla u(x)\cdot z|^{p-2}z^{\top}\nabla^{2}u(x)z$
(15) $+\mathcal{O}(\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}|z|^{p+1}).$
We first start by assuming that $x_{i_{0}}\in\Omega_{n}\cap\Omega^{\prime}$ is fixed (and hence non-random) and let $1\leq k\leq q$ be fixed. Let us estimate as follows:
$n^{k}\varepsilon_{n}^{p+kd}\Delta_{n,\varepsilon_{n}}^{(k,p)}(u)(x_{i_{0}})=\hskip-11.38092pt\sum_{i_{1},\dots,i_{k}=1}^{n}\eta_{\mathrm{p}}(x_{i_{0}},\dots,x_{i_{k}})|(x_{i_{1}}-x_{i_{0}})\cdot\nabla u(x_{i_{0}})|^{p-2}(x_{i_{1}}-x_{i_{0}})\cdot\nabla u(x_{i_{0}})$
(16) $+\frac{(p-1)}{2}\sum_{i_{1},\dots,i_{k}=1}^{n}\eta_{\mathrm{p}}(x_{i_{0}},\dots,x_{i_{k}})|(x_{i_{1}}-x_{i_{0}})\cdot\nabla u(x_{i_{0}})|^{p-2}(x_{i_{1}}-x_{i_{0}})^{\top}\nabla^{2}u(x_{i_{0}})(x_{i_{1}}-x_{i_{0}})$
$+\sum_{i_{1},\dots,i_{k}=1}^{n}\eta_{\mathrm{p}}(x_{i_{0}},\dots,x_{i_{k}})\mathcal{O}\left(\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}|x_{i_{1}}-x_{i_{0}}|^{p+1}\right)$
$=:T_{1}(x_{1},\dots,x_{n})+T_{2}(x_{1},\dots,x_{n})+T_{3}(x_{1},\dots,x_{n})$
$={\mathbb{E}}(T_{1})+{\mathbb{E}}(T_{2})+{\mathbb{E}}(T_{3})+\sum_{i=1}^{3}\left(T_{i}-{\mathbb{E}}(T_{i})\right)$
where we used (15) with $x=x_{i_{0}}$ and $z=x_{i_{1}}-x_{i_{0}}$ for (16).
We now want to estimate $|T_{i}-{\mathbb{E}}(T_{i})|$ for $1\leq i\leq 3$ using Theorem 4.3. For the purpose of the next few equations, for a general function $f(x_{i_{0}},\dots,x_{i_{k}})$, we will write $f(x_{i_{0}},\dots,x_{i_{k}})|_{\{x_{1},\dots,x_{n}\}}$ where the extra subscript $\{x_{1},\dots,x_{n}\}$ indicates that $x_{i_{\ell}}\in\{x_{1},\dots,x_{n}\}$ for $1\leq\ell\leq k$. Let us start by considering
$|T_{3}(x_{1},\dots,x_{i},\dots,x_{n})-T_{3}(x_{1},\dots,\tilde{x_{i}},\dots,x_{n})|$
$\leq\sum_{i_{1},\dots,i_{k}=1}^{n}\Biggl|\left[\eta_{\mathrm{p}}(x_{i_{0}},\dots,x_{i_{k}})\mathcal{O}\left(\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}|x_{i_{1}}-x_{i_{0}}|^{p+1}\right)\right]|_{\{x_{1},\dots,x_{i},\dots,x_{n}\}}$
$-\left[\eta_{\mathrm{p}}(x_{i_{0}},\dots,x_{i_{k}})\mathcal{O}\left(\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}|x_{i_{1}}-x_{i_{0}}|^{p+1}\right)\right]|_{\{x_{1},\dots,\tilde{x}_{i},\dots,x_{n}\}}\Biggr|.$
We note that each term in the latter sum is different from 0 only if there exists $1\leq\ell\leq k$ with $i_{\ell}=i$. By Lemma 4.4, there exists $S^{(n,k)}(i)\leq kn^{k-1}$ such cases and for each of those, the term in the sum can be bounded by $C\varepsilon^{p+1}\|\eta\|^{t(k)}_{\mathrm{L}^{\infty}}\|u\|^{p-1}_{\mathrm{C}^{2}(\mathbb{R}^{d})}$ where $t(k)$ is the number of terms in the double product in $\eta_{\mathrm{p}}$. By Assumptions W.1 and S.1, this leads to:
$|T_{3}(x_{1},\dots,x_{i},\dots,x_{n})-T_{3}(x_{1},\dots,\tilde{x}_{i},\dots,x_{n})|$ $\leq Cn^{k-1}\varepsilon^{p+1}\|u\|_{\mathrm{C}^{3}}^{p-1}$
Similarly,
$|T_{1}(x_{1},\dots,x_{i},\dots,x_{n})-T_{1}(x_{1},\dots,\tilde{x}_{i},\dots,x_{n})|$ $\leq Cn^{k-1}\|\eta\|_{\mathrm{L}^{\infty}}^{t(k)}\varepsilon^{p-1}\|u\|_{\mathrm{C}^{1}}^{p-1}$
and
$|T_{2}(x_{1},\dots,x_{i},\dots,x_{n})-T_{2}(x_{1},\dots,\tilde{x}_{i},\dots,x_{n})|$ $\leq Cn^{k-1}\varepsilon^{p}\|\eta\|_{\mathrm{L}^{\infty}}^{t(k)}\|u\|_{\mathrm{C}^{2}}^{p-1}.$
Using Theorem 4.3, we therefore obtain that ${\mathbb{P}}(|T_{i}-{\mathbb{E}}(T_{i})|\geq t)\leq 2\exp\left(-\frac{t^{2}}{Cn^{2k-1}\varepsilon_{n}^{2p-2}\|u\|_{\mathrm{C}^{3}}}\right)$ and with $t=n^{k}\varepsilon_{n}^{p+kd}\delta\|u\|^{p-1}_{\mathrm{C}^{3}}$,
${\mathbb{P}}(|T_{i}-{\mathbb{E}}(T_{i})|\geq n^{k}\varepsilon_{n}^{p+kd}\delta)\leq 2\exp\left(-Cn\varepsilon_{n}^{2(1+kd)}\delta^{2}\right)$
for $1\leq i\leq 3$.
We now estimate ${\mathbb{E}}(T_{i})$ for $1\leq i\leq 3$. In particular,
(17) $\frac{1}{n^{k}\varepsilon_{n}^{p+kd}}{\mathbb{E}}(T_{3})$ $=\frac{1}{n^{k}\varepsilon_{n}^{p+kd}}\mathcal{O}\left(\varepsilon_{n}^{p+1}\|u\|^{p-1}_{\mathrm{C}^{3}(\mathbb{R}^{d})}{\mathbb{E}}\left(\sum_{i_{1},\dots,i_{k}=1}^{n}\eta_{\mathrm{p}}(x_{i_{0}},\dots,x_{i_{k}})\right)\right)$
$=\mathcal{O}\left(\varepsilon_{n}\|u\|^{p-1}_{\mathrm{C}^{3}(\mathbb{R}^{d})}\frac{1}{\varepsilon_{n}^{kd}}\int_{\Omega^{k}}\eta_{\mathrm{p}}(x_{i_{0}},x_{1},\dots,x_{k})\left[\prod_{\ell=1}^{k}\rho(x_{\ell})\right]\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{1}\right)$
(18) $=\mathcal{O}\left(\varepsilon_{n}\|u\|^{p-1}_{\mathrm{C}^{3}(\mathbb{R}^{d})}\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\right)$
[…]
where we used Assumption W.1 to deduce that $|x_{i_{0}}-x_{i_{1}}|=\mathcal{O}(\varepsilon_{n})$ for (17), Assumption M.2 and the change of variables $z_{j}=(x_{j}-x_{i_{0}})/\varepsilon_{n}$ for $1\leq j\leq k$ for (18) as well as Assumption W.1 for (19).
For $T_{1}$, for $n$ large enough, we proceed as follows:
$\frac{1}{n^{k}\varepsilon_{n}^{p+kd}}{\mathbb{E}}(T_{1})$
$=\frac{1}{\varepsilon_{n}^{p+kd}}\int_{\Omega^{k}}\eta_{\mathrm{p}}(x_{i_{0}},x_{1},\dots,x_{k})$
$\times|(x_{1}-x_{i_{0}})\cdot\nabla u(x_{i_{0}})|^{p-2}(x_{1}-x_{i_{0}})\cdot\nabla u(x_{i_{0}})\left[\prod_{\ell=1}^{k}\rho(x_{\ell})\right]\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{1}$
$=\frac{1}{\varepsilon_{n}}\int_{\otimes_{j}(\{z_{j}\,|\,x_{i_{0}}+\varepsilon_{n}z_{j}\in\Omega\}\cap\mathrm{supp}(\eta))}\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})|z_{1}\cdot\nabla u(x_{i_{0}})|^{p-2}z_{1}\cdot\nabla u(x_{i_{0}})$
(20) $\times\left[\prod_{\ell=1}^{k}\rho(x_{i_{0}}+\varepsilon_{n}z_{\ell})\right]\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}$
$=\frac{1}{\varepsilon_{n}}\int_{\mathrm{supp}(\eta)^{k}}\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})|z_{1}\cdot\nabla u(x_{i_{0}})|^{p-2}z_{1}\cdot\nabla u(x_{i_{0}})\rho(x_{i_{0}})^{k}\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}$
(21) $+\int_{\mathrm{supp}(\eta)^{k}}\hskip-19.91684pt\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})|z_{1}\cdot\nabla u(x_{i_{0}})|^{p-2}z_{1}\cdot\nabla u(x_{i_{0}})\rho(x_{i_{0}})^{k-1}\nabla\rho(x_{i_{0}})(z_{1}+\dots+z_{k})\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}$
$+\mathcal{O}\left(\varepsilon_{n}\int_{\mathrm{supp}(\eta)^{k}}\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})|z_{1}\cdot\nabla u(x_{i_{0}})|^{p-2}z_{1}\cdot\nabla u(x_{i_{0}})\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\right)$
$=\int_{(\mathbb{R}^{d})^{k}}\hskip-8.5359pt\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})|z_{1}\cdot\nabla u(x_{i_{0}})|^{p-2}z_{1}\cdot\nabla u(x_{i_{0}})\rho(x_{i_{0}})^{k-1}\nabla\rho(x_{i_{0}})(z_{1}+\dots+z_{k})\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}$
(22) $+\mathcal{O}\left(\varepsilon_{n}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
$=\rho(x_{i_{0}})^{k-1}\sum_{i=1}^{d}\frac{\partial\rho}{\partial x_{i}}(x_{i_{0}})\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})\psi\left(z_{1}\cdot\nabla u(x_{i_{0}})\right)(z_{1}+\dots+z_{k})_{i}\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}$
(23) $+\mathcal{O}\left(\varepsilon_{n}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
where we used the change of variables $z_{j}=(x_{j}-x_{i_{0}})/\varepsilon_{n}$ for $1\leq j\leq k$ for (20), Lemmas 4.5 and 4.6 for (21), Assumption W.1 as well as the fact that $f(z_{1},\dots,z_{k}):=\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})|z_{1}\cdot\nabla u(x_{i_{0}})|^{p-2}z_{1}\cdot\nabla u(x_{i_{0}})\rho(x_{i_{0}})^{k}$ is odd symmetric for (22), and recalling $\psi(t)=|t|^{p-2}t$. Let $O$ be the orthogonal matrix so that $Oe_{d}=\nabla u(x_{i_{0}})/\|\nabla u(x_{i_{0}})\|_{2}$ where $e_{d}=(0,\dots,0,1)\in\mathbb{R}^{d}$. By the change of variables $\tilde{z}_{j}=O^{\top}z_{j}$ for $1\leq j\leq k$ and noting that $z_{1}\cdot\nabla u(x_{i_{0}})=\tilde{z}_{1}\cdot O^{\top}\nabla u(x_{i_{0}})=(\tilde{z}_{1})_{d}\|\nabla u(x_{i_{0}})\|_{2}$, we can continue our computation from (23):
$\frac{1}{n^{k}\varepsilon_{n}^{p+kd}}{\mathbb{E}}(T_{1})$
$=\|\nabla u(x_{i_{0}})\|_{2}^{p-1}\rho(x_{i_{0}})^{k-1}\sum_{i=1}^{d}\frac{\partial\rho}{\partial x_{i}}(x_{i_{0}})\biggl[\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})\psi\left((\tilde{z}_{1})_{d}\right)$
$\times\sum_{j=1}^{d}(O)_{ij}(\tilde{z}_{1}+\dots+\tilde{z}_{k})_{j}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}\biggr]+\mathcal{O}\left(\varepsilon_{n}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
$=\|\nabla u(x_{i_{0}})\|_{2}^{p-1}\rho(x_{i_{0}})^{k-1}\sum_{i,j=1}^{d}\frac{\partial\rho}{\partial x_{i}}(x_{i_{0}})(O)_{ij}\sum_{r=1}^{k}\int_{(\mathbb{R}^{d})^{k}}\hskip-11.38092pt\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})\psi\left((\tilde{z}_{1})_{d}\right)(\tilde{z}_{r})_{j}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}$
$+\mathcal{O}\left(\varepsilon_{n}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right).$
For $j\neq d$, we note that
$T_{4}$ $:=\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})\psi\left((\tilde{z}_{1})_{d}\right)(\tilde{z}_{r})_{j}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}$
$=\int_{(\mathbb{R})^{k(d-1)}}(\tilde{z}_{r})_{j}\left[\int_{\mathbb{R}^{k}}\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})\psi\left((\tilde{z}_{1})_{d}\right)\,\mathrm{d}(\tilde{z}_{k})_{d}\cdots\mathrm{d}(\tilde{z}_{1})_{d}\right]\,\mathrm{d}(\tilde{z}_{k})_{1:d-1}\cdots\mathrm{d}(\tilde{z}_{1})_{1:d-1}$
(denoting by $(a)_{j:k}$ the elements $a_{j},a_{j+1}\dots,a_{k-1},a_{k}$) and the function $f:\mathbb{R}^{k}\mapsto\mathbb{R}$ defined as
$f(y_{1},\dots,y_{k})=\tilde{\eta}_{\mathrm{p}}((\tilde{z}_{1})_{1:d-1},y_{1},(\tilde{z}_{2})_{1:d-1},y_{2},\dots,(\tilde{z}_{k})_{1:d-1},y_{k})\psi\left(y_{1}\right)$
is odd symmetric for any fixed $(\tilde{z}_{1})_{1:d-1},\dots,(\tilde{z}_{k})_{1:d-1}$ and therefore $T_{4}=0$ and
$\frac{1}{n^{k}\varepsilon_{n}^{p+kd}}{\mathbb{E}}(T_{1})$
$=\|\nabla u(x_{i_{0}})\|_{2}^{p-1}\rho(x_{i_{0}})^{k-1}\sum_{i=1}^{d}\frac{\partial\rho}{\partial x_{i}}(x_{i_{0}})(O)_{id}\sum_{r=1}^{k}\int_{(\mathbb{R}^{d})^{k}}\hskip-8.5359pt\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})\psi((\tilde{z}_{1})_{d})(\tilde{z}_{r})_{d}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}$
$+\mathcal{O}\left(\varepsilon_{n}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
$=\|\nabla u(x_{i_{0}})\|_{2}^{p-1}\rho(x_{i_{0}})^{k-1}\sum_{i=1}^{d}\frac{\partial\rho}{\partial x_{i}}(x_{i_{0}})(O)_{id}(\sigma_{\eta}^{(k,p)}+(k-1)\sigma_{\eta}^{(k,p,2)})+\mathcal{O}\left(\varepsilon_{n}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right).$
By recalling that $(O)_{id}=(\nabla u(x_{i_{0}}))_{i}/\|\nabla u(x_{i_{0}})\|_{2}$, we can conclude:
$\frac{1}{n^{k}\varepsilon_{n}^{p+kd}}{\mathbb{E}}(T_{1})$
(24) $=\|\nabla u(x_{i_{0}})\|_{2}^{p-2}\rho(x_{i_{0}})^{k-1}\nabla\rho(x_{i_{0}})\cdot\nabla u(x_{i_{0}})(\sigma_{\eta}^{(k,p)}+(k-1)\sigma_{\eta}^{(k,p,2)})+\mathcal{O}\left(\varepsilon_{n}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right).$
Let us now tackle $T_{2}$. We can estimate as follows, for $n$ large enough:
$\frac{1}{n^{k}\varepsilon_{n}^{p+kd}}{\mathbb{E}}(T_{2})$
(25) $=\frac{(p-1)}{2}\int_{(\mathbb{R}^{d})^{k}}\hskip-8.5359pt\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})|z_{1}\cdot\nabla u(x_{i_{0}})|^{p-2}z_{1}^{\top}\nabla^{2}u(x_{i_{0}})z_{1}\left[\prod_{\ell=1}^{k}\rho(x_{i_{0}}+\varepsilon_{n}z_{\ell})\right]\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}$
$=\frac{(p-1)}{2}\rho(x_{i_{0}})^{k}\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})|z_{1}\cdot\nabla u(x_{i_{0}})|^{p-2}z_{1}^{\top}\nabla^{2}u(x_{i_{0}})z_{1}\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}$
(26) $+\mathcal{O}\left(\varepsilon_{n}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
$=\frac{(p-1)}{2}\rho(x_{i_{0}})^{k}\|\nabla u(x_{i_{0}})\|_{2}^{p-2}\int_{(\mathbb{R}^{d})^{k}}\hskip-11.38092pt\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})|(\tilde{z}_{1})_{d}|^{p-2}(O\tilde{z}_{1})^{\top}\nabla^{2}u(x_{i_{0}})(O\tilde{z}_{1})\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}$
(27) $+\mathcal{O}\left(\varepsilon_{n}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
$=\frac{(p-1)}{2}\rho(x_{i_{0}})^{k}\|\nabla u(x_{i_{0}})\|_{2}^{p-2}\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})|(\tilde{z}_{1})_{d}|^{p-2}$
$\times\sum_{i,j=1}^{d}(\nabla^{2}u(x_{i_{0}}))_{ij}(O\tilde{z}_{1})_{i}(O\tilde{z}_{1})_{j}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}+\mathcal{O}\left(\varepsilon_{n}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
$=\frac{(p-1)}{2}\rho(x_{i_{0}})^{k}\|\nabla u(x_{i_{0}})\|_{2}^{p-2}\sum_{i,j=1}^{d}(\nabla^{2}u(x_{i_{0}}))_{ij}\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})|(\tilde{z}_{1})_{d}|^{p-2}$
$\times\sum_{r,\ell=1}^{d}(O)_{ir}(O)_{j\ell}(\tilde{z}_{1})_{\ell}(\tilde{z}_{1})_{r}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}+\mathcal{O}\left(\varepsilon_{n}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
$=\frac{(p-1)}{2}\rho(x_{i_{0}})^{k}\|\nabla u(x_{i_{0}})\|_{2}^{p-2}\sum_{i,j=1}^{d}(\nabla^{2}u(x_{i_{0}}))_{ij}\sum_{r,\ell=1}^{d}(O)_{ir}(O)_{j\ell}\int_{(\mathbb{R}^{d})^{k}}\hskip-14.22636pt\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})|(\tilde{z}_{1})_{d}|^{p-2}$
$\times(\tilde{z}_{1})_{\ell}(\tilde{z}_{1})_{r}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}+\mathcal{O}\left(\varepsilon_{n}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
where we used the change of variables $z_{j}=(x_{j}-x_{i_{0}})/\varepsilon_{n}$ for $1\leq j\leq k$ and Lemma 4.6 for (25), Lemma 4.5 and Assumption W.1 for (26) and the change of variables $\tilde{z}_{j}=O^{\top}z_{j}$ for $1\leq j\leq k$ for (27). Similarly to the above, for $\ell\neq r\neq d$,
$T_{5}$ $:=\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})|(\tilde{z}_{1})_{d}|^{p-2}(\tilde{z}_{1})_{\ell}(\tilde{z}_{1})_{r}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}$
$=\int_{(\mathbb{R})^{k(d-1)}}|(\tilde{z}_{1})_{d}|^{p-2}(\tilde{z}_{1})_{\ell}\left[\int_{\mathbb{R}^{k}}\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})(\tilde{z}_{1})_{r}\,\mathrm{d}(\tilde{z}_{k})_{r}\cdots\mathrm{d}(\tilde{z}_{1})_{r}\right]\,\mathrm{d}(\tilde{z}_{k})_{-r}\cdots\mathrm{d}(\tilde{z}_{1})_{-r}$
(denoting by $(a)_{-r}$ the vector $(a_{1},\dots,a_{r-1},a_{r+1},\dots,a_{d})$) and the function $f:\mathbb{R}^{k}\mapsto\mathbb{R}$ defined as
$f(y_{1},\dots,y_{k})=\tilde{\eta}_{\mathrm{p}}((\tilde{z}_{1})_{1:r-1},y_{1},(\tilde{z}_{1})_{r+1:d},(\tilde{z}_{2})_{1:r-1},y_{2},(\tilde{z}_{2})_{r+1:d},\dots,(\tilde{z}_{k})_{r+1:d})y_{1}$
is odd symmetric for any fixed $(\tilde{z}_{1})_{-r},\dots,(\tilde{z}_{k})_{-r}$, so $T_{5}=0$ in this case. By symmetry the case $r\neq\ell\neq d$ follows. We therefore have:
$\frac{1}{n^{k}\varepsilon_{n}^{p+kd}}{\mathbb{E}}(T_{2})$
$=\frac{(p-1)}{2}\rho(x_{i_{0}})^{k}\|\nabla u(x_{i_{0}})\|_{2}^{p-2}\sum_{i,j=1}^{d}(\nabla^{2}u(x_{i_{0}}))_{ij}\sum_{r=1}^{d}(O)_{ir}(O)_{jr}\int_{(\mathbb{R}^{d})^{k}}\hskip-11.38092pt\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})|(\tilde{z}_{1})_{d}|^{p-2}$
$\times(\tilde{z}_{1})_{r}^{2}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}+\mathcal{O}\left(\varepsilon_{n}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
$=\frac{(p-1)}{2}\rho(x_{i_{0}})^{k}\|\nabla u(x_{i_{0}})\|_{2}^{p-2}\sum_{r=1}^{d}(O^{\top}\nabla^{2}u(x_{i_{0}})O)_{rr}\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})|(\tilde{z}_{1})_{d}|^{p-2}$
$\times(\tilde{z}_{1})_{r}^{2}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}+\mathcal{O}\left(\varepsilon_{n}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
$=\frac{(p-1)}{2}\rho(x_{i_{0}})^{k}\|\nabla u(x_{i_{0}})\|_{2}^{p-2}$
$\times\biggl[\mathrm{Tr}(\nabla^{2}u(x_{i_{0}}))\sigma_{\eta}^{(k,p,1)}+(\sigma_{\eta}^{(k,p)}-\sigma_{\eta}^{(k,p,1)})(O^{\top}\nabla^{2}u(x_{i_{0}})O)_{dd}\biggr]$
$+\mathcal{O}\left(\varepsilon_{n}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
$=\frac{(p-1)}{2}\rho(x_{i_{0}})^{k}\|\nabla u(x_{i_{0}})\|_{2}^{p-2}\biggl[\mathrm{Tr}(\nabla^{2}u(x_{i_{0}}))\sigma_{\eta}^{(k,p,1)}$
(28) $+(\sigma_{\eta}^{(k,p)}-\sigma_{\eta}^{(k,p,1)})\frac{1}{\|\nabla u(x_{i_{0}})\|_{2}^{2}}\nabla u(x_{i_{0}})^{\top}\nabla^{2}u(x_{i_{0}})\nabla u(x_{i_{0}})\biggr]+\mathcal{O}\left(\varepsilon_{n}\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
where we used the fact that
$(O^{\top}\nabla^{2}u(x_{i_{0}})O)_{dd}=\sum_{i,j=1}^{n}(\nabla^{2}u(x_{i_{0}}))_{ij}(O)_{id}(O)_{jd}=\sum_{i,j=1}^{n}(\nabla^{2}u(x_{i_{0}}))_{ij}\frac{(\nabla u(x_{i_{0}}))_{i}}{\|\nabla u(x_{i_{0}})\|_{2}}\frac{(\nabla u(x_{i_{0}}))_{j}}{\|\nabla u(x_{i_{0}})\|_{2}}$
for (28).
Combining (24), (28) and (19), we obtain that, with probability $1-6\exp\left(-Cn\varepsilon_{n}^{2(1+kd)}\delta^{2}\right)$, since $\varepsilon_{n}\leq\delta$:
$\Delta_{n,\varepsilon_{n}}^{(k,p)}(u)(x_{i_{0}})$ $=\|\nabla u(x_{i_{0}})\|_{2}^{p-2}\rho(x_{i_{0}})^{k-1}\nabla\rho(x_{i_{0}})\cdot\nabla u(x_{i_{0}})(\sigma_{\eta}^{(k,p)}+(k-1)\sigma_{\eta}^{(k,p,2)})$
$+\frac{(p-1)}{2}\rho(x_{i_{0}})^{k}\|\nabla u(x_{i_{0}})\|_{2}^{p-2}\biggl[\mathrm{Tr}(\nabla^{2}u(x_{i_{0}}))\sigma_{\eta}^{(k,p,1)}$
$+(\sigma_{\eta}^{(k,p)}-\sigma_{\eta}^{(k,p,1)})\frac{1}{\|\nabla u(x_{i_{0}})\|_{2}^{2}}\nabla u(x_{i_{0}})^{\top}\nabla^{2}u(x_{i_{0}})\nabla u(x_{i_{0}})\biggr]$
$+\mathcal{O}\left(\delta\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right).$
As in [14], by taking a union bound on all $x_{i_{0}}\in\Omega_{n}\cap\Omega^{\prime}$, we obtain that
$\biggl|\Delta_{n,\varepsilon_{n}}^{(k,p)}(u)(x_{i_{0}})-\Delta_{\infty}^{(k,p)}(u)(x_{i_{0}})\biggr|\leq\mathcal{O}\left(\delta\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
with probability $1-Cn\exp\left(-Cn\varepsilon_{n}^{2(1+kd)}\delta^{2}\right)$. To conclude the proof, we sum over $1\leq k\leq q$.
When $p=2$, we have $\psi(t)=t$ and directly obtain the estimate (15):
$u(x+z)-u(x)=\nabla u(x)\cdot z+\frac{1}{2}z^{\top}\nabla^{2}u(x)z+\mathcal{O}(\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}|z|^{3}).$
For the remainder of the proof, we proceed exactly as above with $p$ replaced by $2$. ∎
4.2.3 $\Gamma$-convergence
By re-adapting the results in [37, 80], we are able to show the following $\Gamma$-convergence results. In particular, we note that we perform the following decomposition of our problem: we first show $\Gamma$-convergence of a nonlocal version of our continuum energies to their local counterparts; next, we establish $\Gamma$-convergence of the discrete energies to the nonlocal continuum energies.
We will use the following inequality often in our computations. For $a,b\in\mathbb{R}$, $\delta>0$ and $p>1$, there exists a constant $C_{\delta}$ such that
(29) $||c|^{p}-|a|^{p}|\leq C_{\delta}|c-a|^{p}+\delta|a|^{p}.$
We also note that $C_{\delta}\to\infty$ as $\delta\to 0$.
$\Gamma$-convergence of the nonlocal energies
For $v:\Omega\mapsto\mathbb{R}$ and $\varepsilon>0$ we define the nonlocal energies
$\mathcal{E}_{\varepsilon,\mathrm{NL}}^{(k,p)}(v,\eta)=\frac{1}{\varepsilon^{p+kd}}\int_{\Omega^{k+1}}\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{j}-x_{r}|}{\varepsilon}\right)\right]\left|v(x_{1})-v(x_{0})\right|^{p}\prod_{\ell=0}^{k}\rho(x_{\ell})\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{0}$
which are useful intermediary quantities when going from the discrete setting to the continuum one. In this Section, by re-adapting the results in [37], our aim is to prove the $\Gamma$-convergence of our nonlocal energies to the local ones in the continuum. We start with a few technical lemmas used in the subsequent results.
Lemma 4.7 (Integral identity).
Assume that S.1 and W.1 hold. For $k\geq 1$, we have
$\frac{1}{\varepsilon_{n}^{p+dk}}\int_{\Omega^{k+1}}\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{j}-x_{r}|}{\varepsilon_{n}}\right)\right]|x_{1}-x_{0}|^{2p}\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{0}=O(\varepsilon_{n}^{p}).$
Proof.
In the proof $C>0$ will denote a constant that can be arbitrarily large, is independent of $n$ and that may change from line to line.
By using the change of variables $z_{j}=(x_{j}-x_{0})/\varepsilon_{n}$ for $1\leq j\leq k$, we obtain that $(x_{j}-x_{r})/\varepsilon_{n}=z_{j}-z_{r}$ for $1\leq r<j\leq k$. By the latter,
$T_{1}$ $:=\frac{1}{\varepsilon_{n}^{p+dk}}\int_{\Omega^{k+1}}\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{j}-x_{r}|}{\varepsilon_{n}}\right)\right]|x_{1}-x_{0}|^{2p}\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{0}$
$=\frac{\varepsilon_{n}^{2p}}{\varepsilon_{n}^{p}}\int_{\Omega}\int_{\{z_{j}\,|\,x_{0}+\varepsilon_{n}z_{j}\in\Omega\}}|z_{1}|^{2p}\left[\prod_{s=1}^{k}\eta(|z_{s}|)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(|z_{j}-z_{r}|\right)\right]\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\mathrm{d}x_{0}$
$\leq C\varepsilon_{n}^{p}\int_{(\mathbb{R}^{d})^{k}}|z_{1}|^{2p}\left[\prod_{s=1}^{k}\eta(|z_{s}|)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(|z_{j}-z_{r}|\right)\right]\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}$
$=O(\varepsilon_{n}^{p})$
where the last equality follows from Assumption W.1. ∎
Lemma 4.8 (Product identities).
Let $\rho:\mathbb{R}^{d}\mapsto\mathbb{R}$ be a Lipschitz function that is bounded above. For $x_{0},z_{1},\cdots,z_{k}\in\mathbb{R}^{d}$ and $k\geq 1$, we have the following identities:
(30) $\left|\prod_{r=1}^{k}\rho\left(z_{r}+x_{0}\right)-\rho(x_{0})^{k}\right|\leq C(\rho)\sum_{r=1}^{k}|z_{r}|$
and
(31) $\left|\prod_{r=0}^{k}\rho\left(x_{r}+z\right)-\prod_{r=0}^{k}\rho(x_{r})\right|\leq C(\rho)\left|z\right|$
for constants $C(\rho)$ only depending on $\rho$.
Proof.
We only show how to derive (30) as the proof of (31) is similar.
We proceed by induction. For $k=1$, $|\rho(x_{0}+z_{1})-\rho(x_{0})|\leq\textrm{Lip}(\rho)|z_{1}|$. Now assume that (30) holds for k1k-1. We compute as follows:
$\left|\prod_{r=1}^{k}\rho\left(z_{r}+x_{0}\right)-\rho(x_{0})^{k}\right|$ $\leq\left|\prod_{r=1}^{k}\rho\left(z_{r}+x_{0}\right)-\rho(x_{0})\prod_{r=1}^{k-1}\rho\left(z_{r}+x_{0}\right)\right|+\left|\rho(x_{0})\prod_{r=1}^{k-1}\rho\left(z_{r}+x_{0}\right)-\rho(x_{0})^{k}\right|$
$=\left|\prod_{r=1}^{k-1}\rho\left(z_{r}+x_{0}\right)\right|\left|\rho\left(z_{k}+x_{0}\right)-\rho(x_{0})\right|$
$+|\rho(x_{0})|\left|\prod_{r=1}^{k-1}\rho\left(z_{r}+x_{0}\right)-\rho(x_{0})^{k-1}\right|$
$\leq\|\rho\|_{\mathrm{L}^{\infty}}^{k-1}\textrm{Lip}(\rho)\left|z_{k}\right|+\|\rho\|_{\mathrm{L}^{\infty}}C(\rho)\sum_{r=1}^{k-1}|z_{r}|.\qed$
Lemma 4.9 (Pointwise convergence of nonlocal energies).
Assume that S.1, M.1, M.2 and W.1 hold. Let $\{v_{\varepsilon_{n}}\}$ be a sequence of functions in $\mathrm{C}^{2}(\mathbb{R}^{d})$ such that
(32) $\sup_{n\in\mathbb{N}}\{\|\nabla v_{\varepsilon_{n}}\|_{\mathrm{L}^{\infty}(\mathbb{R}^{d})}+\|\nabla^{2}v_{\varepsilon_{n}}\|_{\mathrm{L}^{\infty}(\mathbb{R}^{d})}\}<\infty.$
Suppose that $\rho$ is a positive Lipschitz function and that $\nabla v_{\varepsilon_{n}}\to\nabla v^{*}$ in $\mathrm{L}^{p}(\Omega)$ for some $v^{*}\in\mathrm{C}^{2}(\mathbb{R}^{d})$. Then,
(33) $\lim_{n\to\infty}\mathcal{E}_{\varepsilon_{n},\mathrm{NL}}^{(k,p)}(v_{\varepsilon_{n}},\eta)=\mathcal{E}_{\infty}^{(k,p)}(v^{*}).$
Proof.
In the proof $C>0$ will denote a constant that can be arbitrarily large, is independent of $n$ and that may change from line to line.
For a function $v\in\mathrm{C}^{2}(\mathbb{R}^{d})$ and $x_{0},x_{1}\in\Omega$, we have:
$v(x_{1})-v(x_{0})$ $=\nabla v(x_{0})\cdot(x_{1}-x_{0})+(x_{1}-x_{0})^{T}\nabla^{2}v(c)(x_{1}-x_{0}).$
for some constant $c$ depending on $x_{0}$ and $x_{1}$. Now, define
$H_{\varepsilon_{n}}(v)$ $=\frac{1}{\varepsilon_{n}^{p+dk}}\int_{\Omega^{k+1}}\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{j}-x_{r}|}{\varepsilon_{n}}\right)\right]|\nabla v(x_{0})\cdot(x_{1}-x_{0})|^{p}\left[\prod_{\ell=0}^{k}\rho(x_{\ell})\right]\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{0}.$
We note that by Assumption W.1, we have $H_{\varepsilon_{n}}(v)\leq C\|\nabla v\|_{\mathrm{L}^{\infty}}$. Then, we estimate as follows for $\delta>0$:
$T_{1}:=|\mathcal{E}_{\varepsilon_{n},\mathrm{NL}}^{(k,p)}(v_{\varepsilon_{n}},\eta)-H_{\varepsilon_{n}}(v_{\varepsilon_{n}})|$
$\leq\frac{CC_{\delta}}{\varepsilon_{n}^{p+dk}}\int_{\Omega^{k+1}}\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{j}-x_{r}|}{\varepsilon_{n}}\right)\right]|v_{\varepsilon_{n}}(x_{1})-v_{\varepsilon_{n}}(x_{0})-\nabla v_{\varepsilon_{n}}(x_{0})\cdot(x_{1}-x_{0})|^{p}\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{0}$
(34) $+\delta H_{\varepsilon_{n}}(v_{\varepsilon_{n}})$
$\leq\frac{C_{\delta}\|\nabla^{2}v_{\varepsilon_{n}}\|_{\mathrm{L}^{\infty}(\mathbb{R}^{d})}}{\varepsilon_{n}^{p+dk}}\int_{\Omega^{k+1}}\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{j}-x_{r}|}{\varepsilon_{n}}\right)\right]|x_{1}-x_{0}|^{2p}\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{0}+\delta H_{\varepsilon_{n}}(v_{\varepsilon_{n}})$
(35) $=C_{\delta}O(\varepsilon_{n}^{p})+\delta H_{\varepsilon_{n}}(v_{\varepsilon_{n}})$
where we used (29) and Assumption M.2 for (34) as well as Lemma 4.7 for (35).
Next, we define
$\tilde{H}_{\varepsilon_{n}}(v)$ $=\frac{1}{\varepsilon_{n}^{p+dk}}\int_{\Omega}\int_{\begin{subarray}{c}\{z_{j}\,|\,x_{0}+z_{j}\in\Omega\}\\ \text{ for all $1\leq j\leq k$}\end{subarray}}\left[\prod_{s=1}^{k}\eta\left(\frac{|z_{s}|}{\varepsilon_{n}}\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(\frac{|z_{j}-z_{r}|}{\varepsilon_{n}}\right)\right]\left|\nabla v(x_{0})\cdot z_{1}\right|^{p}$
$\times\rho(x_{0})^{k+1}\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\mathrm{d}x_{0}.$
We note that with the change of variables $z_{j}=x_{j}-x_{0}$ for $1\leq j\leq k$, we have
$H_{\varepsilon_{n}}(v)$ $=\frac{1}{\varepsilon_{n}^{p+dk}}\int_{\Omega}\int_{\begin{subarray}{c}\{z_{j}\,|\,x_{0}+z_{j}\in\Omega\}\\ \text{ for all $1\leq j\leq k$}\end{subarray}}\left[\prod_{s=1}^{k}\eta\left(\frac{|z_{s}|}{\varepsilon_{n}}\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(\frac{|z_{j}-z_{r}|}{\varepsilon_{n}}\right)\right]\left|\nabla v(x_{0})\cdot z_{1}\right|^{p}$
$\times\rho(x_{0})\prod_{t=1}^{k}\rho\left(z_{t}+x_{0}\right)\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\mathrm{d}x_{0}.$
This leads us to
$|H_{\varepsilon_{n}}(v_{\varepsilon_{n}})-\tilde{H}_{\varepsilon_{n}}(v_{\varepsilon_{n}})|\leq\frac{C\|\nabla v_{\varepsilon_{n}}\|_{\mathrm{L}^{\infty}}\|\rho\|_{\mathrm{L}^{\infty}}}{\varepsilon_{n}^{p+dk}}\int_{\Omega}\int_{\begin{subarray}{c}\{z_{j}\,|\,x_{0}+z_{j}\in\Omega\}\\ \text{ for all $1\leq j\leq k$}\end{subarray}}|z_{1}|^{p}$
$\times\left[\prod_{s=1}^{k}\eta\left(\frac{|z_{s}|}{\varepsilon_{n}}\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(\frac{|z_{j}-z_{r}|}{\varepsilon_{n}}\right)\right]\cdot\left|\prod_{t=1}^{k}\rho\left(z_{t}+x_{0}\right)-\rho(x_{0})^{k}\right|\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\mathrm{d}x_{0}$
(36) $\leq\frac{C}{\varepsilon_{n}^{p+dk}}\int_{\Omega}\int_{\begin{subarray}{c}\{z_{j}\,|\,x_{0}+z_{j}\in\Omega\}\\ \text{ for all $1\leq j\leq k$}\end{subarray}}|z_{1}|^{p}\cdot\left[\prod_{s=1}^{k}\eta\left(\frac{|z_{s}|}{\varepsilon_{n}}\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(\frac{|z_{j}-z_{r}|}{\varepsilon_{n}}\right)\right]\cdot\sum_{r=1}^{k}|z_{r}|\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\mathrm{d}x_{0}$
(37) $\leq C\varepsilon_{n}\int_{\Omega}\int_{\begin{subarray}{c}\{\tilde{z}_{j}\,|\,\varepsilon_{n}|\tilde{z}_{j}|\leq\mathrm{diam}(\Omega)\}\\ \text{ for all $1\leq j\leq k$}\end{subarray}}|\tilde{z}_{1}|^{p}\cdot\left[\prod_{s=1}^{k}\eta\left(|\tilde{z}_{s}|\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(|\tilde{z}_{j}-\tilde{z}_{r}|\right)\right]\cdot\sum_{r=1}^{k}|\tilde{z}_{r}|\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}\mathrm{d}x_{0}$
(38) $=O(\varepsilon_{n})$
where we used (30) and (32) for (36), the change of variables $\tilde{z}_{j}=z_{j}/\varepsilon_{n}$ for (37) and Assumption W.1.
We define
$\bar{H}_{\varepsilon_{n}}(v)$ $=\frac{1}{\varepsilon_{n}^{p+dk}}\int_{\Omega}\int_{\begin{subarray}{c}\{z_{j}\,|\,x_{0}+z_{j}\notin\Omega\}\\ \text{ for any $1\leq j\leq k$}\end{subarray}}\left[\prod_{s=1}^{k}\eta\left(\frac{|z_{s}|}{\varepsilon_{n}}\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(\frac{|z_{j}-z_{r}|}{\varepsilon_{n}}\right)\right]\left|\nabla v(x_{0})\cdot z_{1}\right|^{p}$
$\times\rho(x_{0})^{k+1}\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\mathrm{d}x_{0}.$
For the latter, we have:
$\bar{H}_{\varepsilon_{n}}(v_{\varepsilon_{n}})$ $\leq\frac{C}{\varepsilon_{n}^{p+dk}}\int_{\Omega}\int_{\begin{subarray}{c}\{z_{j}\,|\,x_{0}+z_{j}\notin\Omega\}\\ \text{ for any $1\leq j\leq k$}\end{subarray}}|z_{1}|^{p}\cdot\left[\prod_{s=1}^{k}\eta\left(\frac{|z_{s}|}{\varepsilon_{n}}\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(\frac{|z_{j}-z_{r}|}{\varepsilon_{n}}\right)\right]\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\mathrm{d}x_{0}$
(39) $=C\int_{\Omega}\int_{\begin{subarray}{c}\{\tilde{z}_{j}\,|\,x_{0}+\varepsilon_{n}\tilde{z}_{j}\notin\Omega\}\\ \text{ for any $1\leq j\leq k$}\end{subarray}}|\tilde{z}_{1}|^{p}\cdot\left[\prod_{s=1}^{k}\eta\left(|\tilde{z}_{s}|\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(|\tilde{z}_{j}-\tilde{z}_{r}|\right)\right]\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}\mathrm{d}x_{0}$
$\leq C\int_{\Omega}\int_{\begin{subarray}{c}\{\tilde{z}_{j}\,|\,|\tilde{z}_{j}|\geq\frac{\mathrm{dist}(x_{0},\partial\Omega)}{\varepsilon_{n}}\}\\ \text{ for any $1\leq j\leq k$}\end{subarray}}|\tilde{z}_{1}|^{p}\cdot\left[\prod_{s=1}^{k}\eta\left(|\tilde{z}_{s}|\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(|\tilde{z}_{j}-\tilde{z}_{r}|\right)\right]\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}\mathrm{d}x_{0}$
where we used the change of variables $\tilde{z}_{j}=z_{j}/\varepsilon_{n}$ for (39). Now, by using the dominated convergence and Assumption W.1, we get that
(40) $\bar{H}_{\varepsilon_{n}}(v_{\varepsilon_{n}})=o(1).$
We continue by defining:
$\hat{H}_{\varepsilon_{n}}(v):=\bar{H}_{\varepsilon_{n}}(v)+\tilde{H}_{\varepsilon_{n}}(v)$
$=\frac{1}{\varepsilon_{n}^{p+dk}}\int_{\Omega}\int_{(\mathbb{R}^{d})^{k}}\left[\prod_{s=1}^{k}\eta\left(\frac{|z_{s}|}{\varepsilon_{n}}\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(\frac{|z_{j}-z_{r}|}{\varepsilon_{n}}\right)\right]\left|\nabla v(x_{0})\cdot z_{1}\right|^{p}\rho(x_{0})^{k+1}\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\mathrm{d}x_{0}$
(41) $=\int_{\Omega}\int_{(\mathbb{R}^{d})^{k}}\left[\prod_{s=1}^{k}\eta\left(|\tilde{z}_{s}|\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(|\tilde{z}_{j}-\tilde{z}_{r}|\right)\right]|\nabla v(x_{0})\cdot\tilde{z}_{1}|^{p}\rho(x_{0})^{k+1}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}\mathrm{d}x_{0}$
where we used the change of variables $\tilde{z}_{j}=z_{j}/\varepsilon_{n}$ for (41). We also have
$\hat{H}_{\varepsilon_{n}}(v)$ $=\frac{1}{\varepsilon_{n}^{p+dk}}\int_{\Omega}\int_{\mathbb{R}^{dk}}\left[\prod_{s=1}^{k}\eta\left(\frac{|z_{s}|}{\varepsilon_{n}}\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(\frac{|z_{j}-z_{r}|}{\varepsilon_{n}}\right)\right]\left|\nabla v(x_{0})\cdot z_{1}\right|^{p}\rho(x_{0})^{k+1}\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\mathrm{d}x_{0}$
$=\int_{\Omega}\int_{\mathbb{R}^{dk}}\left[\prod_{s=1}^{k}\eta\left(|\tilde{z}_{s}|\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(|\tilde{z}_{j}-\tilde{z}_{r}|\right)\right]\left|\nabla v(x_{0})\cdot\tilde{z}_{1}\right|^{p}\rho(x_{0})^{k+1}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}\mathrm{d}x_{0}$
$=\mathcal{E}_{\infty}^{(k,p)}(v).$
For $\delta>0$, we continue by noting that
$\left|\hat{H}(v_{\varepsilon_{n}})-\mathcal{E}_{\infty}^{(k,p)}(v^{*})\right|$
$\qquad\leq\delta\mathcal{E}_{\infty}^{(k,p)}(v^{*})+CC_{\delta}\int_{\Omega}\int_{(\mathbb{R}^{d})^{k}}|z_{1}|^{p}\cdot\left[\prod_{s=1}^{k}\eta\left(|z_{s}|\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(|z_{j}-z_{r}|\right)\right]$
(42) $\qquad\qquad\times|\nabla v_{\varepsilon_{n}}(x_{0})-\nabla v(x_{0})|^{p}\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\mathrm{d}x_{0}$
$\qquad=\delta\mathcal{E}_{\infty}^{(k,p)}(v^{*})+CC_{\delta}C(\eta)\int_{\Omega}|\nabla v_{\varepsilon_{n}}(x_{0})-\nabla v^{*}(x_{0})|^{p}\,\mathrm{d}x_{0}$
(43) $\qquad=\delta\mathcal{E}_{\infty}^{(k,p)}(v^{*})+C_{\delta}o(\varepsilon_{n})$
where $C(\eta)=\int_{(\mathbb{R}^{d})^{k}}|z_{1}|^{p}\cdot\left[\prod_{s=1}^{k}\eta\left(|z_{s}|\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(|z_{j}-z_{r}|\right)\right]\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}$ which is finite by Assumption W.1 and where we used (29) for (42) as well as the fact that $\nabla v_{\varepsilon_{n}}\to\nabla v^{*}$ in $\mathrm{L}^{p}$ for (43).
We conclude the proof by the following chain of inequalities:
$|\mathcal{E}_{\varepsilon_{n},\mathrm{NL}}^{(k,p)}(v_{\varepsilon_{n}})-\mathcal{E}_{\infty}^{(k,p)}(v^{*})|$ $\leq T_{1}+|H_{\varepsilon_{n}}(v_{\varepsilon_{n}})-\mathcal{E}_{\infty}^{(k,p)}(v^{*})|$
(44) $\leq CC_{\delta}\varepsilon_{n}^{p}+\delta H_{\varepsilon_{n}}(v_{\varepsilon_{n}})+|H_{\varepsilon_{n}}(v_{\varepsilon_{n}})-\tilde{H}_{\varepsilon_{n}}(v_{\varepsilon_{n}})|+|\tilde{H}_{\varepsilon_{n}}(v_{\varepsilon_{n}})-\mathcal{E}_{\infty}^{(k,p)}(v^{*})|$
(45) $\leq CC_{\delta}\varepsilon_{n}^{p}+\delta H_{\varepsilon_{n}}(v_{\varepsilon_{n}})+C\varepsilon_{n}+|\hat{H}_{\varepsilon_{n}}(v_{\varepsilon_{n}})-\mathcal{E}_{\infty}^{(k,p)}(v^{*})|+|\bar{H}_{\varepsilon_{n}}(v_{\varepsilon_{n}})|$
(46) $\leq CC_{\delta}\varepsilon_{n}^{p}+\delta H_{\varepsilon_{n}}(v_{\varepsilon_{n}})+C\varepsilon_{n}+\delta\mathcal{E}_{\infty}^{(k,p)}(v^{*})+C_{\delta}o(\varepsilon_{n})+o(1)$
where we used (35) for (44), (38) for (45) and (40) as well as (43) for (46). By assumption (32), we have that $H_{\varepsilon_{n}}(v_{\varepsilon_{n}})\leq C$ and therefore, by first letting $n\to\infty$ and then $\delta\to 0$, we obtain (33). ∎
Proposition 4.10 ($\liminf$-inequality for the nonlocal energies).
Assume that S.1, M.1, M.2 and W.1 hold. For every $u\in\mathrm{L}^{p}(\mu)$ and sequence $u_{\varepsilon_{n}}\to u$ in $\mathrm{L}^{p}(\mu)$, we have that:
(47) $\liminf_{n\to\infty}\mathcal{E}_{\varepsilon_{n},\mathrm{NL}}^{(k,p)}(u_{\varepsilon_{n}},\eta)\geq\mathcal{E}_{\infty}^{(k,p)}(v).$
Proof.
In the proof $C>0$ will denote a constant that can be arbitrarily large, is independent of $n$, $\delta$ and that may change from line to line.
Since (47) is trivial if $\liminf_{n\to\infty}\mathcal{E}_{\varepsilon_{n},\mathrm{NL}}^{(k,p)}(u_{\varepsilon_{n}},\eta)=\infty$, we might assume without loss of generality (see [90]) that $\sup_{n>0}\mathcal{E}_{\varepsilon_{n},\mathrm{NL}}^{(k,p)}(u_{\varepsilon_{n}},\eta)\leq C$.
We first assume that $\rho$ is Lipschitz. We will be in the same setting as in [37, Theorem 4.1] and therefore let $\Omega^{\prime}$ be compactly contained in $\Omega$. This implies that there exists $\delta^{\prime}>0$ such that $\Omega^{\prime\prime}:=\bigcup_{x\in\Omega^{\prime}}B(x,\delta^{\prime})\subset\Omega$. Furthermore, let $J$ be a positive mollifier supported in $\overline{B(0,1)}$ and for $0<\delta<\delta^{\prime}$ as well as $v\in\mathrm{L}^{p}(\mu)$ we set
$v_{\delta}(x)=\int_{\mathbb{R}^{d}}J_{\delta}(x-z)v(z)\,\mathrm{d}z.$
By [57, Theorem C.16], we have that $v_{\delta}\to v$ in $\mathrm{L}^{p}(\mu)$, $v_{\delta}$ are smooth and in particular, by Young’s convolution inequality, for $\ell\in\{1,2\}$,
(48) $\|\nabla^{\ell}v_{\delta}\|_{\mathrm{L}^{\infty}(\mathbb{R}^{d})}\leq\frac{C}{\delta^{\ell+d}}\|v\|_{\mathrm{L}^{1}(\Omega)}\leq\frac{C}{\delta^{\ell+d}}\|v\|_{\mathrm{L}^{p}(\Omega)}.$
If we therefore set $v=u_{\varepsilon_{n}}$ and $u_{\varepsilon_{n},\delta}:=(u_{\varepsilon_{n}})_{\delta}$ and insert the latter in (48), we obtain
$\sup_{\varepsilon_{n}>0}\sum_{\ell=1}^{2}\|\nabla^{\ell}u_{\varepsilon_{n},\delta}\|_{\mathrm{L}^{\infty}(\mathbb{R}^{d})}\leq C\sum_{\ell=1}^{2}\frac{1}{\delta^{\ell+d}}$
where the last inequality follows from the fact that $u_{\varepsilon_{n}}\to u$ in $\mathrm{L}^{p}(\Omega)$ implies that $\|u_{\varepsilon_{n}}\|_{\mathrm{L}^{p}(\Omega)}\leq C$ uniformly. For fixed $\delta>0$, we deduce that (32) is satisfied. Furthermore,
$\int_{\Omega^{\prime}}|\nabla u_{\varepsilon_{n},\delta}(x)-\nabla u_{\delta}(x)|^{p}\,\mathrm{d}x$ $=\frac{1}{\delta^{d+1}}\int_{\Omega^{\prime}}\left|\int_{B(0,\delta)}(\nabla J)\left(\frac{z}{\delta}\right)\left(u_{\varepsilon_{n}}(x-z)-u(x-z)\right)\,\mathrm{d}z\right|^{p}\,\mathrm{d}x$
$\leq\frac{C}{\delta^{d+1}}\int_{\Omega^{\prime}}\int_{B(0,\delta)}|u_{\varepsilon_{n}}(x-z)-u(x-z)|^{p}\,\mathrm{d}z\mathrm{d}x$
(49) $=\frac{C}{\delta^{d+1}}\int_{\Omega^{\prime}}\int_{B(x,\delta)}|u_{\varepsilon_{n}}(r)-u(r)|^{p}\,\mathrm{d}r\mathrm{d}x$
(50) $\leq\frac{C}{\delta^{d+1}}\int_{\Omega}|u_{\varepsilon_{n}}(x)-u(x)|^{p}\,\mathrm{d}x$
where we used a change of variables for (49) and the definition of $\delta^{\prime}$ as well as Assumption S.1 for (50). We conclude from the latter that $\nabla u_{\varepsilon_{n},\delta}\to\nabla u_{\delta}$ in $\mathrm{L}^{p}(\Omega^{\prime})$ as $\varepsilon_{n}\to 0$ and therefore, by Lemma 4.9,
$\lim_{n\to\infty}\frac{1}{\varepsilon_{n}^{p+kd}}\int_{(\Omega^{\prime})^{k+1}}\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{j}-x_{r}|}{\varepsilon_{n}}\right)\right]\left|u_{\varepsilon_{n},\delta}(x_{1})-u_{\varepsilon_{n},\delta}(x_{0})\right|^{p}\prod_{\ell=0}^{k}\rho(x_{\ell})\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{0}$
(51) $=\sigma_{\eta}\int_{\Omega^{\prime}}\|\nabla u_{\delta}(x_{0})\|_{2}^{p}\,\rho(x_{0})^{k+1}\,\mathrm{d}x_{0}.$
Let us define
$a_{\varepsilon_{n},\delta}$ $=\frac{1}{\varepsilon_{n}^{p+dk}}\int_{\mathbb{R}^{d}}\int_{(\Omega^{\prime\prime})^{k+1}}\frac{1}{\delta^{d}}J\left(\frac{z}{\delta}\right)\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{j}-x_{r}|}{\varepsilon}\right)\right]\left|u_{\varepsilon_{n}}(x_{1})-u_{\varepsilon_{n}}(x_{0})\right|^{p}$
$\times\left(\prod_{\ell=0}^{k}\rho(x_{\ell})-\prod_{\ell=0}^{k}\rho(x_{\ell}+z)\right)\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{0}\mathrm{d}z.$
We now estimate as follows:
$\mathcal{E}_{\varepsilon_{n},\mathrm{NL}}^{(k,p)}(u_{\varepsilon_{n}},\eta)\geq\frac{1}{\varepsilon_{n}^{p+kd}}\int_{(\Omega^{\prime\prime})^{k+1}}\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{j}-x_{r}|}{\varepsilon}\right)\right]\left|u_{\varepsilon_{n}}(x_{1})-u_{\varepsilon_{n}}(x_{0})\right|^{p}\left[\prod_{\ell=0}^{k}\rho(x_{\ell})\right]\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{0}$
$=\frac{1}{\varepsilon_{n}^{p+dk}}\int_{\mathbb{R}^{d}}\int_{(\Omega^{\prime\prime})^{k+1}}\frac{1}{\delta^{d}}J\left(\frac{z}{\delta}\right)\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{j}-x_{r}|}{\varepsilon}\right)\right]\left|u_{\varepsilon_{n}}(x_{1})-u_{\varepsilon_{n}}(x_{0})\right|^{p}$
$\qquad\qquad\times\left[\prod_{\ell=0}^{k}\rho(x_{\ell}+z)\right]\mathrm{d}x_{k}\cdots\mathrm{d}x_{0}\mathrm{d}z+a_{\varepsilon_{n},\delta}$
$\geq\frac{1}{\varepsilon_{n}^{p+dk}}\int_{\mathbb{R}^{d}}\int_{(\Omega^{\prime})^{k+1}}\frac{1}{\delta^{d}}J\left(\frac{z}{\delta}\right)\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|\hat{x}_{j}-\hat{x}_{r}|}{\varepsilon}\right)\right]\left|u_{\varepsilon_{n}}(\hat{x}_{1}-z)-u_{\varepsilon_{n}}(\hat{x}_{0}-z)\right|^{p}$
(52) $\qquad\qquad\times\left[\prod_{\ell=0}^{k}\rho(\hat{x}_{\ell})\right]\mathrm{d}\hat{x}_{k}\cdots\mathrm{d}\hat{x}_{0}\mathrm{d}z+a_{\varepsilon_{n},\delta}$
$\geq\frac{1}{\varepsilon_{n}^{p+dk}}\int_{(\Omega^{\prime})^{k+1}}\left[\prod_{\ell=0}^{k}\rho(\hat{x}_{\ell})\right]\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|\hat{x}_{j}-\hat{x}_{r}|}{\varepsilon}\right)\right]$
(53) $\qquad\qquad\times\left|\int_{\mathbb{R}^{d}}\frac{1}{\delta^{d}}J\left(\frac{z}{\delta}\right)\left(u_{\varepsilon_{n}}(\hat{x}_{1}-z)-u_{\varepsilon_{n}}(\hat{x}_{0}-z)\right)\,\mathrm{d}z\right|^{p}\mathrm{d}\hat{x}_{k}\cdots\mathrm{d}\hat{x}_{0}+a_{\varepsilon_{n},\delta}$
(54) $=\frac{1}{\varepsilon_{n}^{p+dk}}\int_{(\Omega^{\prime})^{k+1}}\left[\prod_{\ell=0}^{k}\rho(\hat{x}_{\ell})\right]\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|\hat{x}_{j}-\hat{x}_{r}|}{\varepsilon}\right)\right]\left|u_{\varepsilon_{n},\delta}(\hat{x}_{1})-u_{\varepsilon_{n},\delta}(\hat{x}_{0})\right|^{p}\mathrm{d}\hat{x}_{k}\cdots\mathrm{d}\hat{x}_{0}+a_{\varepsilon_{n},\delta}$
where we used the change of variables $\hat{x}_{j}=x_{j}+z$ for $0\leq j\leq k$ and the fact that, by definition of $\delta$, $\Omega^{\prime}\subseteq\{w\in\mathbb{R}^{d}\,|\,w+z\in\Omega^{\prime\prime}\text{ for $z\in B(0,\delta)$}\}$ for (52) as well as Jensen’s inequality with probability measure $\nu(A)=\int_{A}\frac{1}{\delta^{d}}J\left(\frac{z}{\delta}\right)\,\mathrm{d}z$ for (53).
Now, using (31), we have
$|a_{\varepsilon_{n},\delta}|$ $\leq\frac{C}{\varepsilon_{n}^{p+dk}}\int_{\mathbb{R}^{d}}\int_{(\Omega^{\prime\prime})^{k+1}}\frac{1}{\delta^{d}}J\left(\frac{z}{\delta}\right)\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{j}-x_{r}|}{\varepsilon}\right)\right]\left|u_{\varepsilon_{n}}(x_{1})-u_{\varepsilon_{n}}(x_{0})\right|^{p}|z|\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{0}\mathrm{d}z$
(55) $\leq\frac{C\delta}{\varepsilon_{n}^{p+dk}}\int_{(\Omega^{\prime\prime})^{k+1}}\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{j}-x_{r}|}{\varepsilon}\right)\right]\left|u_{\varepsilon_{n}}(x_{1})-u_{\varepsilon_{n}}(x_{0})\right|^{p}\prod_{\ell=0}^{k}\rho(x_{\ell})\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{0}$
(56) $\leq C\delta\mathcal{E}_{\varepsilon_{n},\mathrm{NL}}^{(k,p)}(u_{\varepsilon_{n}},\eta)$
where we used Assumption M.2 for (55). We therefore obtain:
$\liminf_{n\to\infty}\mathcal{E}_{\varepsilon_{n},\mathrm{NL}}^{(k,p)}(u_{\varepsilon_{n}},\eta)\geq\liminf_{\delta\to 0}\liminf_{n\to\infty}a_{\varepsilon_{n},\delta}$
(57) $+\liminf_{\delta\to 0}\liminf_{n\to\infty}\frac{1}{\varepsilon_{n}^{p+dk}}\int_{(\Omega^{\prime})^{k+1}}\left[\prod_{\ell=0}^{k}\rho(\hat{x}_{\ell})\right]\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|\hat{x}_{j}-\hat{x}_{r}|}{\varepsilon_{n}}\right)\right]\left|u_{\varepsilon_{n},\delta}(\hat{x}_{1})-u_{\varepsilon_{n},\delta}(\hat{x}_{0})\right|^{p}\mathrm{d}\hat{x}_{k}\cdots\mathrm{d}\hat{x}_{0}$
(58) $=\liminf_{\delta\to 0}\sigma_{\eta}\int_{\Omega^{\prime}}\|\nabla u_{\delta}(x_{0})\|_{2}^{p}\,\rho(x_{0})^{k+1}\,\mathrm{d}x_{0}$
(59) $\geq\sigma_{\eta}\int_{\Omega^{\prime}}\|\nabla u(x_{0})\|_{2}^{p}\,\rho(x_{0})^{k+1}\,\mathrm{d}x_{0}$
where we used (54) for (57), (56) and the fact that the energies $\mathcal{E}_{\varepsilon_{n},\mathrm{NL}}^{(k,p)}(u_{\varepsilon_{n}},\eta)$ are uniformly bounded as well as (51) for (58) and, since $u_{\delta}\to u$ in $\mathrm{L}^{p}(\Omega)$, the fact that $\mathcal{E}_{\infty}^{(k,p)}$ is lower-semicontinuous for (59). We conclude by noting that $\Omega^{\prime}$ was an arbitrary set compactly contained in $\Omega$ so we can take $\Omega^{\prime}\uparrow\Omega$ in (59) to get (47).
Dealing with the case where $\rho$ is not Lipschitz is done analogously to the proof of [37, Theorem 4.1], i.e. relying on the approximation of continuous functions by a monotone sequence of Lipschitz functions and the monotone convergence theorem to deduce the result. ∎
Proposition 4.11 ($\limsup$-inequality for the nonlocal energies).
Assume that S.1, M.1, M.2 and W.1 hold. For every $u\in\mathrm{L}^{p}(\mu)$, there exists a sequence $\{u_{\varepsilon_{n}}\}_{n=1}^{\infty}\subseteq\mathrm{L}^{p}(\mu)$ such that $u_{\varepsilon_{n}}\to u$ in $\mathrm{L}^{p}(\mu)$ and
(60) $\limsup_{n\to\infty}\mathcal{E}_{\varepsilon_{n},\mathrm{NL}}^{(k,p)}(u_{\varepsilon_{n}},\eta)\leq\mathcal{E}_{\infty}^{(k,p)}(u).$
In particular, if $u\in\mathrm{C}^{\infty}(\bar{\Omega})$ then we can choose $u_{\varepsilon_{n}}=u$.
Proof.
In the proof $C>0$ will denote a constant that can be arbitrarily large, is independent of $n$, and that may change from line to line.
We start by noting that (60) is trivial if $\mathcal{E}_{\infty}^{(k,p)}(u)=\infty$ so that we assume $u\in\mathrm{W}^{1,p}(\Omega)$. Furthermore, we are going to apply [37, Remark 2.7], so it is sufficient to verify (60) on a dense subset of $\mathrm{W}^{1,p}(\Omega)$, namely $\mathrm{C}^{\infty}_{\mathrm{c}}(\Omega)$.
We first start by assuming that $\rho$ is Lipschitz. Let us define $u_{n}=u$ and we estimate as follows:
$\mathcal{E}_{\varepsilon_{n},\mathrm{NL}}^{(k,p)}(u,\eta)\leq\frac{1}{\varepsilon_{n}^{p}}\int_{\Omega}\int_{(\mathbb{R}^{d})^{k}}\left[\prod_{s=1}^{k}\eta\left(|z_{s}|\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(|z_{j}-z_{r}|\right)\right]\left|u(x_{0}+\varepsilon_{n}z_{1})-u(x_{0})\right|^{p}$
(61) $\qquad\times\rho(x_{0})\prod_{\ell=1}^{k}\rho\left(x_{0}+\varepsilon_{n}z_{\ell}\right)\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\mathrm{d}x_{0}$
$\,\,\leq\int_{\Omega}\int_{(\mathbb{R}^{d})^{k}}\left[\prod_{s=1}^{k}\eta\left(|z_{s}|\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(|z_{j}-z_{r}|\right)\right]\left|\nabla u(x_{0})\cdot z_{1}\right|^{p}$
$\qquad\times\left[\rho(x_{0})\prod_{\ell=1}^{k}\rho\left(x_{0}+\varepsilon_{n}z_{\ell}\right)\right]\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\mathrm{d}x_{0}$
$\qquad+\varepsilon_{n}\|\rho\|_{\mathrm{L}^{\infty}}^{k+1}\|u\|_{\mathrm{C}^{2}}^{p}\int_{\Omega}\int_{(\mathbb{R}^{d})^{k}}\left[\prod_{s=1}^{k}\eta\left(|z_{s}|\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(|z_{j}-z_{r}|\right)\right]\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\mathrm{d}x_{0}$
$\,\,\leq\int_{\Omega}\int_{(\mathbb{R}^{d})^{k}}\left[\prod_{s=1}^{k}\eta\left(|z_{s}|\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(|z_{j}-z_{r}|\right)\right]\left|\nabla u(x_{0})\cdot z_{1}\right|^{p}$
$\qquad\times\left|\rho(x_{0})\prod_{\ell=1}^{k}\rho\left(x_{0}+\varepsilon_{n}z_{\ell}\right)-\rho(x_{0})^{k+1}\right|\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\mathrm{d}x_{0}+\mathcal{E}_{\infty}^{(k,p)}(u)+C\varepsilon_{n}$
$\,\,\leq C\varepsilon_{n}\int_{\Omega}\int_{(\mathbb{R}^{d})^{k}}\left[\prod_{s=1}^{k}\eta\left(|z_{s}|\right)\right]\left[\prod_{j=1}^{k}\prod_{r=1}^{j-1}\eta\left(|z_{j}-z_{r}|\right)\right]\left|\nabla u(x_{0})\cdot z_{1}\right|^{p}\cdot\sum_{\ell=1}^{k}|z_{\ell}|\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}\mathrm{d}x_{0}$
(62) $\qquad+\mathcal{E}_{\infty}^{(k,p)}(u)+C\varepsilon_{n}$
(63) $\,\,\leq\mathcal{E}_{\infty}^{(k,p)}(u)+C\varepsilon_{n}$
where we used the change of variables $z_{j}=(x_{j}-x_{0})/\varepsilon_{n}$ for $1\leq j\leq k$ for (61), Assumption W.1, (30) for (62) and Assumption W.1 for (63). Taking the limit as $\varepsilon_{n}\to 0$, we obtain (60).
In order to consider general $\rho$, we proceed as in [37, Theorem 4.1] which concludes the proof. ∎
$\Gamma$-convergence of the discrete energies
Proposition 4.12 ($\liminf$-inequality in the ill-posed case).
Assume that S.1, M.1, M.2, W.1, D.1, D.2 and L.2 hold. Then, $\mathbb{P}$-a.s., for every $(\nu,v)\in\mathrm{TL}^{p}(\Omega)$ and $\{(\nu_{n},v_{{}_{n}})\}_{n=1}^{\infty}$ with $(\nu_{n},v_{n})\to(\nu,v)$ in $\mathrm{TL}^{p}(\Omega)$, we have
(64) $\liminf_{n\to\infty}(\mathcal{S}\mathcal{F})_{n,\varepsilon_{n}}^{(q,p)}((\nu_{n},v_{n}))\geq(\mathcal{S}\mathcal{G})_{\infty}^{(q,p)}((\nu,v)).$
Proof.
With probability one, we can assume that the conclusions of Theorem 2.3 hold.
Since (64) is trivial if $\liminf_{n\to\infty}(\mathcal{S}\mathcal{F})_{n,\varepsilon_{n}}^{(q,p)}((\nu_{n},v_{n}))=\infty$, we might assume without loss of generality (see [90]) that $\sup_{n>0}(\mathcal{S}\mathcal{F})_{n,\varepsilon_{n}}^{(q,p)}((\nu_{n},v_{n}))\leq C$. This implies that $\nu_{n}=\mu_{n}$ and, since $(\nu_{n},v_{n})\to(\nu,v)$ in $\mathrm{TL}^{p}(\Omega)$, we have $\nu=\mu$. We start by showing
(65) $\liminf_{n\to\infty}\mathcal{E}_{n,\varepsilon_{n}}^{(k,p)}(v_{n})\geq\mathcal{E}_{\infty}^{(k,p)}(v).$
We follow the three-step decomposition of [37, Theorem 1.1]. First, suppose that $\eta(t)=a$ if $0\leq t\leq b$ and $\eta(t)=0$ else where $a$ and $b$ are positive constants. Define $\tilde{\varepsilon}_{n}=\varepsilon_{n}-\frac{2\|T_{n}-\mathrm{Id}\|_{\mathrm{L}^{\infty}}}{b}$. From [80, Lemma 4.2], we know that $\frac{\varepsilon_{n}}{\tilde{\varepsilon}_{n}}\to 1$ and
$\eta\left(\frac{|x-y|}{\tilde{\varepsilon}_{n}}\right)\leq\eta\left(\frac{|T_{n}(x)-T_{n}(y)|}{\varepsilon_{n}}\right).$
Using a change of variables and the above, we obtain that
$\mathcal{E}_{n,\varepsilon_{n}}^{(k,p)}(v_{n})$ $\geq\frac{1}{\varepsilon_{n}^{p+kd}}\int_{\Omega^{k+1}}\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{j}-x_{r}|}{\tilde{\varepsilon}_{n}}\right)\right]\left|v_{n}\circ T_{n}(x_{1})-v_{n}\circ T_{n}(x_{0})\right|^{p}\left[\prod_{\ell=0}^{k}\rho(x_{\ell})\right]\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{0}$
$=\left(\frac{\tilde{\varepsilon}_{n}}{\varepsilon_{n}}\right)^{p+kd}\mathcal{E}_{\tilde{\varepsilon}_{n},\mathrm{NL}}^{(k,p)}(v_{n}\circ T_{n},\eta).$
Since $u_{n}\to u$ in $\mathrm{TL}^{p}(\Omega)$, we have that $u_{n}\circ T_{n}\to u$ in $\mathrm{L}^{p}(\Omega)$ and we can therefore use Proposition 4.10 to deduce that:
$\liminf_{n\to\infty}\mathcal{E}_{n,\varepsilon_{n}}^{(k,p)}(u_{n})\geq\liminf_{n\to\infty}\left(\frac{\tilde{\varepsilon}_{n}}{\varepsilon_{n}}\right)^{p+kd}\mathcal{E}_{\tilde{\varepsilon}_{n},\mathrm{NL}}^{(k,p)}(v_{n}\circ T_{n},\eta)\geq\mathcal{E}_{\infty}^{(k,p)}(v).$
Our next step is to assume that $\eta=\sum_{k=1}^{\ell}\eta_{l}$ satisfies Assumption W.1 where $\eta_{l}$ are functions of the type considered in the above. Then, as in [37, Theorem 1.1], we use the linearity of the integral to obtain (64).
Our final step is to let $\eta$ be a general function satisfying Assumption W.1. Then, as in [37, Theorem 1.1], we use the monotone convergence theorem and approximation of $\eta$ by functions as in the previous step to obtain (65).
Now by subadditivity of the $\liminf$ we can conclude (64). ∎
Proposition 4.13 ($\limsup$-inequality in the well-posed case).
Assume that S.1, M.1, M.2, W.1, D.1, D.2 and L.2 hold. Then, $\mathbb{P}$-a.s., for every $(\nu,v)\in\mathrm{TL}^{p}(\Omega)$, there exists a sequence $\{(\nu_{n},v_{{}_{n}})\}_{n=1}^{\infty}$ with $(\nu_{n},v_{n})\to(\nu,v)$ in $\mathrm{TL}^{p}(\Omega)$ and
(66) $\limsup_{n\to\infty}\mathcal{F}_{n,\varepsilon_{n}}^{(k,p)}((\nu_{n},v_{n}))\leq\mathcal{F}_{\infty}^{(k,p)}((\nu,v)).$
In particular, for $1\leq k\leq q$ and $v\in\mathrm{C}^{\infty}(\bar{\Omega})$ with $v(x_{i})=y_{i}$ for $i\leq N$, we can pick $\{(\nu_{n},v_{n})\}_{n=1}^{\infty}=\{(\mu_{n},v|_{\Omega_{n}})\}_{n=1}^{\infty}$.
Proof.
With probability one, we can assume that the conclusions of Theorem 2.3 hold.
In the proof $C>0$ will denote a constant that can be arbitrarily large, is independent of $n$, and that may change from line to line.
We start by noting that (66) is trivial if $\mathcal{F}_{\infty}^{(k,p)}(\nu,v)=\infty$ so that we assume $\mathcal{F}_{\infty}^{(k,p)}(\nu,v)<\infty$ which implies that $\nu=\mu$, $v\in\mathrm{W}^{1,p}(\Omega)$ and $v(x_{i})=\ell_{i}$ for all $i\leq N$. Furthermore, we are going to apply [37, Remark 2.7], so it is sufficient to verify (66) on a dense subset of $\{\mu\}\times\mathrm{W}^{1,p}(\Omega)$, namely we consider $(\mu,v)\in\{\mu\}\times\mathrm{C}^{\infty}(\bar{\Omega})$ with $v(x_{i})=\ell_{i}$ for all $i\leq N$. We let $\nu_{n}=\mu_{n}$ and $v_{n}=v|_{\Omega_{n}}$ so $v_{n}(x_{i})=\ell_{i}$ for all $i\leq N$ and (66) is equivalent to
$\limsup_{n\to\infty}\mathcal{E}_{n,\varepsilon_{n}}^{(k,p)}(v_{n})\leq\mathcal{E}_{\infty}^{(k,p)}(v).$
The fact that $(\mu_{n},v_{n})\to(\mu,v)$ in $\mathrm{TL}^{p}(\Omega)$ follows analogously from what is shown in [90, Proposition 4.17]: it relies on the fact that $v$ is uniformly continuous as well as $\|T_{n}-\mathrm{Id}\|_{\mathrm{L}^{\infty}}\to 0$.
We follow the three-step decomposition of [37, Theorem 1.1]. First, suppose that $\eta(t)=a$ if $0\leq t\leq b$ and $\eta(t)=0$ else where $a$ and $b$ are positive constants. Define $\tilde{\varepsilon}_{n}=\varepsilon_{n}+\frac{2\|T_{n}-\mathrm{Id}\|_{\mathrm{L}^{\infty}}}{b}$. From [37, Theorem 1.1], we know that $\frac{\varepsilon_{n}}{\tilde{\varepsilon}_{n}}\to 1$ and similarly to the previous proposition
(67) $\mathcal{E}_{n,\varepsilon_{n}}^{(k,p)}(v_{n})\leq\left(\frac{\tilde{\varepsilon}_{n}}{\varepsilon_{n}}\right)^{p+kd}\mathcal{E}_{\tilde{\varepsilon}_{n},\mathrm{NL}}(v_{n}\circ T_{n},\eta).$
Let $\delta>0$ and estimate as follows:
$T_{1}$ $:=|\mathcal{E}_{\tilde{\varepsilon}_{n},\mathrm{NL}}^{(k,p)}(v,\eta)-\mathcal{E}_{\tilde{\varepsilon}_{n},\mathrm{NL}}^{(k,p)}(v\circ T_{n},\eta)|$
$\leq\frac{1}{\tilde{\varepsilon}_{n}^{p+kd}}\int_{\Omega^{k+1}}\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{j}-x_{r}|}{\tilde{\varepsilon}_{n}}\right)\right]\left[\prod_{\ell=0}^{k}\rho(x_{\ell})\right]$
$\qquad\qquad\times\left|\left|v(x_{1})-v(x_{0})\right|^{p}-\left|v_{n}\circ T_{n}(x_{1})-v_{n}\circ T_{n}(x_{0})\right|^{p}\right|\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{0}$
$\leq\frac{C_{\delta}}{\tilde{\varepsilon}_{n}^{p+kd}}\int_{\Omega^{k+1}}\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{j}-x_{r}|}{\tilde{\varepsilon}_{n}}\right)\right]\left[\prod_{\ell=0}^{k}\rho(x_{\ell})\right]$
(68) $\qquad\qquad\times\left|v(x_{1})-v\circ T_{n}(x_{1})-(v(x_{0})-v\circ T_{n}(x_{0}))\right|^{p}\,\mathrm{d}x_{k}\cdots\mathrm{d}x_{0}+\delta\mathcal{E}_{\tilde{\varepsilon}_{n},\mathrm{NL}}^{(k,p)}(v,\eta)$
$\leq CC_{\delta}\left(\frac{\|\mathrm{Id}-T_{n}\|_{\mathrm{L}^{\infty}}}{\tilde{\varepsilon}_{n}}\right)^{p}+\delta\mathcal{E}_{\tilde{\varepsilon}_{n},\mathrm{NL}}^{(k,p)}(v,\eta)$
where we used (29) for (68) and the fact that $v\in\mathrm{C}^{\infty}(\bar{\Omega})$ as well as Assumptions W.1 and M.2. We obtain:
(69) $\limsup_{n\to\infty}\mathcal{E}_{n,\varepsilon_{n}}^{(k,p)}(v_{n})\leq\limsup_{n\to\infty}\left(\frac{\tilde{\varepsilon}_{n}}{\varepsilon_{n}}\right)^{p+kd}\mathcal{E}_{\tilde{\varepsilon}_{n},\mathrm{NL}}^{(k,p)}(v\circ T_{n},\eta)$
(70) $\leq\limsup_{n\to\infty}\left(\frac{\tilde{\varepsilon}_{n}}{\varepsilon_{n}}\right)^{p+kd}\frac{1}{1+\delta}\left(\mathcal{E}_{\tilde{\varepsilon}_{n},\mathrm{NL}}^{(k,p)}(v,\eta)+CC_{\delta}\left(\frac{\|\mathrm{Id}-T_{n}\|_{\mathrm{L}^{\infty}}}{\tilde{\varepsilon}_{n}}\right)^{p}\right)$
(71) $\leq\frac{1}{1+\delta}\mathcal{E}_{\infty}^{(k,p)}(v)$
where we used (67) for (69), (68) for (70) and the fact that the recovery sequence in Proposition 4.11 for $v$ was $v$ for (71). Letting $\delta\to 0$ proves (66) for $\eta$ in this form.
Next we proceed as in Proposition 4.12 or [37, Theorem 1.1]: by assuming that $\eta=\sum_{k=1}^{\ell}\eta_{l}$ satisfies Assumption W.1 where $\eta_{l}$ are functions of the type considered in the above, we use the linearity of the integral to deduce (66); assuming that $\eta$ is a general function satisfying Assumption W.1, we approximate $\eta$ by functions of the type considered in the previous step and use the monotone convergence theorem to conclude. ∎
Using the result of Proposition 4.13, we can prove the next straightforward corollary. The key point to note is that, since we have the same recovery sequence for all $1\leq k\leq q$, we just apply the subadditivity of $\limsup$ to conclude.
Corollary 4.14 ($\limsup$-inequality for the sum of semi-supervised energies in the well-posed case).
Assume that S.1, M.1, M.2, W.1, D.1, D.2 and L.2 hold. Then, $\mathbb{P}$-a.s., for every $(\nu,v)\in\mathrm{TL}^{p}(\Omega)$, there exists $\{(\nu_{n},v_{{}_{n}})\}_{n=1}^{\infty}$ with $(\nu_{n},v_{n})\to(\nu,v)$ in $\mathrm{TL}^{p}(\Omega)$ such that:
$\limsup_{n\to\infty}\left(\mathcal{S}\mathcal{F}\right)_{n,\varepsilon_{n}}^{(k,p)}((\nu_{n},v_{n}))\leq\left(\mathcal{S}\mathcal{F}\right)_{\infty}^{(k,p)}((\nu,v)).$
The following proofs use arguments from [80]. For Proposition 4.15, the sum of all energies $\{\mathcal{F}_{n,\varepsilon_{n}}^{(k,p)}\}_{k=1}^{q}$ has to be considered directly as we plan to use the uniform convergence results for the $k=1$ case from [80, Lemma 4.5]. In Proposition 4.16, we show that $n\varepsilon_{n}^{p}\to\infty$ is the common lower bound for all energies $\{\mathcal{F}_{n,\varepsilon_{n}}^{(k,p)}\}_{k=1}^{q}$ in order for them to converge to the ill-posed continuum objective functions.
Proposition 4.15 ($\liminf$-inequality for the sum of semi-supervised energies in the well-posed case).
Assume that S.1, M.1, M.2, W.1, D.1, D.2 and L.2 hold. Assume that $n\varepsilon_{n}^{p}\to 0$. Then, $\mathbb{P}$-a.s., for every $(\nu,v)\in\mathrm{TL}^{p}(\Omega)$ and $\{(\nu_{n},v_{{}_{n}})\}_{n=1}^{\infty}$ with $(\nu_{n},v_{n})\to(\nu,v)$ in $\mathrm{TL}^{p}(\Omega)$, we have:
(72) $\liminf_{n\to\infty}\left(\mathcal{S}\mathcal{F}\right)_{n,\varepsilon_{n}}^{(q,p)}((\nu_{n},v_{n}))\geq\left(\mathcal{S}\mathcal{F}\right)_{\infty}^{(q,p)}((\nu,v)).$
Proof.
With probability one, we can assume that the conclusions of Proposition 4.12 and [80, Lemma 4.5] hold.
In the proof $C>0$ will denote a constant that can be arbitrarily large, is independent of $n$, and that may change from line to line.
First, by the same argument as in Proposition 4.12, we can assume that $\sup_{n\geq 1}\left(\mathcal{S}\mathcal{F}\right)_{n,\varepsilon_{n}}^{(k,p)}((\nu_{n},v_{n}))\leq C$ and therefore $\nu_{n}=\mu_{n}$ and $\nu=\mu$. In particular, we also have that $\mathcal{E}_{n,\varepsilon_{n}}^{(1,p)}(v_{n})\leq C$ and, by [80, Lemma 4.5], we deduce the existence of a continuous function $\hat{v}$ such that for any $\Omega^{\prime}\subset\subset\Omega$, $\max_{\{i\leq n_{k}\,|\,x_{i}\in\Omega^{\prime}\}}|v_{n_{k}}(x_{i})-\hat{v}(x_{i})|\to 0$: this implies that $\hat{v}(x_{i})=\ell_{i}$ for all $i\leq N$ with probability one. We also note that $v=\hat{v}$ (in particular, $v(x_{i})=\ell_{i}$ for all $i\leq N$) and (72) reduces to proving
$\liminf_{n\to\infty}\sum_{k=1}^{q}\lambda_{k}\mathcal{E}_{n,\varepsilon_{n}}^{(k,p)}(v_{n})\geq\sum_{k=1}^{q}\lambda_{k}\mathcal{E}_{\infty}^{(k,p)}(v)=\sum_{k=1}^{q}\lambda_{k}\mathcal{G}_{\infty}^{(k,p)}((\mu,v)).$
By Proposition 4.12, we know that $\liminf_{n\to\infty}\mathcal{E}_{n,\varepsilon_{n}}^{(k,p)}(v_{n})\geq\mathcal{G}_{\infty}((\mu,v))$ so that:
$\liminf_{n\to\infty}\sum_{k=1}^{q}\lambda_{k}\mathcal{E}_{n,\varepsilon_{n}}^{(k,p)}(v_{n})$ $\geq\sum_{k=1}^{q}\lambda_{k}\liminf_{n\to\infty}\mathcal{E}_{n,\varepsilon_{n}}^{(k,p)}(v_{n})$
$\geq\sum_{k=1}^{q}\lambda_{k}\mathcal{G}_{\infty}^{(k,p)}((\mu,v)).\qed$
Proposition 4.16 ($\limsup$-inequality in the ill-posed case).
Assume that S.1, M.1, M.2, W.1, D.1, D.2 and L.2 hold. Assume that $n\varepsilon_{n}^{p}\to\infty$. Then, $\mathbb{P}$-a.s., for every $(\nu,v)\in\mathrm{TL}^{p}(\Omega)$, there exists $\{(\nu_{n},v_{{}_{n}})\}_{n=1}^{\infty}$ with $(\nu_{n},v_{n})\to(\nu,v)$ in $\mathrm{TL}^{p}(\Omega)$ such that:
(73) $\limsup_{n\to\infty}\mathcal{F}_{n,\varepsilon_{n}}^{(k,p)}((\nu_{n},v_{n}))\leq\mathcal{G}_{\infty}^{(k,p)}((\nu,v)).$
In particular, for $1\leq k\leq q$ and $v\in\mathrm{C}^{\infty}(\bar{\Omega})$, we can pick $\{(\nu_{n},v_{n})\}_{n=1}^{\infty}=\{(\mu_{n},\hat{v}_{n})\}_{n=1}^{\infty}$ where $\hat{v}_{n}(x_{i})=y_{i}$ for $i\leq N$ and $\hat{v}_{n}=v|_{\Omega_{n}}$ else.
Proof.
With probability one, we can assume that the conclusions of Proposition 4.13 hold.
In the proof $C>0$ will denote a constant that can be arbitrarily large, is independent of $n$, and that may change from line to line.
We start by noting that (73) is trivial if $\mathcal{G}_{\infty}^{(k,p)}((\nu,v))=\infty$ so that we assume $\mathcal{G}_{\infty}^{(k,p)}((\nu,v))<\infty$ which implies that $\nu=\mu$ and $v\in\mathrm{W}^{1,p}(\Omega)$. Furthermore, we are going to apply [37, Remark 2.7], so it is sufficient to verify (73) on a dense subset of $\{\mu\}\times\mathrm{W}^{1,p}(\Omega)$, namely we consider $(\mu,v)\in\{\mu\}\times\mathrm{C}^{\infty}(\bar{\Omega})$. We let $\nu_{n}=\mu_{n}$ and $\hat{v}_{n}=v|_{\Omega_{n}}$.
By repeating the proof of Proposition 4.13, we can show that $(\nu_{n},\hat{v}_{n})\to(\nu,v)$ in $\mathrm{TL}^{p}(\Omega)$ and
(74) $\limsup_{n\to\infty}\mathcal{E}_{n,\varepsilon_{n}}^{(k,p)}(\hat{v}_{n})\leq\mathcal{G}_{\infty}^{(k,p)}((\mu,v)).$
The subtlety of (74) compared to (66) is that $\hat{v}_{n}$ does not necessarily satisfy $\hat{v}_{n}(x_{i})=\ell_{i}$ for all $i\leq N$ since this condition is not imposed on $v$. We note that $\|\hat{v}_{n}\|_{\mathrm{L}^{\infty}}\leq C$ since $\hat{v}_{n}=v|_{\Omega_{n}}$ and $v\in\mathrm{C}^{\infty}(\bar{\Omega})$.
Define $(\mu_{n},v_{n})$ with
$v_{n}(x_{i})=\begin{cases}y_{i}&\text{if $i\leq N$,}\\ \hat{v}_{n}(x_{i})&\text{else.}\end{cases}$
Again, we have $\|v_{n}\|_{\mathrm{L}^{\infty}}\leq C$ and, using the arguments of [90, Proposition 4.24], we can show that $(\mu_{n},v_{n})\to(\mu,v)$ in $\mathrm{TL}^{p}(\Omega)$. Since $v_{n}(x_{i})=\ell_{i}$ for all $i\leq N$, in order to show (73), it therefore is sufficient to show that
$\lim_{n\to\infty}\left(\underbrace{\mathcal{F}_{n,\varepsilon_{n}}^{(k,p)}(\mu_{n},v_{n})-\mathcal{E}_{n,\varepsilon_{n}}^{(k,p)}(\hat{v}_{n})}_{=:T_{2}}\right)=0.$
We estimate as follows:
$|T_{2}|$ $\leq\frac{1}{n^{k+1}\varepsilon_{n}^{p+kd}}\sum_{i_{0},\cdots,i_{k}=1}^{n}\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{i_{j}}-x_{i_{r}}|}{\varepsilon_{n}}\right)\right]\left||v_{n}(x_{i_{1}})-v_{n}(x_{i_{0}})|^{p}-|\hat{v}_{n}(x_{i_{1}})-\hat{v}_{n}(x_{i_{0}})|^{p}\right|$
By definition of $v_{n}$, for $(i_{0},\cdots,i_{k})\in S:=\{(i_{0},\cdots,i_{k})\,|\,N\leq i_{j}\leq n\text{ for all $0\leq j\leq 1$}\}$, the corresponding term in the above sum vanishes. This means that we need to consider all indices in
$S^{c}=\{(i_{0},\cdots,i_{k})\,|\,\text{there exists $0\leq j\leq 1$ such that $1\leq i_{j}\leq N$}\}.$
Summing over all sets in $S^{c}$ therefore yields:
$T_{2}$ $\leq\frac{1}{n^{k+1}\varepsilon_{n}^{p+kd}}\sum_{t=0}^{1}\sum_{i_{t}=1}^{N}\sum_{\begin{subarray}{c}i_{s}=1\\ s\neq t\end{subarray}}^{n}\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{i_{j}}-x_{i_{r}}|}{\varepsilon_{n}}\right)\right]\left||v_{n}(x_{i_{1}})-v_{n}(x_{i_{0}})|^{p}-|\hat{v}_{n}(x_{i_{1}})-\hat{v}_{n}(x_{i_{0}})|^{p}\right|$
(75) $\leq\frac{C}{n\varepsilon_{n}^{p}}\sum_{t=0}^{1}\sum_{i_{t}=1}^{N}\frac{1}{n^{k}\varepsilon_{n}^{dk}}\sum_{\begin{subarray}{c}i_{s}=1\\ s\neq t\end{subarray}}^{n}\left[\prod_{j=1}^{k}\prod_{r=0}^{j-1}\eta\left(\frac{|x_{i_{j}}-x_{i_{r}}|}{\varepsilon_{n}}\right)\right]$
$\leq\frac{C}{n\varepsilon_{n}^{p}}\sum_{t=0}^{1}\sum_{i_{t}=1}^{N}\frac{1}{n^{k}\varepsilon_{n}^{dk}}\sum_{\begin{subarray}{c}i_{s}=1\\ s\neq t\end{subarray}}^{n}\left[\prod_{j=1}^{k}\eta\left(\frac{|x_{i_{j}}-x_{i_{j-1}}|}{\varepsilon_{n}}\right)\right].$
For $t\in\{0,1\}$, using $\eta(s)=0$ for all $|s|>1$,
$\frac{1}{n^{k}\varepsilon_{n}^{dk}}\sum_{\begin{subarray}{c}i_{s}=1\\ s\neq t\end{subarray}}^{n}\left[\prod_{j=1}^{k}\eta\left(\frac{|x_{i_{j}}-x_{i_{j-1}}|}{\varepsilon_{n}}\right)\right]$
$\leq\frac{\eta(0)^{k}}{n^{k}\varepsilon_{n}^{dk}}\#\{(i_{0},\cdots,i_{t-1},i_{t+1},\cdots,i_{k})\,|\,|x_{i_{j}}-x_{i_{j-1}}|<\varepsilon_{n}\text{ for $0\leq j\leq k$}\}.$
Now, for an element in $(i_{0},\cdots,i_{k})\in\{(i_{0},\cdots,i_{t-1},i_{t+1},\cdots,i_{k})\,|\,|x_{i_{j}}-x_{i_{j-1}}|<\varepsilon_{n}\text{ for $0\leq j\leq k$}\}=:\hat{S}$, we have $x_{i_{t-1}}\in B(x_{i_{t}},\varepsilon_{n})$, $x_{i_{t-2}}\in B(x_{i_{t-1}},\varepsilon_{n})$ until $x_{i_{0}}\in B(x_{i_{1}},\varepsilon_{n})$ as well as $x_{i_{t+1}}\in B(x_{i_{t}},\varepsilon_{n})$, $x_{i_{t+2}}\in B(x_{i_{t+1}},\varepsilon_{n})$ until $x_{i_{k}}\in B(x_{i_{k-1}},\varepsilon_{n})$. Hence $x_{i_{j}}\in B(x_{i_{t}},k\varepsilon_{n})$ for all $j$. This shows that
$\#\hat{S}$ $\leq\sum_{\begin{subarray}{c}z_{1},\cdots,z_{k}\in\Omega_{n}\end{subarray}}\prod_{j=1}^{k}\mathds{1}_{B(x_{i_{t}},bk\varepsilon_{n})}(z_{j})$
$=\left(n\mu_{n}(B(x_{i_{t}},k\varepsilon_{n}))\right)^{k}.$
Using the latter, we continue estimating:
$\frac{1}{n^{k}\varepsilon_{n}^{dk}}\sum_{\begin{subarray}{c}i_{s}=1\\ s\neq t\end{subarray}}^{n}\prod_{j=1}^{k}\eta\left(\frac{|x_{i_{j}}-x_{i_{j-1}}|}{\varepsilon_{n}}\right)$ $\leq\frac{\eta(0)^{k}}{n^{k}\varepsilon_{n}^{dk}}\#\hat{S}$
$\leq C\left(\varepsilon_{n}^{-d}\mu_{n}(x_{i_{t}},k\varepsilon_{n})\right)^{k}$
$=C\left(\varepsilon_{n}^{-d}\int_{\Omega}\mathds{1}_{\{|T_{n}(x)-x_{i_{t}}|<k\varepsilon_{n}\}}\rho(x)\,\mathrm{d}x\right)^{k}$
$\leq C\left(\varepsilon_{n}^{-d}\int_{\Omega}\mathds{1}_{\{|x-x_{i_{t}}|<k\varepsilon_{n}-\|T_{n}-\mathrm{Id}\|_{\mathrm{L}^{\infty}}\}}\rho(x)\,\mathrm{d}x\right)^{k}$
(76) $\leq C\left(\textrm{Vol}(B(0,1))\left(\frac{k\varepsilon_{n}-\|T_{n}-\mathrm{Id}\|_{\mathrm{L}^{\infty}}}{\varepsilon_{n}}\right)^{d}\right)^{k}$
(77) $\leq C$
where we used Assumption M.2 for (76) and Assumption L.2 for (77).
Inserting (77) in (75), we obtain that
$T_{2}$ $\leq\frac{C}{n\varepsilon_{n}^{p}}$
from which we deduce that $T_{2}\to 0$ and (73). ∎
The next corollary is the analogue to Corollary 4.14 and is proved in the same manner.
Corollary 4.17 ($\limsup$-inequality for the sum of semi-supervised energies in the ill-posed case).
Assume that S.1, M.1, M.2, W.1, D.1, D.2 and L.2 hold. Assume that $n\varepsilon_{n}^{p}\to\infty$. Then, $\mathbb{P}$-a.s., for every $(\nu,v)\in\mathrm{TL}^{p}(\Omega)$, there exists $\{(\nu_{n},v_{{}_{n}})\}_{n=1}^{\infty}$ with $(\nu_{n},v_{n})\to(\nu,v)$ in $\mathrm{TL}^{p}(\Omega)$ such that:
$\limsup_{n\to\infty}\left(\mathcal{S}\mathcal{F}\right)_{n,\varepsilon_{n}}^{(k,p)}((\nu_{n},v_{n}))\leq\left(\mathcal{S}\mathcal{G}\right)_{\infty}^{(k,p)}((\nu,v)).$
We conclude with a lemma summarizing our $\Gamma$-convergence results for our semi-supervised objectives.
Lemma 4.18 ($\Gamma$-convergence of energies).
Assume that S.1, M.1, M.2, W.1, D.1, D.2 and L.2 hold. If $n\varepsilon_{n}^{p}\to 0$, then, $\mathbb{P}$-a.e, $(\mathcal{S}\mathcal{F})_{n,\varepsilon_{n}}^{(k,p)}$ $\Gamma$-converges to $(\mathcal{S}\mathcal{F})_{\infty}^{k,p}$ in $\mathrm{TL}^{p}(\Omega)$. If $n\varepsilon_{n}^{p}\to\infty$, then, $\mathbb{P}$-a.s., $(\mathcal{S}\mathcal{F})_{n,\varepsilon_{n}}^{(k,p)}$ $\Gamma$-converges to $(\mathcal{S}\mathcal{G})_{\infty}^{k,p}$ in $\mathrm{TL}^{p}(\Omega)$.
Compactness and proof of Theorem 3.4
The following lemma is inspired by [90].
Lemma 4.19 (Uniform bound of energies of minimizers).
Assume that S.1, M.1, M.2, W.1, D.1, D.2 and L.2 hold. Let $(\mu_{n},u_{n})$ be minimizers of $(\mathcal{S}\mathcal{F})_{n,\varepsilon_{n}}^{(q,p)}$. Then, $\mathbb{P}$-a.s., there exists $C>0$ such that
$\sup_{n>0}(\mathcal{S}\mathcal{F})_{n,\varepsilon_{n}}^{(q,p)}((\mu_{n},u_{n}))\leq C.$
Proof.
With probability one, we can assume that the conclusions of Lemma 4.18 hold.
In the proof $C>0$ will denote a constant that can be arbitrarily large, is independent of $n$, and that may change from line to line. We are going to follow the proof of [90, Lemma 4.25]
Let $v\in\mathrm{C}^{\infty}_{\mathrm{c}}(\Omega)$ be a function that interpolates the points $\{(x_{i},\ell_{i})\}_{i=1}^{\infty}$. Then, $v\in\mathrm{W}^{1,p}(\Omega)$ so in particular, by Assumption M.2, there exists $C_{0}$ such that
$(\mathcal{S}\mathcal{F})_{\infty}^{(q,p)}((\mu,v))<C_{0}.$
By Lemma 4.18, we can pick a recovery sequence $\{v_{n}\}_{n=1}^{\infty}$ for $v$ such that:
$\lim_{n\to\infty}h_{n}$ $:=\lim_{n\to\infty}\sup_{m\geq n}(\mathcal{S}\mathcal{F})^{(q,p)}_{m,\varepsilon_{m}}((\mu_{m},v_{m}))$
$=\limsup_{n\to\infty}(\mathcal{S}\mathcal{F})_{n,\varepsilon_{n}}^{(q,p)}((\mu_{n},v_{n}))$
$\leq(\mathcal{S}\mathcal{F})_{\infty}^{(q,p)}((\mu,v))$
$<C_{0}$
(since $(\mathcal{S}\mathcal{F})_{\infty}^{(q,p)}((\mu,v))=(\mathcal{S}\mathcal{G})_{\infty}^{(q,p)}((\mu,v))$). Let $h:=\limsup_{n\to\infty}(\mathcal{S}\mathcal{F})^{(q,p)}_{n,\varepsilon_{n}}((\mu_{n},v_{n}))$ and let $\bar{\varepsilon}=C_{0}-h>0$. Then, there exists $n_{0}$ such that for all $n\geq n_{0}$, $h_{n}-h<\bar{\varepsilon}/2$, which means that
$h_{n}=\sup_{m\geq n}(\mathcal{S}\mathcal{F})^{(q,p)}_{m,\varepsilon_{m}}(v_{m})<h+\bar{\varepsilon}/2<C_{0}.$
Using the latter, we have
$\sup_{n>0}(\mathcal{S}\mathcal{F})^{(q,p)}_{n,\varepsilon_{n}}((\mu_{n},v_{n}))$ $=\hskip-1.42262pt\max\hskip-2.27621pt\left\{(\mathcal{S}\mathcal{F})^{(q,p)}_{1,\varepsilon_{1}}((\mu_{1},v_{1})),\dots,(\mathcal{S}\mathcal{F})^{(q,p)}_{n_{0},\varepsilon_{n_{0}}}((\mu_{n_{0}},v_{n_{0}})),\sup_{n\geq n_{0}}(\mathcal{S}\mathcal{F})^{(q,p)}_{n,\varepsilon_{n}}((\mu_{n},v_{n}))\right\}$
$\leq C.$
Since $\{u_{n}\}_{n=1}^{\infty}$ are minimizers, we use $(\mathcal{S}\mathcal{F})^{(q,p)}_{n,\varepsilon_{n}}((\mu_{n},u_{n}))\leq(\mathcal{S}\mathcal{F})^{(q,p)}_{n,\varepsilon_{n}}((\mu_{n},v_{n}))$ to conclude. ∎
Proof of Theorem 3.4.
With probability one, we can assume that the conclusions of Lemmas 4.18 and 4.19 hold.
In the proof $C>0$ will denote a constant that can be arbitrarily large, is independent of $n$, and that may change from line to line.
By Lemma 4.19, we know that there exists $C$ such that $\sup_{n>0}(\mathcal{S}\mathcal{F})_{n,\varepsilon_{n}}^{(k,p)}((\mu_{n},u_{n}))<C$ and in particular, $\mathcal{E}_{n,\varepsilon_{n}}^{(1,p)}(u_{n})$ is uniformly bounded. Furthermore, analogously to what is described in the proof of [80, Theorem 2.1], we know that $\sup_{n>0}\|u_{n}\|_{\mathrm{L}^{\infty}}<C$ with probability 1. We can therefore apply [80, Proposition 4.4] to obtain a subsequence $\{n_{r}\}_{r=1}^{\infty}$ and $(\mu,u)\in\mathrm{TL}^{p}(\Omega)$ such that $(\mu_{n_{r}},u_{n_{r}})\to(\mu,u)$ in $\mathrm{TL}^{p}(\Omega)$.
  1. 1.
    Since $n\varepsilon^{p}_{n}\to 0$, by [80, Lemma 4.5], we know that $u$ is continuous and, for every $\Omega^{\prime}\subset\subset\Omega$, we have that $\max_{\{s\leq n_{r}\,|\,x_{s}\in\Omega^{\prime}\}}|u(x_{s})-u_{n_{r}}(x_{s})|\to 0$ and, with probability 1, $u(x_{i})=\ell_{i}$ for all $i\leq N$. By Lemma 4.18 and Proposition 2.6, we also have that $(\mu,u)$ is a minimizer of $(\mathcal{S}\mathcal{F})_{\infty}^{(q,p)}$. Finally, by the uniqueness of the minimizer of $(\mathcal{S}\mathcal{F})_{\infty}^{(q,p)}$, we conclude that the whole sequence $(\mu_{n},u_{n})$ converges to $(\mu,n)$ in $\mathrm{TL}^{p}(\Omega)$ and for every $\Omega^{\prime}\subset\subset\Omega$, we have that $\max_{\{s\leq n\,|\,x_{s}\in\Omega^{\prime}\}}|u(x_{s})-u_{n}(x_{s})|\to 0$.
  2. 2.
    By Lemma 4.18, Proposition 2.6 and the assumption $n\varepsilon_{n}^{p}\to\infty$, $(\mu,u)$ is a minimizer of $(\mathcal{S}\mathcal{G})_{\infty}^{(k,p)}$. ∎
4.3 Higher-order hypergraph learning
The proofs in this section are simple corollaries from the results in [90]. In contrast to the discrete-continuum nonlocal-continuum local decomposition used for the proofs in Section 4, everything in this section relies on spectral convergence results between the discrete Laplace operators $\Delta_{n,\varepsilon_{n}}$ and its continuum counterpart $\Delta_{\rho}$.
For our first result, the proof follows from an application of [90, Proposition 4.21]. The key observation is that for any $\varepsilon_{n}$, $p>0$ and $v\in\mathrm{C}^{\infty}(\bar{\Omega})$ with $v(x_{i})=y_{i}$ for $i\leq N$, we can pick $\{(\nu_{n},v_{n})\}_{n=1}^{\infty}=\{(\mu_{n},v|_{\Omega_{n}})\}_{n=1}^{\infty}$ and hence the same recovery sequence for $\mathcal{J}^{(p_{k})}_{n,\Delta_{n,\varepsilon_{n}^{(k)}}}$ with $1\leq k\leq q$. This allows us to use the subadditivity of $\limsup$ to deduce the result.
Proposition 4.20 ($\limsup$-inequality for the sum of semi-supervised energies in the well-posed case).
Assume that S.2, M.1, M.2, W.1, D.1 and D.2 hold. Let $q\geq 1$, $P=\{p_{k}\}_{k=1}^{q}\subseteq\mathbb{R}$ with $p_{1}\leq\cdots\leq p_{q}$ and $E_{n}=\{\varepsilon_{n}^{(k)}\}_{k=1}^{q}$ with $\varepsilon_{n}^{(1)}>\cdots>\varepsilon_{n}^{(q)}$. Assume that $\varepsilon_{n}^{(q)}$ satisfies L.3 and that $\rho\in\mathrm{C}^{\infty}$. Then, $\mathbb{P}$-a.s., for every $(\nu,v)\in\mathrm{TL}^{2}(\Omega)$, there exists a sequence $\{(\nu_{n},v_{{}_{n}})\}_{n=1}^{\infty}$ with $(\nu_{n},v_{n})\to(\nu,v)$ in $\mathrm{TL}^{2}(\Omega)$ and
$\limsup_{n\to\infty}(\mathcal{S}\mathcal{J})_{n,E_{n}}^{(q,P)}((\nu_{n},v_{n}))\leq(\mathcal{S}\mathcal{J})_{\infty}^{(q,P)}((\nu,v)).$
The next result is shown analogously to Proposition 4.15. In particular, one relies on the compactness result [90, Proposition 4.13]: we only require that our smallest length-scale $\varepsilon_{n}^{(q)}$ satisfies the appropriate upper bound and that the its associated power $p_{q}$ scales correctly with the dimension of $\Omega$. Then, the problem reduces to using the superadditivity of $\liminf$ and [90, Theorem 4.14].
Proposition 4.21 ($\liminf$-inequality for the sum of semi-supervised energies in the well-posed case).
Assume that S.2, M.1, M.2, W.1, D.1 and D.2 hold. Let $q\geq 1$, $P=\{p_{k}\}_{k=1}^{q}\subseteq\mathbb{R}$ with $p_{1}\leq\cdots\leq p_{q}$ and $E_{n}=\{\varepsilon_{n}^{(k)}\}_{k=1}^{q}$ with $\varepsilon_{n}^{(1)}>\cdots>\varepsilon_{n}^{(q)}$. Assume that $\varepsilon_{n}^{(q)}$ satisfies L.3, that $n\cdot(\varepsilon_{n}^{(q)})^{p_{q}/2-1/2}$ is bounded and that $p_{q}>\frac{5}{2}d+4$. Then, $\mathbb{P}$-a.s., for every sequence $\{(\nu_{n},v_{{}_{n}})\}_{n=1}^{\infty}\subseteq\mathrm{TL}^{2}(\Omega)$ with $(\nu_{n},v_{n})\to(\nu,v)$ in $\mathrm{TL}^{2}(\Omega)$, we have
$\liminf_{n\to\infty}(\mathcal{S}\mathcal{J})_{n,E_{n}}^{(q,P)}((\nu_{n},v_{n}))\geq(\mathcal{S}\mathcal{J})_{\infty}^{(q,P)}((\nu,v)).$
For the next result, we again rely on the fact that [90, Proposition 4.24] implies that the same recovery can be chosen for all $\mathcal{J}_{n,\Delta_{n,\varepsilon_{n}^{(k)}}}^{(p_{k})}$ with $1\leq k\leq q$. In particular, we need to assume that all $\varepsilon_{n}^{(k)}$ satisfy an appropriate lower bound and that their associated powers $p_{k}$ scale correctly with the dimension of $\Omega$. We then conclude using the subadditivity of $\limsup$.
Proposition 4.22 ($\limsup$-inequality for the sum of semi-supervised energies in the ill-posed case).
Assume that S.2, M.1, M.2, W.1, D.1 and D.2 hold. Let $q\geq 1$, $P=\{p_{k}\}_{k=1}^{q}\subseteq\mathbb{R}$ with $p_{1}\leq\cdots\leq p_{q}$ and $E_{n}=\{\varepsilon_{n}^{(k)}\}_{k=1}^{q}$ with $\varepsilon_{n}^{(1)}>\cdots>\varepsilon_{n}^{(q)}$. Assume that $\rho\in\mathrm{C}^{\infty}$ and that for $\varepsilon_{n}^{(q)}$ satisfies L.2 as well as $n(\varepsilon_{n}^{(q)})^{2p_{q}}\to\infty$. Then, $\mathbb{P}$-a.s., for every $(\nu,v)\in\mathrm{TL}^{2}(\Omega)$, there exists a sequence $\{(\nu_{n},v_{{}_{n}})\}_{n=1}^{\infty}$ with $(\nu_{n},v_{n})\to(\nu,v)$ in $\mathrm{TL}^{2}(\Omega)$ and
$\limsup_{n\to\infty}(\mathcal{S}\mathcal{J})_{n,E_{n}}^{(q,P)}((\nu_{n},v_{n}))\leq(\mathcal{S}\mathcal{K})_{\infty}^{(q,P)}((\nu,v)).$
Summarizing all our previous results and using the subadditivity of $\liminf$ in conjunction with [90, Proposition 4.22], we obtain the following result.
Lemma 4.23 ($\Gamma$-convergence of energies).
Assume that S.2, M.1, M.2, W.1, D.1, D.2 hold. Let $q\geq 1$, $P=\{p_{k}\}_{k=1}^{q}\subseteq\mathbb{R}$ with $p_{1}\leq\cdots\leq p_{q}$ and $E_{n}=\{\varepsilon_{n}^{(k)}\}_{k=1}^{q}$ with $\varepsilon_{n}^{(1)}>\cdots>\varepsilon_{n}^{(q)}$. Assume that $\rho\in\mathrm{C}^{\infty}$.
  1. 1.
    Assume that $\varepsilon_{n}^{(q)}$ satisfies L.3, that $n\cdot(\varepsilon_{n}^{(q)})^{p_{q}/2-1/2}$ is bounded and that $p_{q}>\frac{5}{2}d+4$. Then, $\mathbb{P}$-a.s., $(\mathcal{S}\mathcal{J})_{n,E_{n}}^{(q,P)}$ $\Gamma$-converges to $(\mathcal{S}\mathcal{J})_{\infty}^{(q,P)}$.
  2. 2.
    Assume that $\varepsilon_{n}^{(q)}$ satisfies L.2 as well as $n(\varepsilon_{n}^{(q)})^{2p_{q}}\to\infty$. Then, $\mathbb{P}$-a.s., $(\mathcal{S}\mathcal{J})_{n,E_{n}}^{(q,P)}$ $\Gamma$-converges to $(\mathcal{S}\mathcal{K})_{\infty}^{(q,P)}$.
Proof of Theorem 3.5.
The proof is analogous to the proof of Theorem 3.4.
In particular, for the well-posed case, we use [90, Proposition 4.17 and Lemma 4.25] to obtain a uniform bound on the $\mathrm{L}^{2}$-norms of $u_{n}$. Then, uniform and $\mathrm{TL}^{2}$-convergence of a subsequence of $u_{n}$ to some continuous $u$ follows from [90, Proposition 4.13 and Theorem 4.14]. By the uniqueness of the minimizer, Lemma 4.23 and Proposition 2.6, the result follows.
For the ill-posed case, convergence of a subsequence in $\mathrm{TL}^{2}$ to some $u$ follows from [90, Theorem 4.14]. Again, Lemma 4.23 and Proposition 2.6 allow us to conclude. ∎
5 Numerical Experiments
Multiscale Laplace learning has demonstrated strong empirical performance on point cloud data, outperforming many existing graph-based semi-supervised learning methods [61]. Building on this, and given that we approximate HOHL using this framework, our evaluation focuses on sensitivity analyses. In particular, we show that choosing exponents $p_{\ell}=\ell$ leads to improved performance over constant-exponent settings, highlighting the benefit of applying higher-order regularization at finer scales.
In particular, we include the following experiments on four datasets of various sizes and difficulty: iris [29], digits [4], Salinas A [1], MNIST [56]. We summarize all the notation used in Table 2.
$q$-Experiment.
For $1\leq q\leq 5$, we test (6) with various configurations of weight and power coefficients. In particular, first, we pick $\varepsilon^{(1)}\geq\varepsilon^{(2)}\geq\varepsilon^{(3)}\geq\varepsilon^{(4)}\geq\varepsilon^{(5)}$ and build the Laplacian matrices $\Delta_{n,\varepsilon^{(\ell)}}$ for $1\leq\ell\leq 5$. Then, we setup different models with constant coefficients $\lambda_{\ell}=1$ (CC), slowly increasing coefficients $\lambda_{\ell}=\ell$ (SC) or quickly increasing coefficients $\lambda_{\ell}=\ell^{2}$ (QC) as well as constant powers $p_{\ell}=1$ (CP) or increasing powers $p_{\ell}=\ell$ for $1\leq\ell\leq 4$ (IP). The aim of this experiment is to analyze the performance of our model as a function of $q$.
$j$-Experiment.
For $1\leq q\leq 3$, we pick $\varepsilon_{n}^{(1)}\geq\varepsilon_{n}^{(2)}\geq\varepsilon_{n}^{(3)}$ and build the Laplacian matrices $\Delta_{n,\varepsilon^{(\ell)}}$ for $1\leq\ell\leq 3$. Then, we set up different models for $1\leq j\leq 4$ with $\lambda_{1}=1$, $\lambda_{2}=j^{2}$, $\lambda_{3}=(j+1)^{2}$ (VQC($q$) – we let $q\in\{2,3\}$) as well as $p_{1}=1$, $p_{2}=2$ and $p_{3}=3$. The aim of this experiment is to compare the performance of our model as a function of the coefficients $\lambda_{\ell}$ for a fixed $q$.
Term / Abbreviation $q$-Experiment $j$-Experiment
Aim of experiment Analysis of HOHL as a function of maximum powers $q$ Analysis of HOHL as a function of coefficients $\lambda_{\ell}$
$\ell$ Index over scales $1\leq\ell\leq q$ Same meaning
$q$ Number of Laplacians $1\leq q\leq 5$ Number of Laplacians $2\leq q\leq 3$
$j$ Coefficients $\lambda_{\ell}$ are a function of parameter $1\leq j\leq 4$
$\varepsilon^{(\ell)}$ Scale for $\ell$-th $\varepsilon$-graph Laplacian Same meaning
$k^{(\ell)}$ Scale for $\ell$-th kkNN-graph Laplacian Same meaning
$\Delta_{n,\varepsilon^{(\ell)}}$ $\ell$-th $\varepsilon$-graph Laplacian Same meaning
$\lambda_{\ell}$ Fixed or increasing ($1$ or $\ell$ or $\ell^{2}$) Varies with $j$: $\lambda_{1}=1,\lambda_{2}=j^{2},\lambda_{3}=(j+1)^{2}$
$p_{\ell}$ Constant or increasing ($1$ or $\ell$) Increasing: $p_{\ell}=\ell$
$C$ $\lambda_{\ell}=1$
SC $\lambda_{\ell}=\ell$
QC $\lambda_{\ell}=\ell^{2}$
CP $p_{\ell}=1$
IP $p_{\ell}=\ell$ $p_{l}=\ell$
VQC($q$) For $q=2$: $\lambda_{1}=1$, $\lambda_{2}=j^{2}$. For $q=3$: $\lambda_{1}=1$, $\lambda_{2}=j^{2}$, $\lambda_{3}=(j+1)^{2}$.
Table 2: Terminology used in the $q$- and $j$-experiments.
We always use the full dataset as nodes in the graph construction. For MNIST, we use the same data embedding as in [16]. For the smaller datasets, iris and digits, we use $\varepsilon$-graphs with weights $w_{\varepsilon^{(\ell)},ij}=\exp\left(\frac{-4|x_{i}-x_{j}|^{2}}{\left(\varepsilon^{(\ell)}\right)^{2}}\right)$. To illustrate that our model works with different (hyper)graph types and in order to speed-up computations, we rely on $k$-nearest neighbors (kkNN) graphs for the large datasets (naturally substituting the sequence $\varepsilon^{(1)}\geq\varepsilon^{(2)}\geq\varepsilon^{(3)}\geq\varepsilon^{(4)}\geq\varepsilon^{(5)}$ with $k^{(1)}\geq k^{(2)}\geq k^{(3)}\geq k^{(4)}\geq k^{(5)}$) with weights $w_{k^{(\ell)},ij}=\exp\left(\frac{-4|x_{i}-x_{j}|^{2}}{d_{k^{(\ell)}}(x_{i})^{2}}\right)$ where $d_{k^{(\ell)}}(x_{i})$ denotes the distance from $x_{i}$ to its $k^{(\ell)}$-th nearest neighbor. Each experiment is conducted over 100 trials and we report the mean accuracy and standard deviation (in brackets) of our results in percentages. In particular, for each trial we re-sample labelled points to be used as fixed constraints in the learning (the same constraints as in (2)). We vary the labelling rate from $0.02$ to $0.8$ (for the Salinas A dataset, the rate parameter goes from 1 to 100 and denotes the number of labelled points per class).
Since we show that hypergraph learning problem and HOHL behave asymptotically like graph problems in Theorems 3.3, 3.4 and 3.5 (see Figure 2), it is relevant to consider other graph algorithms for fair comparisons: Laplace learning [99], Poisson learning [16], Fractional Laplace (FL) learning [90] with $s=2$ and $s=3$ (only on iris), Weighted Nonlocal Laplacian (WNLL) [75], $p$-Laplace learning [30], Random Walk (RW) [97], Centered Kernel (CK) [59], Sparse LP (SLP) [54] and Properly Weighted Graph-Laplacian [18]. Lastly, we note that no methodical hyperparameter optimization has been performed for the choice of $\varepsilon^{(\ell)}$ and $k^{(\ell)}$. The result highlights are displayed in Tables 3, 4, 5, 6, 7 and 8. The complete results can be found in Appendix 7.
We note that our proposed models mostly outperform the other graph SSL models, especially when the labelling rate, $q$ and $j$ are large enough (see also Tables 10, 12, 14 and 16). This signifies that the multiscale/hypergraph structure can be efficiently leveraged when solving a discrete learning problem on point clouds. Furthermore, our method is robust with respect to the choice of parameters $p_{\ell}$ and $\lambda_{\ell}$.
The model configurations with increasing powers (IP) outperform all other versions of our model (5) (see also Tables 9, 11, 13 and 15). This confirms the idea that the choice of $p_{\ell}=\ell$ is more efficient than $p_{\ell}=1$. In particular, this means that our proposed HOHL model (5)/(6) is more effective than the hypergraph learning model (2) because of higher-regularization.
Finally, we note that a choice of $q=2,3$ is often sufficient to have a significant increase in performance (see Tables 9, 11 and 13) which is analogous to observations in [61]. This is an important finding as higher values of $q$ imply a notable increase in computation time: for each additional $\varepsilon^{(\ell)}$ or $k^{(\ell)}$, one has to compute a new set of hyperedges/a new Laplacian as well as an increasingly costly matrix product. A similar statement can be made for $j$ (see Tables 10 and 12): increasing the weights $\lambda_{\ell}$ is effective up to a certain point at which the performance of our models does not improve. For this reason, we only consider a restricted set of $q$ and $j$ when dealing with larger datasets.
Rate Laplace Poisson IP-QC CP-QC IP-SC CP-SC IP-$C$ CP-$C$
0.02 11.96 (4.03) 78.81 (2.98) 22.57 (9.14) 15.02 (5.8) 20.91 (8.57) 15.46 (5.4) 18.96 (7.82) 13.79 (5.43)
0.05 19.35 (6.62) 84.87 (1.63) 61.81 (7.17) 37.24 (7.55) 58.56 (7.5) 31.54 (9.11) 52.93 (7.74) 24.84 (7.85)
0.10 42.87 (7.4) 87.13 (1.12) 81.57 (3.51) 60.04 (7.23) 80.78 (3.71) 54.66 (7.07) 78.93 (4.26) 50.4 (6.7)
0.20 68.58 (4.38) 87.61 (0.94) 89.12 (1.5) 85.79 (2.17) 89.06 (1.5) 82.83 (2.57) 88.82 (1.47) 79.01 (3.19)
0.30 82.1 (2.02) 87.58 (0.74) 91.74 (0.87) 90.98 (1.02) 91.75 (0.87) 89.44 (1.15) 91.73 (0.88) 87.57 (1.28)
0.50 88.3 (1.11) 87.85 (0.78) 93.87 (0.71) 93.39 (0.81) 93.89 (0.7) 92.45 (0.86) 93.89 (0.71) 91.37 (0.92)
0.80 89.73 (1.43) 87.88 (1.42) 94.98 (0.99) 94.33 (1.13) 94.96 (0.98) 93.3 (1.2) 94.91 (0.96) 92.18 (1.21)
Table 3: Accuracy of various SSL methods on the digits dataset for the $q$-experiment with $q=3$. We pick $\varepsilon^{(\ell)}=100^{2-\ell}$ for $1\leq\ell\leq 5$. Proposed methods are in bold.
Rate Laplace Poisson WNLL Properly $p$-Lap RW CK IP-VQC (2) IP-VQC (3)
0.02 12.20 (4.75) 79.00 (2.75) 67.07 (6.07) 78.29 (3.14) 77.83 (3.23) 30.17 (11.33) 60.00 (4.17) 25.16 (9.35) 24.25 (9.65)
0.05 20.42 (7.03) 84.61 (1.72) 69.20 (4.38) 83.11 (2.08) 82.50 (2.19) 32.00 (5.96) 66.19 (3.73) 62.69 (6.84) 61.96 (6.85)
0.10 41.62 (6.59) 86.73 (1.36) 80.73 (3.07) 87.67 (1.45) 87.45 (1.51) 31.95 (5.56) 71.98 (2.73) 81.51 (3.66) 81.25 (3.61)
0.20 68.47 (4.79) 87.61 (0.99) 86.21 (1.53) 89.04 (0.97) 88.93 (1.00) 40.94 (4.75) 78.25 (1.53) 89.49 (1.09) 89.41 (1.10)
0.30 82.17 (2.32) 87.62 (0.80) 88.00 (1.20) 89.81 (0.87) 89.74 (0.89) 44.89 (5.34) 82.11 (0.81) 91.83 (0.86) 91.79 (0.83)
0.50 88.18 (1.00) 87.84 (0.96) 89.04 (1.00) 89.98 (1.00) 89.94 (0.99) 37.33 (2.51) 85.67 (0.98) 93.79 (0.91) 93.77 (0.90)
0.80 89.65 (1.49) 87.88 (1.40) 89.68 (1.45) 89.97 (1.42) 89.97 (1.41) 33.93 (1.16) 88.34 (1.39) 94.91 (1.01) 94.93 (1.00)
Table 4: Accuracy of various SSL methods on the digits dataset for the $j$-experiment with $j=2$. We pick $\varepsilon^{(\ell)}=100^{2-\ell}$ for $1\leq\ell\leq 5$. Proposed methods are in bold.
Rate Laplace Poisson IP-QC CP-QC IP-SC CP-SC IP-$C$ CP-$C$
1 58.08 (8.37) 57.12 (7.32) 60.98 (7.28) 59.25 (7.54) 59.73 (7.89) 59.00 (7.85) 58.81 (8.09) 58.67 (8.07)
2 66.85 (5.49) 57.32 (6.44) 67.75 (5.42) 67.45 (5.44) 67.26 (5.60) 67.32 (5.46) 66.85 (5.77) 67.22 (5.46)
5 73.46 (2.31) 56.83 (5.31) 73.59 (2.36) 73.65 (2.35) 73.61 (2.42) 73.63 (2.34) 73.58 (2.48) 73.59 (2.27)
10 75.86 (1.82) 56.08 (5.31) 76.09 (1.88) 76.21 (1.81) 76.15 (1.84) 76.14 (1.83) 76.15 (1.84) 76.06 (1.83)
20 77.61 (1.15) 56.20 (4.25) 78.52 (1.51) 78.14 (1.18) 78.42 (1.43) 78.02 (1.17) 78.26 (1.31) 77.87 (1.15)
50 79.60 (0.88) 56.44 (3.93) 80.95 (0.91) 80.37 (0.89) 80.83 (0.94) 80.18 (0.90) 80.64 (0.93) 80.00 (0.90)
100 80.86 (0.57) 56.06 (2.98) 82.47 (0.70) 81.82 (0.56) 82.33 (0.62) 81.61 (0.56) 82.10 (0.61) 81.35 (0.55)
Table 5: Accuracy of various SSL methods on the Salinas A dataset for the $q$-experiment with $q=3$. We pick $k^{(1)}=50$, $k^{(2)}=30$, $k^{(3)}=20$ and $k^{(4)}=10$. Proposed methods are in bold.
Rate Laplace Poisson WNLL Properly $p$-Lap RW CK IP-VQC (2) IP-VQC (3)
1 59.28 (8.54) 58.31 (6.46) 64.13 (6.05) 64.10 (6.04) 60.26 (5.44) 63.10 (5.14) 28.50 (5.98) 61.88 (7.11) 62.23 (6.78)
2 66.82 (5.35) 56.76 (7.03) 67.54 (5.04) 67.42 (5.10) 64.65 (5.13) 66.94 (4.76) 33.05 (6.65) 67.53 (5.07) 67.68 (5.12)
5 73.74 (2.71) 55.56 (5.89) 73.42 (3.07) 73.14 (3.15) 72.26 (3.07) 73.70 (2.60) 46.37 (5.32) 73.94 (2.84) 73.86 (2.85)
10 75.88 (1.67) 56.49 (5.18) 75.81 (1.73) 75.32 (1.81) 74.80 (1.85) 75.98 (1.73) 55.54 (4.27) 76.23 (1.76) 76.14 (1.81)
20 77.44 (1.37) 55.99 (4.62) 78.23 (1.40) 77.56 (1.58) 77.51 (1.61) 77.99 (1.22) 66.04 (3.10) 78.34 (1.31) 78.40 (1.37)
50 79.58 (0.94) 56.69 (4.19) 80.87 (0.90) 80.21 (0.93) 80.36 (0.88) 79.10 (0.85) 75.21 (1.76) 80.87 (0.98) 80.98 (1.01)
100 80.96 (0.73) 55.83 (2.75) 82.10 (0.63) 81.88 (0.66) 82.12 (0.61) 79.27 (0.70) 79.82 (1.02) 82.41 (0.72) 82.53 (0.75)
Table 6: Accuracy of various SSL methods on the Salinas A dataset for the $j$-experiment with $j=2$. We pick $k^{(1)}=50$, $k^{(2)}=30$, $k^{(3)}=20$ and $k^{(4)}=10$. Proposed methods are in bold.
Rate Laplace Poisson IP-QC CP-QC IP-SC CP-SC IP-$C$ CP-$C$
0.02 97.07 (0.07) 96.80 (0.06) 97.36 (0.07) 97.29 (0.08) 97.26 (0.07) 97.24 (0.07) 97.19 (0.07) 97.19 (0.07)
0.05 97.37 (0.05) 96.85 (0.04) 97.64 (0.06) 97.59 (0.06) 97.56 (0.05) 97.54 (0.06) 97.50 (0.05) 97.48 (0.06)
0.10 97.58 (0.04) 96.85 (0.04) 97.82 (0.04) 97.77 (0.04) 97.76 (0.04) 97.74 (0.04) 97.70 (0.04) 97.69 (0.04)
0.20 97.81 (0.04) 96.87 (0.04) 98.01 (0.04) 97.98 (0.04) 97.97 (0.04) 97.95 (0.04) 97.92 (0.04) 97.91 (0.04)
0.30 97.92 (0.04) 96.87 (0.05) 98.10 (0.04) 98.07 (0.04) 98.07 (0.04) 98.05 (0.04) 98.02 (0.04) 98.02 (0.05)
0.50 98.08 (0.06) 96.87 (0.08) 98.24 (0.06) 98.21 (0.06) 98.21 (0.06) 98.19 (0.06) 98.18 (0.06) 98.17 (0.06)
0.80 98.25 (0.09) 96.90 (0.12) 98.38 (0.09) 98.36 (0.10) 98.37 (0.09) 98.34 (0.09) 98.34 (0.09) 98.32 (0.09)
Table 7: Accuracy of various SSL methods on the MNIST dataset for the $q$-experiment with $q=3$. We pick $k^{(\ell)}=30-(\ell-1)\cdot 10$ for $1\leq\ell\leq 3$. Proposed methods are in bold.
Rate Laplace Poisson WNLL Properly $p$-Lap RW CK IP-VQC (2) IP-VQC (3)
0.02 97.06 (0.09) 96.79 (0.07) 96.55 (0.09) 94.76 (0.17) 94.48 (0.17) 97.15 (0.10) 95.34 (0.16) 97.31 (0.09) 97.34 (0.09)
0.05 97.37 (0.06) 96.85 (0.05) 97.20 (0.05) 94.49 (0.12) 95.49 (0.10) 97.37 (0.07) 96.46 (0.08) 97.62 (0.05) 97.64 (0.05)
0.10 97.59 (0.04) 96.86 (0.04) 97.58 (0.05) 95.59 (0.08) 96.88 (0.06) 97.45 (0.05) 97.18 (0.06) 97.80 (0.04) 97.82 (0.04)
0.20 97.80 (0.04) 96.87 (0.04) 97.86 (0.04) 97.08 (0.05) 97.71 (0.04) 97.50 (0.05) 97.68 (0.04) 97.99 (0.04) 98.00 (0.04)
0.30 97.92 (0.05) 96.87 (0.05) 97.98 (0.05) 97.61 (0.06) 97.88 (0.05) 97.51 (0.05) 97.88 (0.05) 98.10 (0.05) 98.10 (0.05)
0.50 98.08 (0.06) 96.86 (0.06) 98.11 (0.06) 98.01 (0.06) 98.07 (0.06) 97.51 (0.06) 98.09 (0.06) 98.24 (0.06) 98.24 (0.05)
0.80 98.22 (0.10) 96.87 (0.14) 98.23 (0.10) 98.22 (0.11) 98.23 (0.10) 97.52 (0.13) 98.24 (0.11) 98.37 (0.11) 98.37 (0.11)
Table 8: Accuracy of various SSL methods on the MNIST dataset for the $j$-experiment with $j=2$. We pick $k^{(\ell)}=30-(\ell-1)\cdot 10$ for $1\leq\ell\leq 3$. Proposed methods are in bold.
6 Conclusion
We have conducted a rigorous continuum analysis of hypergraph-based semi-supervised learning, revealing that classical hypergraph models—despite their combinatorial expressiveness—converge asymptotically to first-order graph-based methods. Specifically, through both pointwise and variational convergence analyses, we established that these discrete hypergraph models approach weighted Sobolev $\mathrm{W}^{1,p}$ regularization in the limit. Crucially, our results identify a previously undocumented discrepancy between the operator obtained via pointwise convergence and that derived from the variational limit. This divergence underscores the importance of considering both modes of convergence when analyzing discrete-to-continuum limits in graph- and hypergraph-based learning. Furthermore, we demonstrated that the transition between meaningful regularization and trivial smoothing is sensitive to the parameters chosen in the hypergraph construction.
To address the inherent limitations of classical hypergraph approaches, we introduced HOHL—a variational model that enforces higher-order smoothness by penalizing powers of graph Laplacians at multiple scales derived from the hypergraph structure. Our theoretical results characterize the well- and ill-posedness of HOHL, and show that it converges to genuinely higher-order Sobolev-type energies. Additionally, we argued that for point clouds embedded in a metric space, multiscale Laplacian learning is an appropriate surrogate for HOHL, thereby providing a principled theoretical foundation for this latter class of methods. Extensions to more general hypergraph settings, along with algorithmic and computational developments, are discussed in [88].
Empirical evaluations confirm the practical effectiveness of HOHL: by leveraging multiscale structure and higher-order regularization, the model achieves strong performance across a range of standard SSL benchmarks.
By analyzing hypergraph learning within a unified continuum framework, our work also offers a systematic classification of regularization-based SSL algorithms (see Figure 2) and opens new avenues for principled model design grounded in asymptotic analysis. Promising future directions include extending our framework to additional hypergraph-based methods [28, 58, 32, 50, 72].